A Bibliography of Algol 68 Publications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 May 2026 Version 1.09

Title word cross-reference

(a, b) [581, 1069]. (a, bγ) [1069]. (a, bγ − γ) [1069]. (a, bσ) [581]. (a, bσ − σ) [581]. 25 [754]. 29 [813]. 3 [272]. 68 [796, 855]. 9 [379]. Ax = λBx [616]. cos(x)/x [257, 488]. d [425]. e^{x}/x [257, 488]. Eig3 [162]. F [340]. J [214, 215]. λ [1201]. LR [617]. n [568, 75]. N ≤ 8 [1091]. Eig [277]. QD [153]. QR [267]. sin(x)/x [257, 488]. t [492, 493, 495, 496, 680, 1003, 1004, 936]. U(a, b, z) [1054]. X [303].

-Algorithm [153]. -calculus [1201]. -Distribution [492, 495, 1003, 936, 340]. -method [267]. -multiprogramming [339]. -Quantiles [493, 496, 1004]. -R [796, 855]. -th [75].

0.5 [430]. 0.90 [750]. 06 [541]. 06-07.01.1972 [541]. 07.01.1972 [541].

1 [784, 1116, 88, 41, 1079, 1105, 1120, 1094, 1097, 693, 1286, 284]. 10 [607]. 1108 [732]. 11e [483]. 11th [483]. 123 [303]. 14.00 [966]. 15 [320]. 1538-1972 [621]. 16 [74]. 163 [122]. 179 [312]. 180 [303]. 181 [303]. 191 [724, 140]. 192 [139]. 1960 [9]. 1969 [381, 1277]. 1970 [1278]. 1972 [707, 708]. 1974 [746, 1280]. 1975 [970, 971, 1282]. 1976 [1289, 801, 1285, 1287]. 1977 [1288]. 1978 [1291, 970, 971]. 1979 [1292, 970, 971]. 1981-1982 [1058].

2 [1214, 114, 390, 1216, 693, 1227]. 2.1 [256, 294, 1109, 1282, 1287]. 20-31 [746]. 2005 [1225]. 209 [303]. 222 [164]. 225 [159]. 226 [160, 303]. 22nd [1302, 1285]. 24 [373]. 244 [181]. 25 [36, 1282, 541]. 25-26.09.1993 [1177]. 252 [216]. 259 [210]. 25th [1285, 541]. 26 [218]. 260 [214]. 261 [215]. 268 [231, 431]. 272 [303, 229, 356]. 282 [257]. 29-30 [1277]. 291 [266]. 292 [407]. 299 [301, 361]. 2nd [1299, 750, 1255]. 2nd-class [1255].

3 [1224, 961, 78, 1126]. 3-486-21551-5 [961]. 30 [1287]. 300 [299, 700]. 301 [287, 313]. 304 [332, 302, 303]. 31.01-04.02.2005 [1221]. 3100 [490, 550]. 314 [605]. 322 [340]. 335 [338]. 338 [365]. 339 [366, 433]. 34 [716]. 345 [432]. 349 [440]. 35 [79]. 36 [46]. 360 [507, 567, 630, 631]. 363 [396, 626]. 367 [606]. 37 [33]. 370/158 [797]. 38 [34]. 39 [67]. 395 [492, 495, 936]. 396 [493, 496, 936]. 3e [348]. 3rd [1279].

4 [248, 829, 169, 634, 635, 185, 360, 63]. 40 [60]. 404 [565]. 409 [579]. 411 [570]. 412 [538, 701]. 414 [547]. 426 [602, 709]. 440 [673]. 442 [679]. 445 [668]. 447 [681]. 455 [682]. 457 [667]. 459 [696, 776]. 464 [692]. 465 [680]. 469 [685]. 47 [392]. 471 [674]. 48 [640]. 480 [726]. 482 [727]. 49 [641]. 4th [441].

5 [961]. 50 [258]. 5500 [350]. 56 [691]. 57 [51]. 58 [141].

6 [612, 214]. 6- [214]. 6.50 [723]. 60 [247, 452, 1116, 326, 24, 1213, 1214, 664, 10, 11, 111, 112, 786, 704, 706, 788, 824, 1275, 528, 335, 530, 29, 82, 288, 466, 115, 752, 202, 37, 605, 606, 89, 794, 839, 903, 13, 40, 41, 42, 90, 91, 116, 118, 120, 483, 93, 121, 1249, 797, 545, 546, 94, 1079, 1105, 1120, 258, 125, 717, 348, 349, 166, 167, 298, 905, 217, 677, 850, 402, 552, 678, 721, 491, 763, 764, 260, 101, 1107, 54, 1064, 56, 57, 1047, 222, 58, 804, 805, 132, 59, 178, 306, 868, 503, 627, 1216, 684, 410, 504]. 60 [226, 227, 628, 180, 309, 231, 17, 102, 103, 137, 138, 236, 918, 1012, 1194, 421, 422, 511, 424, 363, 364, 1097, 1274, 189, 1209, 142, 143, 191, 144, 68, 431, 580, 238, 145, 71, 73, 960, 273, 76, 195, 738, 890, 927, 107, 241, 149, 656, 820, 324, 244, 448, 21, 152, 929, 1126, 221, 1138]. 60-System [324]. 6000 [657]. 61 [26]. 6500 [465, 532]. 66 [784, 693, 1286]. 67 [783, 801, 297, 318]. 68 [744, 1301, 745, 453, 525, 663, 783, 784, 1264, 327, 328, 382, 785, 455, 456, 595, 526, 458, 459, 597, 665, 461, 826, 385, 463, 464, 386, 387, 1159, 827, 791, 828, 899, 934, 389, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476, 533, 534, 535, 601, 792, 391, 603, 1160, 1270, 756, 900, 713, 830, 902, 833, 539, 540, 1076, 1061, 996, 837, 757, 1271, 842, 395, 485, 904, 972, 801, 398, 399, 400, 489, 548, 549, 718, 846, 941, 974, 1002, 1294, 847, 676, 720, 848, 1046, 1030, 853, 943, 854, 494].

68 [351, 497, 857, 858, 859, 977, 978, 1005, 404, 623, 866, 354, 768, 867, 409, 1163, 1189, 560, 410, 504, 947, 725, 411, 505, 1094, 412, 563, 629, 869, 1165, 1166, 564, 308, 357, 358, 413, 506, 1265, 1266, 414, 415, 508, 569, 632, 1083, 912, 913, 914, 950, 982, 418, 810, 728, 952, 420, 1052, 953, 811, 871, 954, 983, 1096, 362, 573, 638, 731, 919, 872, 873, 874, 693, 1281, 1286, 1268, 513, 577, 514, 515, 957, 958, 775, 516, 878, 435, 436, 733, 879, 816, 923, 882, 1251, 650, 883, 585, 1019, 737, 887]. 68 [653, 447, 523, 589, 377, 524, 659, 660, 661, 739, 1055, 1071, 592, 1087, 1272, 997, 962, 367, 368, 369, 370, 371, 442, 443, 444, 520, 588, 652, 778, 817, 818, 889, 926, 1127, 441, 518, 587, 884, 651, 594, 705, 966, 484, 500]. 68-modes [947]. 68-R [750, 663, 828, 539, 540, 904, 524, 659, 660, 739]. 68-Rt [953]. 68C [845]. 68’R [891, 825]. 6th [1295].

7.35 [740]. 7090 [211]. 74 [716]. 74-34 [716].

80 [772]. 85 [753]. 86 [773]. ’88 [1297, 84]. 89 [85].

9 [121, 215]. 9- [215]. 90 [86]. ’91 [1298]. ’95 [1302].

A. [1103]. A1 [685]. AAD [1019, 1012, 1169]. Aarhus [1295, 1295]. Abbildungen [961]. ABC [1150]. ABS12 [101]. Abstract [706, 844, 1106, 1007, 1166, 986, 1196, 790, 768, 1109, 1122]. Abstraction [1178, 1199, 1123, 1192]. Academic [723, 1058, 1115, 1144]. Acceptive [1045]. Acceptors [457]. access [1137]. Accessing [1029]. according [379]. account [919]. Accuracy [273, 159]. Accurate [688]. ACM [2, 602, 667, 538, 668, 673, 674, 547, 905, 492, 493, 495, 496, 679, 680, 681, 682, 55, 685, 565, 726, 727, 570, 7, 692, 921, 579, 696]. ACM-GAMM [2]. actions [110]. active [1196]. Activity [904]. Actuele [1222]. ADA [1019, 1225, 1107, 1094, 1037]. Adaptive [673, 1032, 906, 637]. Address [989, 994, 1012, 1014, 1019, 1022, 494]. Adjustment [513]. Administration [237]. Adressenberechnung [494]. Advanced [1283, 746]. Affix [559]. Affix-Grammars [559]. Affordable [1255]. Aflink [897]. Again [630]. Age [1259, 1131]. aggregate [952]. ähnlich [872]. AI [1073]. aid [96]. aims [12, 234]. Airy [287, 313]. Akademie [258]. Akademie-Verlag [258]. AL [715]. ALCOR [211, 192]. ALEPH [1001]. Alfons [300]. Algebra [760, 512, 1240, 735, 736]. Algebraic [2, 977, 1, 993]. algebras [1145]. ALGOL [1288, 379, 155, 1277, 594, 326, 785, 111, 705, 528, 792, 3, 605, 606, 41, 1271, 608, 609, 670, 92, 161, 545, 546, 487, 127, 128, 165, 398, 167, 171, 99, 494, 405, 909, 623, 1280, 867, 307, 728, 102, 360, 511, 1278, 872, 1285, 1245, 513, 647, 1084, 438, 439, 961, 282, 283, 740, 652, 926, 747, 822, 990, 751, 896, 966, 390, 708, 289, 336, 967, 1026, 87, 484, 345, 346, 800, 258, 221, 627, 807, 1266, 1259, 185, 311, 421, 422, 424, 1136, 732, 956, 1138, 193, 881, 322, 323, 324, 376]. Algol [1087, 997, 987, 247, 198, 378, 1301, 1154, 379, 745, 453, 22, 23, 110, 381, 783, 784, 24, 964, 1213, 1214, 328, 382, 25, 785, 111, 112, 113, 786, 455, 595, 458, 459, 597, 156, 12, 1275, 329, 330, 331, 384, 1238, 1242, 529, 461, 826, 335, 530, 30, 82, 288, 249, 385, 463, 464, 386, 387, 1159, 466, 827, 791, 389, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476, 601, 115, 391, 537, 710, 603, 202, 756, 900, 479, 158, 205, 713, 37, 540, 607, 251, 89, 839, 996, 757, 4, 13, 41]. ALGOL [42, 90, 91, 117, 118, 120, 252, 253, 670, 121, 206, 290, 291, 341, 342, 343, 796, 672, 45, 613, 545, 546, 395, 485, 486, 14, 94, 1079, 1105, 1120, 96, 162, 347, 255, 293, 845, 972, 163, 256, 294, 125, 349, 97, 126, 938, 296, 801, 297, 398, 399, 400, 489, 166, 49, 50, 298, 5, 905, 846, 941, 217, 1294, 847, 676, 720, 848, 850, 170, 99, 1030, 943, 218, 402, 552, 678, 854, 129, 403, 494, 763, 764, 130, 174, 260, 351, 101, 497, 619, 304, 620, 498, 859, 15, 54, 621, 1064, 352, 499, 978, 6]. ALGOL [56, 57, 1047, 222, 58, 223, 622, 305, 404, 225, 132, 946, 866, 624, 59, 178, 306, 354, 768, 868, 409, 503, 1216, 355, 561, 261, 1149, 410, 504, 226, 227, 179, 947, 1031, 411, 505, 180, 412, 869, 1165, 1166, 506, 688, 414, 415, 508, 632, 1083, 633, 309, 913, 914, 950, 982, 16, 1267, 231, 417, 418, 810, 728, 952, 420, 17, 102, 103, 135, 136, 137, 138, 183, 186, 233, 235, 236, 918, 1012, 511, 1034, 871, 983, 263, 265, 1096, 362, 573, 1278, 638, 731, 919, 574, 575, 7, 187, 363, 872, 1097, 642, 1274, 189]. ALGOL [874, 1014, 1201, 190, 141, 693, 1281, 1286, 65, 142, 143, 1245, 1268, 144, 192, 68, 431, 1037, 513, 314, 514, 515, 1282, 1287, 958, 105, 580, 238, 775, 145, 1178, 1192, 516, 436, 582, 922, 1018, 1053, 19, 269, 316, 733, 879, 71, 583, 882, 73, 194, 960, 1133, 1251, 650, 106, 271, 272, 883, 585, 274, 76, 1019, 146, 887, 277, 653, 279, 372, 445, 1070, 195, 738, 890, 927, 373, 148, 320, 107, 241, 77, 149, 891, 447, 820, 321, 151, 242, 243, 244, 280, 281, 284, 285, 523, 448, 152, 377, 524, 659]. ALGOL [661, 1055, 108, 197, 1272, 593, 246, 780, 962, 147, 275, 318, 367, 368, 369, 370, 371, 442, 443, 444, 520, 588, 652, 778, 817, 818, 889, 926, 1127, 441, 518, 587, 884, 319, 1023, 1196, 1199, 450, 1246, 1247, 452, 80, 81, 109, 154, 525, 663, 1116, 1264, 327, 248, 664, 894, 10, 11, 456, 526, 596, 704, 200, 706, 788, 824, 383, 599, 157, 665, 825, 333, 286, 114, 29, 897, 1244, 828, 899, 934, 533, 534, 535, 1180, 829, 752, 1160, 1187, 1270, 478]. Algol [83, 204, 1207, 830, 902, 669, 833, 539, 87, 88, 1076, 794, 903, 1061, 837, 1077, 1078, 40, 116, 483, 93, 1248, 1249, 842, 797, 612, 393, 254, 904, 716, 124, 292, 717, 348, 47, 350, 548, 549, 718, 168, 169, 1243, 974, 1002, 551, 1254, 300, 1046, 677, 259, 401, 852, 853, 1226, 721, 491, 554, 618, 100, 855, 857, 858, 1107, 556, 977, 1005, 683, 1065, 804, 805, 176, 406, 501, 1163, 1189, 560, 684, 1108, 807, 1067, 725, 628, 1094, 563, 629, 564, 308, 357, 358, 413, 1265, 569, 912, 981, 634, 635, 359]. Algol [1250, 1231, 1052, 133, 184, 1194, 1135, 421, 423, 1236, 1260, 1190, 1204, 953, 1035, 425, 811, 954, 920, 639, 364, 1125, 1205, 873, 188, 1209, 429, 576, 267, 191, 237, 577, 957, 876, 877, 581, 1069, 878, 365, 366, 432, 433, 437, 648, 985, 649, 270, 317, 816, 923, 1112, 1143, 239, 273, 737, 1102, 1173, 654, 655, 656, 1174, 375, 150, 589, 21, 196, 660, 739, 1071, 592, 1087, 929, 245, 449, 1126, 651, 1024, 250, 175, 723, 512, 740, 750, 707, 937, 172, 173, 500]. ALGOL- [405, 307]. Algol-60 [751, 336, 322, 786, 41, 118, 120, 1079, 1105, 1120, 503, 236, 1097, 491, 503, 452, 706, 788, 29, 752, 491, 804, 805]. ALGOL-60-Übersetzer [41]. Algol-60-Unterprogrammen [896]. ALGOL-68 [713, 653]. ALGOL-68-ähnlich [872]. ALGOL-68-like [872]. ALGOL-based [251, 613, 393, 829, 429]. ALGOL-E [624]. ALGOL-GENIUS [179]. ALGOL-like [1178, 1070, 126, 499, 529, 205, 946, 1031, 265, 190, 1192, 1133, 1135, 1180, 1187, 1207, 556, 1112, 1143]. ALGOL-procedure [65]. ALGOL-Procedures [379, 670, 350]. ALGOL-Programme [961, 271, 272]. ALGOL-Programming [316]. ALGOL-Prozeduren [379, 670]. Algol-Systems [376]. ALGOL-Verschwörung [1238, 1242]. Algol-W [981]. ALGOL-Wörterbuch [99]. Algol60 [928]. ALGOL68 [851, 1148, 984]. ALGOL68C [865]. algological [1228]. Algols [971, 915, 916, 917, 951, 970]. Algorithm [378, 153, 666, 536, 602, 667, 709, 538, 668, 834, 835, 836, 838, 795, 544, 843, 14, 759, 673, 674, 547, 616, 617, 906, 492, 493, 495, 496, 679, 680, 762, 1081, 681, 856, 766, 682, 860, 557, 622, 685, 911, 687, 565, 726, 568, 727, 570, 692, 813, 579, 696, 960, 590, 591, 798, 598, 790, 114, 531, 994, 996, 799, 615, 976, 134, 572, 645, 514, 697, 197, 379, 332, 287, 477, 709, 33, 34, 753, 392, 481, 36, 84, 85, 86, 159, 160, 338, 340, 122, 164, 210, 257, 396, 488, 46, 214, 215, 216, 299, 51, 301, 302]. Algorithm [303, 1003, 1004, 407, 626, 724, 686, 60, 771, 181, 229, 356, 772, 773, 231, 361, 691, 640, 641, 266, 312, 313, 63, 139, 140, 67, 431, 365, 366, 432, 433, 776, 440, 74, 272, 700, 701, 79, 936]. Algorithmen [379, 452, 670, 695]. Algorithmic [326, 327, 111, 786, 1025, 466, 603, 484, 1249, 1290, 623, 17, 103, 138, 1194, 104, 1293, 442, 443, 778, 817, 818, 1127, 10, 11, 112, 113, 605, 606, 91, 258, 98, 358, 236, 7, 1282, 1287, 585, 1021, 1272, 275, 318, 367, 368, 369, 370, 371, 444, 520, 588, 652, 889, 926, 991]. algorithmique [744, 603]. Algorithmische [111, 623, 258, 102, 314, 652, 926, 991]. Algorithms [452, 381, 753, 1077, 1078, 671, 1223, 543, 212, 942, 401, 764, 681, 772, 773, 691, 640, 641, 813, 20, 929, 379, 972, 763, 650, 78]. Algorithmus [598, 572, 697, 379, 754, 272]. algoritmicheskom [444, 520]. Algoritmichniyat [585]. aller [598]. allgemeinen [849, 779]. Allocation [535, 1032, 68, 389, 476, 58]. Alps [1245]. Also [705, 605, 606]. alternate [32]. alternating [1104]. Alternative [1092, 704]. am [1087]. ambiguity [402]. American [920]. among [988]. Amsterdam [746, 1058, 705, 740, 541, 1247, 1115, 1130, 1221, 1222, 1293]. analitik [1034, 1035]. Analyse [346, 438]. Analysis [467, 533, 969, 95, 123, 498, 682, 1039, 1220, 1129, 276, 347, 491, 220, 868, 503, 16, 265, 984, 708, 707]. Analytic [781]. Analyzing [1124]. and/or [649]. Andenken [963]. Andersen [114]. Andrew [937]. Angeles [1291]. angewandten [405]. Angles [1153]. Anglia [1285]. Anleitung [487]. anniversary [541]. annotation [840]. Announcement [20]. Annual [1300, 1298]. Anpassungsoperationen [513]. ANSI [1188]. antisymmetric [584, 697]. antisymmetrische [697]. antisymmetrischen [584]. APL [479, 1079, 1105, 1120, 1097, 984, 922, 1018, 1053]. Appendix [117].

Application [305, 628, 1071, 792, 996, 768, 72, 305]. Applications [423, 1285, 1269, 790, 347, 984, 986]. Applied [970, 971, 1273, 755, 394]. Approach [1042, 791, 981, 262, 734, 704, 993, 624, 1015]. approximants [766]. Approximate [510]. Approximation [547, 683, 1065, 730, 295, 975]. Approximations [296]. Arbeit [1245]. Arbitrary [771, 957, 366, 433, 440, 924, 36]. architecture [850]. architectures [1254]. Areas [748, 749]. Argument [204]. Arguments [210, 860, 392]. Arithmetic [47, 350, 974, 1002, 1279, 1295, 685, 1050, 189, 957, 273, 108, 929, 672, 53, 622, 1263, 445]. Arithmetik [800, 324]. Array [761]. arrays [976, 68]. Arsenal [108]. art [1156]. artificial [613, 519]. Aspects [964, 1187, 1273, 950, 473, 642]. assembler [825]. Assembly [529, 45]. assert [809]. Assertional [1219]. assertions [1020]. assessment [723]. assessments [1085]. assigning [810]. Assignment [14, 531, 862, 908, 694]. assignments [1173]. Associated [392]. Association [605, 66]. Associative [393]. Assumptions [521]. Asymptotic [84]. Atlas [304, 309]. ATOL [930]. Attempted [586]. attribute [878]. aufbau [572]. Aufl. [487]. Auflage [99]. Auflösung [452]. Augments [19]. August [1292, 1277, 1282, 1287]. Ausdrucksmittel [733]. Auswertung [491, 503]. Autocode [309]. Autocorrelation [641]. Automate [1073]. automated [72]. Automatic [528, 31, 712, 207, 208, 1048, 952, 183, 235, 639, 1254, 873, 698, 494, 271]. automating [711]. automation [1139, 715]. Automatische [494, 271]. automatisering [1139]. autres [434]. Available [542, 705]. avoiding [558, 625]. Axiom [935]. Axiomatic [1113, 1128].

B [605, 606, 258, 723, 500, 961, 528, 465, 532, 716, 350, 803]. B-5500 [350]. B-6500 [465, 532]. B-splines [803]. Babbage [1254]. Babuska [421, 511]. Babuska’sche [421, 422, 511]. Backus [605, 606, 258]. bacteriological [1228]. BAIL [774]. balanced [600, 843]. Band [528]. bandwidth [557]. barbara [1069, 581]. Based [393, 977, 560, 432, 989, 829, 251, 613, 1150, 429, 1098, 1017, 107]. Basic [1060, 338, 1239, 653, 53, 325, 607, 1079, 1105, 1120, 574, 575, 1097, 928]. basis [238]. Bauer [250, 605, 606, 258, 175]. Baumann [250, 175]. BC [131]. BCD [141]. Bd [405]. be [490, 550, 1008, 509]. Beam [702]. Beamstrukturierte [702]. beens [884]. before [1043]. beginner [1011]. beginning [793]. Begriffe [733]. Beispiel [1087]. Beliebigen [336]. BELL [271, 272, 271, 272]. BELL-Programmen [271, 272]. Bemer [38]. Benutzeranleitung [390]. Ber [51]. Berechnung [598]. berekenen [372]. Bericht [926, 111, 258, 623, 102, 652]. Berlin [528, 258, 867]. Beschreibung [424]. Besprungen [528]. Bessel [477, 960]. best [1225]. beta [164, 1082, 312]. Between [227, 241, 663, 988, 226, 519]. bewerteten [670]. beyond [1261]. Bibliography [525]. bij [1111]. Bilingual [868]. Binary [668, 686, 640, 641, 843]. Binding [1083]. biofilms [1228]. Biografie [1245]. biography [1245]. Bisection [63]. BIT [248, 114, 612, 169, 634, 635, 185]. Bitwise [70]. Block [1092, 980, 948, 770]. Block-Structured [1092, 948].

Blocks [955, 237, 56]. BNF [1124]. Bo [612]. Bob [38]. bombing [27]. Bond [750, 1084]. Bonus [636]. Book [594, 705, 750, 966, 707, 708, 250, 967, 1026, 605, 606, 608, 609, 484, 254, 486, 487, 970, 971, 346, 937, 171, 172, 173, 175, 221, 723, 500, 405, 1259, 512, 314, 1084, 1138, 269, 316, 438, 439, 193, 961, 740]. Boolean [612, 259, 54]. Boor [803]. boucles [363]. Bound [1102]. boundary [519]. bounded [1009]. Bounds [510, 375, 635]. Brailsford [967, 1026]. Brian [1218]. Briandais [28]. brief [901, 1161, 1164, 633, 1171]. Bronchitis [618]. Brook [1289]. browsing [1118]. Buchbesprechung [258]. Built [1243]. Bulgarian [585]. Bulletin [111, 290, 341, 342, 343, 111, 291]. Burger [1104]. Burroughs [1244, 350].

C [594, 705, 966, 605, 606, 608, 609, 258, 221, 500, 405, 316, 1079, 1105, 1120, 1097, 1171, 1099]. C.U.P [937]. C6 [365, 366, 432, 433]. CA [1291, 1297]. CAIN [27]. Calculating [517, 803, 134, 372]. Calculation [81, 1176, 83, 552, 678, 721, 803, 1152, 44, 494, 64, 75]. calculus [743, 1075, 1117, 614, 1204, 1201]. calibrating [333]. California [1302]. Call [241]. called [509]. CAMAC [784, 693, 1286]. CAMAC-IML [784, 693, 1286]. Cambridge [937, 1151]. Can [892, 1198, 1008]. CAP [1151]. Capabilities [895]. capacitated [687]. care [1183]. Carl [345]. Carl-Erik [345]. Carlo [35, 673]. Carolina [1300]. case [1235, 809, 821]. categories [1182]. Category [1180, 717]. Category-Theoretic [1180]. Causal [1262]. Causality [1262]. cautious [906]. CDC [490, 550, 657]. Center [1117, 1041, 1075, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222]. Centraal [797]. central [349]. centrale [349]. Centre [1040, 541, 746, 1059, 1293]. Centrum [1041, 1075, 1117, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 1169]. Centum [1059]. certain [856, 325]. Certification [248, 332, 477, 392, 407, 626, 724, 771, 634, 635, 134, 361, 63, 277]. cEVM [1061]. chance [1193]. Changing [898]. chaos [1185]. chapters [358]. Character [857, 401]. Characteristics [808, 755]. Charleston [1300]. chart [731, 73]. Chartreuse [1287]. Chebychev [579]. Chebyshev [295, 547]. Check [577, 106]. checker [551]. Checking [1029, 628, 233, 180, 16, 883]. checkout [883]. chemistry [89]. Chi [1081, 798, 301, 361]. Chi-Squared [1081, 798, 361, 301]. Chichester [1026]. Chinese [564]. Chisholm [766]. Choice [451]. Cholesky [646]. Chr [114]. Church [226, 227]. Circuits [696, 776]. Class [224, 228, 1255]. Classes [748, 749, 777, 485, 925]. Classroom [1068]. Clebsch [216]. Clenshaw [637]. Cliques [667]. CLU [1161]. Clustering [695]. co [1255]. co- [1255]. COBOL [1079, 1105, 1120, 633, 1097, 279]. Code [782, 969, 223, 560, 186, 657, 823, 1254, 676, 1149, 412, 1182, 1220]. coding [31]. coefficients [216, 67]. coercions [823, 470, 471]. coinduction [1253]. collaboration [991]. collateral [429]. Collection [535, 589, 454, 476, 395, 506, 880, 523]. Collector [490, 550, 569, 469, 414]. colloquium [1139]. Colorado [1292]. colours [1015]. Combinations [806]. Combinatorial [228]. combined [45]. Combining [16]. comes [1254]. commands [1000]. Comment [651, 37]. Comments [537, 56, 192, 417]. Committee [4, 7, 3]. Common [783, 801, 12]. Communications [605]. Companion [573, 638]. Comparative [1228, 737, 387]. compared [1037]. comparing [1079, 1105, 1120, 1097]. Comparison [784, 965, 75, 900, 479, 507, 633, 693, 1281, 1286, 923, 1281]. Compatible [767]. compilateur [474]. Compilation [363, 268, 349, 723, 948, 364, 821, 363]. Compile [1029]. Compiled [206]. Compiler [1292, 782, 155, 1283, 88, 57, 1047, 560, 1266, 235, 732, 191, 577, 322, 1213, 1214, 249, 474, 53, 225, 865, 768, 1216, 135, 136, 1098, 882, 883, 997, 800]. Compilern [336]. Compilers [480, 223, 689, 233, 424, 655, 820, 664, 704, 389, 61, 876]. Compiling [931, 610, 549, 54, 1124, 1114, 25, 514]. complementary [81, 83, 303]. complete [598]. completeness [1134]. Complex [204, 565, 108, 267, 396, 626]. complexity [790, 310]. composed [1033]. Composite [970, 971]. composition [475, 534, 651]. compound [581, 1069]. Compressed [689]. Computation [109, 528, 124, 581, 1069, 960, 1054]. Computational [752]. computations [1074]. Computer [378, 1040, 1041, 1059, 1075, 1117, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 895, 529, 205, 606, 970, 971, 1243, 760, 802, 52, 1279, 1295, 224, 804, 805, 1280, 1276, 1215, 1208, 262, 1259, 1286, 1304, 521, 657, 1086, 929, 319, 746, 1130, 989, 388, 89, 43, 1079, 1105, 1120, 850, 1284, 868, 1230, 1052, 1097, 1274, 814, 72, 963]. computer-oriented [43]. Computerentwicklung [963]. Computers [929, 1074, 99]. Computing [605, 1239, 296, 66, 1174, 662, 1139, 619, 726, 507]. CONA [894]. concept [989, 461]. Concepts [895, 765, 1211, 653, 872, 325]. concerning [945, 873]. concerns [1123]. Concurrent [1234, 1201, 1119]. conditional [54]. conditions [426]. Conference [1288, 1291, 1299, 1300, 1302, 1301, 1154, 1155, 747, 12, 1280, 1278, 1304, 1297, 1277, 905, 1285, 921, 1282, 1287, 740, 1089, 2]. Confidential [59]. confluent [1054, 139]. conic [1179]. Conjugate [942]. conservatively [1190]. considerations [380]. Considered [841, 847]. Conspiracy [1238, 1242]. Constraints [397, 979]. Constructing [212, 689, 45, 408]. Construction [1292, 1283, 1086, 249, 711, 843, 501, 1098]. Constructors [1033]. Constructs [935, 743]. Context [213, 458, 468, 601, 305, 426, 516]. Context-Free [213, 468, 601, 305, 516]. context-sensitive [458]. Continuations [1142, 1202]. Continuous [547, 295]. Contribution [80, 81, 109, 83, 124, 292, 168, 259, 401, 176, 184, 185, 270, 317, 239, 320, 242, 280, 281, 282, 245, 248, 634, 635]. contributions [1012]. Control [452, 327, 938, 939, 940, 1254, 1056, 1057]. control-flow [1056, 1057]. controlled [159]. Conventions [178, 246]. Convergence [536]. Conversational [894, 268]. Conversion [791, 271, 272]. Convex [754].

Convolution [432]. COPILOT [734]. CORAL [693, 1286, 784]. correction [1121]. Correctness [350, 1235]. Correlation [67]. Correspondence [156, 227, 981, 144, 647, 194, 1128, 226, 984]. Corrigenda [881]. Cosine [84, 86]. Cost [671]. Coulomb [481, 299, 407, 700]. Count [1233]. counting [1193]. coupling [216]. Course [1283, 716, 221, 982, 746, 1059, 1075, 1117, 1131, 1146, 1157, 1179, 1185, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1233, 1237, 596, 990, 913, 1098, 13, 483, 937]. cover [937]. Creation [1076, 1236]. Creed [125, 144]. Criteria [852]. Critical [965, 753, 633, 60]. criticisms [24]. Cross [453, 825, 391, 581, 1069]. cross-referencing [391]. CROUT [74]. cubic [759]. cultural [1139]. Culture [1246, 1250]. cultuurgeschiedenis [1139]. Current [1222]. curriculum [802, 814]. Cursus [13, 483]. Curtis [637]. Curve [302, 303, 579, 332]. CWI [1139]. cycles [670]. Czechoslovakia [1248].

D [967, 1026, 161, 937, 127, 128, 193, 1257]. D. [486]. D1 [673]. dagelijks [1212]. d’ALGOL [463, 464, 473, 474, 489, 504, 516, 435]. Dallas [1279]. Dangling [247, 132]. Daniel [127, 128, 161]. dans [349, 363]. DAP [1077, 1078]. DAP-Algol [1077, 1078]. Darstellung [909]. Dasgupta [1259]. Data [1106, 618, 1093, 179, 1094, 690, 141, 695, 879, 1041, 930, 788, 30, 844, 615, 938, 1108, 877, 1133, 986, 885]. Database [895]. Datastructuren [885]. Datenorganisation [879]. day [1139]. debugger [774, 1056, 1057]. decidable [1231]. Decision [335, 1005, 978, 279]. declarations [514]. Decomposition [478, 296, 1263]. DECsystem [607]. DECsystem-10 [607]. Defect [1121]. Defense [42, 261, 195]. defined [401]. Defining [394, 14, 429]. Definite [168, 169, 779, 557]. definiten [779]. Definition [1147, 177, 914, 956, 586, 243, 284, 285, 603, 1043, 1294, 410, 283, 603, 375]. Definitions [1090, 373, 892, 255, 293, 1008]. Dekker [512]. delivered [1144]. demands [706, 752]. demonstration [910]. Denmark [1295]. Dennis [1258]. Denotation [466]. Denotational [954, 1072, 1009]. Denver [1292]. Department [1280]. dependent [1232]. dereferencing [862, 908]. derivations [994]. Derivatives [833, 488, 960, 257]. descent [1043]. described [713, 239]. Describing [264]. Description [526, 1275, 383, 599, 528, 543, 131, 887, 781, 901, 540, 994, 402, 507, 633, 878, 435, 593, 240]. Descriptions [212]. Design [788, 205, 119, 1235, 262, 1167, 1191, 1086, 745, 992, 388, 993, 846, 864, 868, 1011, 135, 136, 919, 1037, 240, 971, 970]. designational [148]. detecting [670]. Detection [969]. determination [831, 832]. determinism [1183]. deux [305]. d’évaluation [463, 464]. developed [1228]. Developing [799]. Development [1025, 1296, 1077, 1078, 1240, 320, 242, 280, 281, 282, 991, 918, 920, 963]. Developments [1151, 613]. Deviate [679]. Device [636, 1125]. device-independent [1125]. Dezimale [800]. diagnosis [96]. diagnostic [876]. diagram [32]. Diagrams [378]. Diagraphen [670]. Dialect [791, 15]. dialects [304, 1268]. dialogovaëiìa [1061]. dictionary [99]. DIDYMUS [828]. Diego [1297]. Dies [1208]. Difference [241, 1117]. Differences [663, 984]. different [1228, 799]. differential [755, 576, 581, 1069, 698]. differentierekening [1117]. Digital [89, 99, 1276, 1217]. Digital-Rechenanlagen [99]. digraphs [670]. Dijkstra [1213, 1214, 1216, 1208]. Dimensional [695, 566, 1137, 1104]. Diplomarbeit [956]. Directed [57, 1047, 530, 409]. direction [1104]. directions [1282, 1287]. discover [461]. Discovering [1258]. Discrete [579, 775]. Discussion [527, 548]. display [1135, 437]. displays [1080]. Dissertation [1258]. distributed [1177, 850, 1015]. Distribution [340, 303, 492, 495, 1003, 229, 773, 936, 160, 356]. distributions [643]. div [1147]. Diversity [892]. dlëiìa [1061]. DM [528]. DM.44 [961]. Do [892]. document [371]. Documentation [793, 137]. Documents [1241, 580]. Dominators [911]. don’t [1186, 1183]. double [1263, 581, 1069, 372, 445]. double-differential [581, 1069]. double-length [445]. Doubly [813, 637]. doubly-adaptive [637]. dr [348]. Draft [351, 436, 367, 1127, 358, 275, 318, 368, 369, 441]. Drawing [675, 849]. Driven [931, 690]. dubbele [372, 445]. dumping [883]. d’un [474]. Durch [956, 99]. during [1058, 1115, 1144]. Dutch [1175, 1041, 1059, 1075, 1117, 1130, 1131, 1139, 1146, 1156, 1157, 1179, 1185, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1233, 1237, 13, 483, 619, 1111, 372, 445, 1153, 885]. Dynamic [1032, 1140, 831, 832, 887].

E1 [726]. E2 [547, 579]. Early [1260]. East [1285]. Easter [1176, 1152]. Easy [1102]. EBCDIC [676]. Ed [740]. Edited [111]. Editing [618, 1150]. edition [483, 99]. Editor [1176, 231, 1246, 605, 1150, 431, 1271, 24, 25, 28, 202, 158, 37, 38, 39, 42, 44, 217, 222, 62, 141, 66, 76, 77, 149, 448]. edits [580]. edn [750, 966]. Edsger [1208]. Eduard [963, 963]. Éducation [346]. Eeuw [1131]. EFDM [1093]. effect [1255]. effectiveness [853]. effects [995, 1119]. Efficiencies [213]. Efficiency [714, 71]. Efficient [681, 722, 813, 237]. Effort [1248]. ego [1034, 1035]. Eigenschaften [398]. Eigenvalue [1063, 162, 277]. Eigenvalues [184, 185, 692, 267]. Eigenvectors [278, 184, 185]. eight [1074]. eighteenth [1298]. ëiìazyka [1061]. Eindhoven [1222]. eindige-volume [1111]. Einem [324, 598]. einfache [545, 546]. Einführung [872, 314, 718]. einheitlicher [909]. Einige [398]. Einsatz [751]. EISPACK [1063]. Ekman [345, 316]. elaborate [249]. electoral [762]. Electric [92, 93]. Electrologica [1213, 1214, 1216, 1235]. Elegant [1167]. Element [209, 903]. Elementary [254, 696, 776, 1133, 865]. Elements [450, 893]. elemzese [868]. Elimination [933, 949, 245]. Elliott [100, 130, 174]. Ellis [1026]. else [132, 247]. Embedded [1250]. Embracing [1248]. emulation [1235]. End [1265, 1264]. endlichen [572]. Engineering [1209]. England [1285]. English [92, 198, 93]. Enhanced [1253]. Enkele [1111, 1041]. Entrées [382]. entry [978]. Enumerability [777, 925]. Enumerable [777, 416, 315, 925, 886]. Environment [1095, 268, 461, 1118, 1150, 509]. Environmental [388]. Environments [1090]. epsilon [197]. equal [683, 1065]. equality [578]. Equation [452, 153, 836, 838]. equational [1183]. Equations [80, 666, 755, 835, 975, 1121, 856, 557, 576, 245, 1104, 698]. equidistant [759]. equipment [125, 144]. equivalence [600, 932]. equivalent [1009]. Equivalents [198]. Era [1208]. ERF [303]. Erik [345]. Erinnerung [111]. Errata [92, 282, 283]. Erratum [625, 185]. Error [452, 204, 396, 883, 485, 303, 626, 303, 510]. errors [29]. Erstellung [424]. erweiterte [99]. especially [12, 494, 984]. essence [1014]. Estimating [714, 654]. Estimation [618]. ETH [963, 963]. Euclidean [1147, 507]. EULER [283, 243, 284, 285, 1121]. EUROGRAPHICS [1089]. Europe [1236]. European [1221, 918, 1281]. evaluating [491, 503]. Evaluation [783, 801, 239, 819, 463, 464, 1177, 650, 995, 1119, 84, 85, 86]. even [1104]. Event [931, 775]. Event-Driven [931]. Events [1136]. every [315]. EVERYBODY [129]. everyday [1212]. EVM [1062]. examinations [1228]. Example [228, 1087]. examples [105]. Exception [1088, 1113, 1128, 1136]. Exception-Handling [1088]. Excerpt [846]. Execution [268, 910, 1082, 1230, 429, 654]. exemplarische [718]. exemplified [720]. Existence [315, 115]. Experience [88, 1050, 953, 661, 90, 1068]. Experimental [1206, 943, 624, 430]. Experimentele [1206]. explanations [105]. explicit [939, 940]. Exponential [674]. expressing [938]. Expression [84, 1113, 1128, 733, 239]. Expression-Oriented [1113, 1128]. Expressions [612, 189, 1196, 1199, 25, 259, 53, 54, 437, 148]. expressive [861, 907]. Extended [706, 804, 1093, 1166, 989, 99, 497, 805, 1108, 1263]. Extending [542, 1190]. Extensible [959, 586, 413, 729]. extension [531, 672, 101, 187, 190]. Extensions [1044, 4, 780]. Extrapolation [755].

F [250, 967, 1026, 605, 606, 258, 171, 175]. F2 [692]. face [1229]. Fachbegriffe [99]. facilitating [263]. facilities [930, 876]. Facit [88]. Facit-Algol [88]. factorization [1010, 646]. Factors [366, 433, 388]. facts [1040]. FAIL [742]. fair [1000]. fall [1074]. familiar [105]. Families [97]. far [1255]. Fast [942, 911, 730, 835, 365, 366, 432, 433]. Faults [554, 647]. Feasability [1107]. Feature [1217]. Features [360, 928, 757]. February [9]. Fehlererfassung [452]. Fehlerschranken [510]. Feliciano [250, 175]. Fermi [109, 124]. Few [1195, 448]. Fewer [629]. fibrational [1226]. Field [1029, 685, 1058, 1145, 1184]. Fields [924]. Fifty [1261, 1221]. fifty-second [1221]. figures [1040]. File [973, 1006]. Final [247, 801, 368]. Finding [153, 536, 667, 911, 165]. Finite [685, 947, 572, 1111]. Finitely [1031]. Finiteness [521]. First [477, 1249, 982, 755, 392, 937, 913, 1172]. Fit [579]. Five [1210, 655, 99]. fixed [789]. FL [1298]. Flächenstücken [849]. Floating [542, 688, 672, 98, 1235]. Floating-Point [542, 688, 672, 98, 1235]. Flow [671, 911, 1207, 1056, 1057]. flowchart [386, 968]. flowcharts [802]. flows [1111, 584]. Flüsse [584]. folding [566]. Ford [1218]. Foreword [1013]. form [32, 292]. Formal [1090, 1008, 177, 914, 243, 284, 285, 459, 1294, 947, 642, 434, 283, 240]. Format [1241]. Formats [781, 449]. formatted [900]. Formatting [437]. formels [434]. Formelübersetzung [8]. Former [1304]. Forms [397]. Formula [996, 18, 683, 1065, 1068, 8, 698]. formula-manipulation [698]. formulae [187]. Forsythe [1191]. FORTRAN [608, 609, 405, 909, 307, 647, 971, 378, 381, 28, 462, 530, 900, 479, 607, 89, 1079, 1105, 1120, 403, 619, 868, 355, 561, 633, 1034, 574, 575, 1097, 1037, 922, 1018, 1053, 279, 738, 890, 927, 321, 596, 822, 990, 897, 999, 716, 300, 554, 555, 915, 916, 917, 951, 1049, 1035, 873, 188, 876, 581, 1069, 985, 649, 1174, 970]. FORTRAN-Programmen [405, 307]. Foundations [746]. Four [350]. Fourier [365, 366, 432, 433, 220]. framework [1226]. France [1250]. Francisco [1302]. Fredholm [835, 975]. Free [213, 468, 601, 305, 516]. Frei [896]. French [328, 382, 463, 464, 473, 474, 603, 349, 489, 305, 504, 363, 516, 434]. Frequency [680]. frequently [29]. Fresnel [81, 83, 84, 85, 86, 181]. Freudenthal [378]. Fröberg [316, 345]. Front [1265, 1264]. Full [1199, 1178, 1192]. fuller [145]. fully [1196, 1122]. fun [1255]. FUNARG [509]. Function [287, 536, 396, 942, 51, 565, 773, 230, 640, 641, 266, 109, 36, 159, 160, 122, 124, 401, 303, 865, 626, 509, 313, 1054, 239, 509]. Functional [1093, 1108, 1190, 1110]. Functions [1147, 477, 481, 210, 1239, 547, 299, 552, 678, 721, 303, 860, 771, 229, 425, 264, 440, 960, 1202, 700, 392, 164, 295, 803, 491, 554, 407, 503, 356, 647, 995]. functor [1182]. functors [844]. fundamental [989]. fünf [99]. further [163].

G [594, 705, 750, 966, 608, 609, 970, 971, 165, 405, 314, 1084]. G7 [727]. Galerkin [835, 836, 838]. game [1196]. Games [1170, 1233]. GAMM [2, 7]. Gamma [565, 266, 159]. Ganzzahligen [754]. Garbage [454, 535, 490, 550, 569, 589, 469, 476, 395, 506, 414, 880, 523]. Gatlinburg [1240]. Gauss [303]. Gaussian [645, 245]. GCC [1264]. Gebrauch [487]. gegevens [1041]. geld [1233]. Gemischt [754]. Gemischt-Ganzzahligen [754]. General [469, 179, 197, 779, 611, 544, 1036, 429, 639]. Generalised [1223]. Generalization [222, 150, 151, 243, 284, 285, 202, 163, 283]. Generalized [353, 147, 831, 832, 64]. generate [1182]. Generating [834, 689, 976]. Generation [782, 766, 223, 657, 418, 952, 572]. generative [887]. generators [472]. Genes [1259]. GENIUS [179]. Gentrification [1255]. Geometric [827]. Geometry [1197, 1262, 1156]. German [307, 379, 452, 785, 111, 787, 598, 991, 1238, 1242, 528, 754, 41, 670, 545, 546, 398, 399, 718, 849, 99, 491, 553, 494, 623, 503, 949, 728, 102, 421, 510, 511, 572, 730, 427, 428, 872, 1245, 8, 513, 733, 879, 271, 272, 584, 697, 702, 375, 1087, 779, 963, 652, 926]. Germany [1282]. getal [1157]. gezicht [1229]. GIER [155, 135, 136, 235, 360]. given [1058, 645]. GKS [1083, 1087]. Glasgow [1288]. Gleichungssysteme [452, 406]. Gleitpunktrechnung [896]. global [684]. GLYPNIR [769]. GMFW [1139]. GMFW-CWI [1139]. GNU [1266]. go [558, 625]. Golden [1131]. gone [1255]. Good [935, 133]. Gordan [216]. Goto [133]. Gouden [1131]. GPS [134]. Grading [207, 208, 183]. grammaire [305, 516]. Grammaires [434]. Grammar [213, 1217, 115, 94, 1001, 863, 945, 305, 427, 428, 515, 516, 887, 337]. Grammars [559, 955, 457, 458, 1045, 719, 502, 416, 434, 886]. Grammatik [427, 428]. Graph [667, 538, 796, 681, 911, 696, 776, 701, 972, 1015, 984]. graph-analysis [984]. graph-theoretic [1015]. GRAPHEX68 [757]. Graphical [757]. Graphics [804, 1095, 805]. graphs [482]. Grau [528, 25]. gravimetrii [985]. Green [605, 606, 258]. Grew [1210]. grösserer [1087]. Group [378, 498, 682, 1281, 1282, 1287, 1221, 572, 192]. Grundlehren [528]. Gruppe [572]. Guaranteed [536]. guarded [1000]. Guessing [1170]. Guest [1246]. Guide [161, 127, 128, 420, 811, 193, 659, 739, 1055, 1079, 1105, 1120, 1097, 649, 660, 750, 1084].

H [594, 705, 991, 528, 966, 707, 708, 605, 606, 608, 609, 258, 500, 405, 681, 570, 696]. H. [991, 405]. H.M.S.O [750]. Hague [1218]. Half [88]. Halstead [1071]. Hammarling [1218]. Handbook [1079, 1105, 1120, 1276, 1097, 607, 528]. Handling [219, 359, 1088, 1113, 1128, 973, 1136]. Hankel [122]. Hans [608, 609]. hard [937]. hardback [1026]. Hardware [1023, 562, 125, 676, 848, 144]. harmful [847]. Hash [560]. Hash-Code [560]. Hashed [939, 940]. Hawgood [269]. Heap [455, 1024]. Heidelberg [528]. Heinz [1245, 1245]. held [746, 1115, 1285, 1282, 1287]. Hellgren [612]. Herausgegeben [111]. Herschel [487, 439]. heuristic [52]. Hewet [1059]. hierarchies [1027]. High [783, 801, 1050, 1245, 462, 944, 1006, 1051, 1056, 1057, 893]. high-level [462, 1051, 1056, 1057]. higher [720, 494]. higher-level [720]. Hill [528]. Historical [1254, 1175, 1214, 1216]. historisch [1175]. History [1291, 1299, 1158, 1303, 1148, 1166, 1267, 1301, 1154, 1139, 905, 1165, 921, 1022]. Hoare [935]. Hochalpen [1245]. Hoeken [1153]. Hoffman [512]. höhere [494, 909]. Holiday [1059, 1075, 1117, 1131, 1146, 1157, 1179, 1185, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1233, 1237]. Holland [705, 966, 740]. Hopcroft [795]. HOPL [1299, 1210]. HOPL-II [1299]. Hopscotch [1104]. Hor [1138]. Horwood [1026]. Householder [286, 334]. hun [1153]. Hungarian [868]. hybrid [553, 429]. hybride [553]. Hydra [390, 311]. Hydra-X8 [311]. HYGOL [553]. Hypergeometric [724, 139, 1054, 140]. HYPERLISP [1016]. hypotheses [1066].

I-O [256, 294]. I. [314]. I/O [1183, 246]. IAzyk [1034]. iazykami [1034]. ibid [606]. IBM [797, 507, 567, 630, 631, 1260]. ICL1900 [825]. ICR [182, 232]. Idealised [1201]. idealized [1196, 1199, 1231]. Identification [592, 400]. Identifier [466, 560]. identifiers [394]. IFIP [1278, 167, 1109, 1282, 1287, 656, 1022]. Ignorance [847]. Ihre [956]. II [1299, 1072, 1303, 41, 943, 619, 807, 1067, 227, 136, 285, 512]. III [155, 581, 1069]. Ilem [1102]. Illiac [769]. ILLINOIS [211, 1289, 1110]. illustrations [859]. Im [390, 779]. imaginary [392]. IML [784, 693, 1286]. Impact [1251]. Implementation [378, 1247, 964, 527, 825, 537, 92, 93, 1290, 1106, 765, 620, 1280, 560, 411, 505, 508, 1277, 1043, 330, 461, 388, 470, 471, 473, 829, 347, 399, 489, 855, 858, 978, 1006, 56, 866, 867, 308, 357, 506, 415, 632, 1052, 871, 1278, 1274, 583, 740, 740]. Implementations [965, 1268]. Implemented [490, 550, 850]. Implementierung [399, 956, 376]. Implementing [455, 1108, 1096, 972, 948, 694]. implicant [976]. implicit [1104]. importance [461]. Impossible [935]. improved [99]. Improvement [870, 1048]. improving [71]. in- [1051]. Includes [381]. including [745]. Incomplete [164, 312]. Incremental [268, 1217, 675]. Incrementally [610]. independence [1232]. Independent [119, 688, 276, 768, 1125, 888]. Indexes [1217]. inductive [1020]. Industrial [1286, 101]. industry [1221]. Inference [840]. infinite [418]. Influence [1240]. Informal [869, 441, 518, 587, 1277, 594, 705, 966]. Informatica [1041, 1075, 1117, 541, 1059, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222]. Information [898, 1207, 1021]. initial [755]. inner [507, 372, 374]. Input [338, 121, 166, 178, 690, 130, 656, 449, 382]. Input-Output [338, 166, 178, 656]. input/output [130, 449]. insbesondere [494]. instances [880]. instruction [851]. instructional [741]. instructions [694]. Instruments [1062]. Integer [754, 622]. Integral [204, 85, 86, 169, 301, 302, 303, 1081, 798, 332, 975, 683, 1065, 361]. Integrals [84, 674, 168, 181, 81, 83]. Integrated [1095]. integration [615]. Intellectual [521]. intelligence [613]. intensional [861, 907]. Interaction [1262, 1252]. Interactive [1090, 734, 1068, 1070, 1021, 1056, 1057]. Interconnecting [1129]. interest [44]. interface [897]. Interference [65, 1143]. intermediate [1182]. International [1284, 1280, 1304, 1293, 747, 2, 1, 1286]. Interpolating [552, 678, 721, 726, 803]. interpolation [787, 759, 401, 787]. Interpretation [6]. interpretative [978]. Interpreted [1031]. interpreter [530, 1149, 1110]. interpreters [871]. interrogation [650]. Intersection [305, 305]. Interval [350, 974, 1002, 1050, 929, 598, 491, 503, 598]. Intervall [324]. Intervall-Arithmetik [324]. Intervallarithmetik [751, 987]. Intervallmäßerung [503]. Intervallmäßige [491]. interview [1218, 1100]. Introducing [922, 1225]. Introduction [1154, 594, 157, 665, 385, 708, 1203, 345, 1294, 171, 725, 914, 435, 316, 21, 1246, 705, 387, 1159, 966, 718, 1161, 1164, 869, 234, 872, 105, 1171, 1101, 441, 518, 587, 707, 250, 175]. Introductory [899, 934, 967, 1026]. Intval [987]. invariant [683, 1065]. inverse [278]. inversion [39, 779]. investigations [1087]. Invited [989, 994, 1012, 1014, 1019, 1022]. involving [418]. inwendig [372]. IO [1027]. IO- [1027]. IPIP [740]. irasa [868]. ISBN [961]. ISO [621, 676, 412]. ISO-code [676, 412]. ISO/R [621]. Isolation [521]. Issues [1011, 146]. Italy [1177]. Iteration [209, 353, 1231, 278]. Iterative [286, 334, 75]. IV [769, 738, 890, 927].

J [605, 606, 486, 970, 971, 258, 171, 500, 512, 740]. J. [605, 606, 269]. J3B [783, 801]. J6 [538]. January [1297, 1298, 1302, 1300]. Java [1219]. Jazyk [1035]. jazykami [1035]. je [1233]. JEC [1061]. Journal [606]. JOVIAL [901, 801, 1037, 783, 649]. July [1278, 740, 801]. Jumps [1202]. June [1291, 1289, 1295, 1280]. Just [1205].

K. [254]. Kac [1115, 1144, 1145]. Kahan [421, 422, 511, 421]. Kahan-Babuska [421]. Kahan-Babuska’sche [421, 422]. Karlsruhe [311, 323]. Katz [605, 606, 258]. KDF [121]. KDF9 [92, 200, 537, 93, 172, 173]. KDF9ALGOL [419]. kegelsneden [1179]. kernel [975]. Kerner [258, 314]. Ketnyelvu [868]. key [99]. Kilter [697]. kind [392, 975, 477]. know [1186, 129]. Konvergenz [751]. konvexen [754]. Koster [500]. kunde [1156]. kunst [1156]. Kunzi [608, 609, 405]. kwadratische [1179].

L [250, 605, 606, 258, 175, 500, 740]. Laguerre [617]. lake [1228]. Lambda [614, 226, 227, 1204, 743]. Lambda-calculus [1204, 743]. Lambda-Notation [226, 227]. Lanczos [616]. langage [744, 603, 349, 305, 363, 364]. langages [305]. Langmaack [528]. Language [783, 784, 326, 327, 111, 786, 1090, 383, 599, 1025, 529, 965, 1180, 603, 205, 935, 484, 1249, 1188, 1290, 393, 801, 1273, 1162, 623, 179, 562, 567, 630, 231, 262, 1050, 17, 103, 138, 1194, 1236, 264, 1170, 1191, 104, 693, 1178, 653, 522, 658, 703, 1086, 1113, 1241, 442, 443, 778, 817, 818, 1127, 110, 930, 10, 11, 112, 113, 12, 991, 462, 388, 389, 1257, 482, 605, 606, 251, 1027, 837, 757, 91, 43, 613, 258, 349, 938, 48, 905, 846, 720, 677, 553, 621, 1064, 305, 864, 1008, 1082, 769, 358, 413]. language [631, 1011, 420, 102, 236, 871, 265, 7, 363, 872, 429, 921, 1286, 430, 431, 1017, 1192, 1085, 585, 735, 736, 1070, 195, 1021, 107, 149, 374, 741, 1272, 893, 240, 275, 318, 367, 368, 369, 370, 371, 444, 520, 588, 652, 889, 926, 1172, 2, 1, 872, 1281, 959, 1128]. Language-comparison [693, 1281]. language-oriented [735, 736]. Languages [1291, 1299, 1300, 1302, 247, 784, 1158, 1303, 480, 714, 119, 610, 543, 97, 126, 1092, 224, 1215, 1167, 1260, 1210, 644, 1211, 145, 1143, 777, 1174, 1297, 1298, 743, 745, 785, 458, 1043, 1186, 932, 993, 1207, 607, 1079, 1105, 1120, 973, 939, 940, 98, 905, 99, 944, 494, 52, 1006, 861, 907, 305, 979, 624, 868, 980, 948, 633, 1051, 1135, 7, 1097, 190, 921, 693, 1281, 1286, 1282, 1287, 814, 1112, 1133, 925, 886, 1183, 325, 1293, 519]. Langzahl [800, 956]. Langzahl-Triplex-Algol [956]. Large [537, 303, 646]. Large-scale [646]. Larger [210, 860, 1087]. last [918]. Later [1151, 1019]. Lattice [946]. Lattices [1010]. Lawes [486]. Lazy [1230]. learn [1198]. learner [874]. learner-programmers [874]. learning [72]. Least [286, 789, 296, 334, 836, 838, 165]. lectures [1058, 1074, 1144, 1145]. Left [819]. Legendre [860, 392, 210]. Leitfäden [405]. lengte [445]. length [445]. Letter [24, 25, 28, 1176, 202, 37, 38, 42, 44, 217, 222, 62, 141, 66, 76, 77, 149, 448]. Letters [9, 39]. Level [1050, 955, 462, 720, 1006, 502, 1051, 1056, 1057, 893]. leven [1212]. lex [253]. libraries [852, 1269]. Library [1076, 799, 1030]. Lie [378]. life [1212]. Lifetimes [824]. Like [1180, 126, 499, 529, 1187, 205, 1207, 45, 556, 946, 1031, 265, 872, 190, 1192, 1112, 1133, 1143, 1135, 1178, 1070]. limited [978]. Lindsey [594, 705, 966]. Line [981, 551]. line-syntax [551]. lineaire [885]. Linear [452, 80, 286, 397, 176, 1240, 666, 334, 835, 836, 838, 544, 165, 856, 571, 1204, 646, 735, 736, 245, 885]. Lineare [406]. linearer [452]. Linearity [1196]. lineinykh [985]. linguistic [1284]. LINUS [741]. LISP [633, 509, 1037, 995]. List [270, 200, 875, 274, 99]. list-processing [200]. lists [45]. literary [1284]. Literaturzitaten [627]. LITHP [274]. Litt [87]. Little [1210, 87]. LL [859]. LLGAUSS [114]. Local [472, 979, 1181, 1143, 411, 505, 1122, 1168]. Localities [808]. Locality [1247]. Locations [841]. Logarithm [266]. Logic [1141, 938, 861, 907, 864]. Logical [61, 842]. London [723]. long [643]. long-tailed [643]. look [932]. loop [968, 310]. Loops [717, 363]. Lorentz [683, 1065]. Lorentz-invariant [683, 1065]. Lost [1258]. low [333]. LPMODSIMPLEX [248, 634, 635]. LR [1223, 859, 408]. LTPL [1281]. LTPL-European [1281].

M [750, 250, 605, 606, 258, 175, 221, 500, 1084, 938]. M1 [602]. maat [1153]. Machine [1023, 119, 223, 688, 186, 929, 662, 45, 545, 546, 799, 768, 888]. Machine-Code [186]. machine-independent [888]. machine-like [45]. Machinery [605]. machines [788, 1137]. made [1008]. Maechanik [405]. magnetic [30]. Mailloux [500]. Maintenance [981, 1217]. make [105]. Making [120, 1163, 1189, 562, 273, 145, 40]. Malcolme [486]. Malcolme-Lawes [486]. management [826]. mandy [1069, 581]. manipulating [187]. Manipulation [5, 681, 977, 996, 1068, 64, 887, 107, 698]. Manitoba [1280]. Mantissenlänge [896]. Manual [756, 582, 648, 699, 323]. Manuel [744]. Many [730]. map [566]. map-folding [566]. mapping [1036]. March [1288, 1285]. Mark [1115, 1144]. market [1203]. Marriage [570]. Martin [707, 708]. MARY [642]. Maschinenintervallarithmetik [336]. masses [683, 1065]. matching [857]. mate [568]. Mathematical [1296, 608, 609, 1273, 1293, 1058, 1184, 607, 347, 355, 561, 746, 1145, 541]. Mathematics [1040, 1041, 1059, 1075, 1117, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1273, 1130, 1233, 1175]. Mathematik [605, 606, 405]. Mathematisch [1169]. mathematischen [528, 405, 307]. Matrices [617, 924, 531, 184, 185]. Matrix [1063, 296, 692, 1039, 779, 267, 67, 278, 779]. May [746]. MC [541, 997]. MC-25 [541]. McCarthy [605, 606, 258]. McCracken [161, 127, 128, 193]. McCreight [790]. McGettrick [937]. MDN [258]. means [134, 733, 698]. measure [1153]. Measurements [460, 706]. Mechanical [1243]. Mechanism [1088]. Media [1217]. Meek [970, 971]. Meetkunde [1156, 1197]. member [919]. mémoire [349]. Memoriam [1103]. memories [1137]. Memory [1243, 770, 111, 454, 706, 349, 571, 1036, 963]. Merge [602, 709]. Metalanguage [203, 954]. Metalinguistic [198]. Method [153, 610, 168, 169, 421, 755, 762, 176, 408, 180, 511, 637, 267, 1111, 145, 754, 730]. methode [1111]. Methoden [405, 307]. Methodist [1279]. Methods [608, 609, 1290, 1239, 269, 523, 589, 1177, 836, 838, 670, 165, 355, 561, 75, 698]. Metodika [985]. Meulen [705, 966]. Meyer [790]. MG [1111]. MG/eindige [1111]. MG/eindige-volume [1111]. Michael [1100]. Micro [959]. Microprogram [893]. might [884]. might-have-beens [884]. MIMD [1129]. Min [346]. mini [745]. Minimal [671]. Minimum [220]. MINIX [1224]. Minty [379, 379]. Mises [773]. missing [595]. mission [27]. Mitarb [258]. Mitarbeit [991]. Mixed [754]. MLISP [344]. mod [1147]. Mode [400, 577, 595, 950]. Model [1093, 628, 743, 1108, 1183]. Modeling [1177]. Modelle [961]. modelling [827]. Models [1106, 1129, 988, 461, 517]. modes [380, 947, 418, 426, 578]. modest [780]. Modified [786, 903, 379, 176, 645, 122]. modifizierten [379]. Modula [1227]. Modula-2 [1227]. Modular [1224, 1088]. modulo [425]. moments [645]. momentum [683, 1065]. momentum-space [683, 1065]. Monday [1285]. money [1233]. Monitoring [904]. Montague [863, 945]. Monte [35, 673]. Moody [1145]. Moore [379]. motives [12]. move [568]. moves [568]. Muelen [594]. Multi [1262, 1082, 566, 1137, 891]. Multi-dimensional [566, 1137]. Multi-sequential [1082]. Multi-token [1262]. Multidimensional [673]. multigrid [1030, 1121]. Multilanguage [543, 1114]. Multipass [560]. Multiple [622, 734, 810]. multiplication [220, 67]. Multiplicators [451]. Multiply [813]. multiprocessing [263]. multiprogramming [339]. Multithreaded [1219]. München [961]. Munich [1278, 740, 1282]. Munkres [531]. MUTANT [430]. Mutual [664, 704]. Myers [970, 971].

N [967, 1026, 352, 583, 593]. nach [379]. NAG [1076]. Näherungswerten [510]. Name [241]. names [411, 505]. Naming [1015]. narrative [377]. National [526]. Nationale [346]. Natural [552, 678, 721, 726, 803, 519]. Naur [605, 606, 258, 9, 111]. NEAT [574, 575]. necessary [461]. negative [670]. negativer [670]. NELIAC [15, 131]. nelineinykh [985]. Nested [610, 1046, 983]. nesting [300]. nestings [949]. Netherlands [1293, 1221]. Nets [1259]. networks [584, 697]. Netzwerke [697]. Netzwerken [584]. Neue [311]. Neural [1259]. Newton [1132]. nicht [779]. Nicot [254]. nieuw [1197]. Niklaus [1195]. No [111, 705, 750, 966, 1026, 937, 723, 740, 80, 81, 109, 248, 114, 83, 606, 290, 291, 341, 342, 343, 124, 292, 168, 259, 401, 621, 176, 634, 635, 184, 185, 270, 317, 239, 245]. nodes [759]. Non [1143, 115, 304, 1183, 590]. non-determinism [1183]. non-existence [115]. Non-Interference [1143]. non-recursive [590]. non-standard [304]. nondeterminism [1009]. nonexistence [94]. nonlinear [1121]. Nonsymmetric [929]. Normal [303, 679, 229, 292, 356, 332, 160, 302, 303]. North [705, 966, 740]. North-Holland [705, 740]. Northwest [66]. Norwegian [87]. Norwich [1285]. Notation [226, 227, 892, 446]. Note [115, 795, 169, 170, 1190, 191, 777, 1214, 158, 856, 132, 1216, 416, 694, 925, 656, 821]. Notes [155, 480, 162, 558, 625, 737, 277, 1018, 1053]. nothing [1186]. Notices [1176]. notion [998]. November [1279]. Novye [1061]. nuclear [581, 1069]. nucleus [713]. Nullstellen [598]. Number [1256, 981, 375, 1157, 98, 767, 232]. Numbering [806]. Numbers [451, 929]. Numerical [1289, 355, 561, 512, 269, 1240, 276, 12, 837, 799, 943, 64, 1054, 239, 735, 736, 986, 608, 609, 165]. Numerics [1245]. Numerik [1245]. Numérique [438, 346]. Numerische [605, 606, 405, 307]. Numerischen [751]. nyelvu [868].

O [256, 294, 258, 314, 1183, 246]. Oak [1289]. ob [444, 520]. Oberon [1227]. Object [206, 1252, 1174, 389, 1149]. Object-Oriented [1174]. objects [1033, 1132]. obratnykh [985]. observations [1051]. Obtain [935]. occasion [541, 1293]. occurrences [394]. occurring [29]. October [801, 1293]. odd [1104]. odd-even [1104]. off [62]. Official [110]. OI [1027]. OI-hierarchies [1027]. Oklahoma [747, 851]. old [1197]. Oldenburg [961]. O’Leary [1100]. om [87]. omission [597]. OMNITAB [619]. onbewezen [1200]. One [210, 860, 695, 932, 1009, 514, 821, 842]. One-Dimensional [695]. one-pass [514, 821]. op [1193]. opaque [675]. open [38]. Opening [1022]. Operand [233]. Operating [90, 249, 252, 253]. operational [1009]. Operations [924, 761, 1235, 513, 70]. Operator [95, 123, 400]. Operators [592, 952]. Optimierung [405, 307, 754]. optimised [823]. Optimization [608, 609, 543, 1020, 355, 561, 770, 1135, 639]. Optimized [792, 1056, 1057]. option [1203]. Optional [1028]. ORACLE [30]. Order [783, 801, 1178, 755, 944, 1231, 1192]. Ordered [695]. Ordering [450]. Ordinary [960]. Organisation [349]. organisatorische [1041]. Organization [879, 201, 395, 506, 349]. organizational [1041]. organized [746, 1074]. orientated [864]. Oriented [600, 543, 1174, 1113, 1128, 43, 48, 872, 735, 736]. Origins [1257]. Orlando [1298]. Orsay [882]. Orthogonal [240, 1134, 1172]. orthogonality [1173]. Osnovnye [653]. other [434]. oud [1197]. out-of-kilter-algorithm [697]. Out-Of-Kilter-Algorithmus [697]. Outer [498, 374]. Output [338, 166, 178, 121, 130, 1051, 1125, 656, 449]. Outputs [382]. Outward [1102]. Overa [685]. overview [1099].

P [111, 750, 991, 605, 606, 608, 609, 258, 405, 185]. P. [111, 991]. Package [1063, 462, 972, 803]. pages [705, 750, 966, 1026, 346, 937, 723, 740]. pain [77]. Panel [527, 548]. paper [105]. paperback [1026, 937, 705]. Papers [117, 1215, 1292, 1300, 1302, 162, 277]. Parallel [199, 1074, 1180, 1077, 1078, 969, 1177, 843, 263, 694]. Parameters [933, 955, 1028, 944, 501, 1123]. Parametricity [1181, 1168]. Parasol [1162]. Parenthesis [533, 467]. Paris [882]. Paris-XI [882]. parity [834]. Parlett [162, 277]. Parser [689, 447]. Parser-Generating [689]. Parsing [1223, 212, 1043, 859, 409, 1220]. Part [528, 512, 283, 1177, 1116, 842, 226, 227, 135, 136, 284, 285, 1126]. Partial [450, 755, 576, 698]. particles [683, 1065]. particular [429]. Partition [425].

Partitions [813]. Partsch [991, 991]. PASCAL [909, 785, 782, 783, 1079, 1105, 1120, 801, 1097, 1017, 1018, 1053, 522, 658, 703, 785, 842, 999, 1046, 1164, 1094, 1037, 923, 1102, 657]. PASCAL-based [1017]. pass [514, 821]. passive [1199]. path [60]. Pattern [668, 686, 857]. patterns [829]. PDP11 [825]. PDP11/ICL1900 [825]. Pearl [784, 693, 1281]. Peck [500, 740]. pedagogical [814]. Pegasus [191]. Penultimate [358, 369]. PEP [1070]. Pepper [991]. Performance [1177, 235]. Period [801]. Periodic [759, 942, 666]. Perlis [605, 606, 258]. Permutational [691]. Permutations [806]. Persistent [1042, 1024, 1118]. personal [919, 1169]. perspectief [1175]. perspective [1175]. Perspectives [1174]. Peter [970, 971, 9]. phase [918]. Philip [1084]. Phrase [555, 620, 115, 94]. physical [89]. Physicist [1208]. physics [1115, 1144]. Piece [1249]. pieces [849]. Pierre [1287]. Pierre-de-Chartreuse [1287]. pivoting [74]. PL [970, 971, 381, 784, 900, 479, 1079, 1105, 1120, 1094, 633, 915, 916, 917, 951, 1097, 921, 693, 1037, 737]. PL-I [951]. PL/1 [970, 971, 784, 1079, 1105, 1120, 1094, 1097]. PL/I [381, 900, 479, 633, 915, 916, 917, 921, 693, 1037, 737]. PL360 [567, 630, 631]. PLAIN [1021]. Planarity [795, 722]. planning [27, 611]. pleasure [77]. Plotter [538, 701, 675]. plotting [796]. plus [1059]. Point [542, 688, 672, 98, 1235]. points [789]. Policy [381]. Polish [677]. Politics [1260]. Polygamma [771, 440]. polymorphic [1060, 1204]. Polynomial [153, 353]. Polynomials [510, 317, 1010]. Polynomwurzeln [510]. Ponyaiya [653]. POPL [1297, 1298, 1302]. Portability [537, 1289]. portable [871]. PORTAL [1017]. positiv [779]. Positive [153, 779, 557]. Possibilities [206]. Possible [273, 1226, 1142]. possible-world [1226, 1142]. Poue [363]. power [861, 907]. powerful [930]. pp [248, 605, 606, 612, 169, 634, 635]. Präcompiler [956]. Practical [380, 1187, 604, 953, 811, 661, 1008, 408]. practice [1175]. praktijk [1175]. Praktijkcursus [348]. praktischen [487]. Precedence [95, 123]. precisie [372]. Precision [542, 771, 957, 440, 672, 1263, 372, 986]. predicate [938]. Predicates [1096]. Preis [258]. Preliminary [9, 480, 611, 415, 1]. Prelude [846]. prepared [1109, 1282, 1287]. preprint [897, 861, 862, 863, 945]. Preprints [1299]. preprocessor [304]. prepublication [790]. Presentations [161]. presented [1292, 1300, 1302, 1145]. President [55]. Press [723]. preview [377]. Price [705, 750, 966, 1026, 937, 723, 740]. pricing [1203]. prime [976]. primer [91, 524]. primes [134]. Principles [1297, 1298, 1300, 1302, 469, 993, 639]. probability [1115, 1144]. Problem [902, 48, 616, 864, 570, 12, 531, 755, 1091, 830, 619, 947, 687, 509]. Problem-oriented [48]. Problem-Solving [902, 830]. Problems [671, 1063, 126, 549, 1031, 725, 228, 781, 945, 137, 873, 165]. Procedural [1132, 944]. Procedure [162, 296, 688, 366, 432, 277, 165, 949, 65, 433, 303, 229, 356]. Procedures [379, 338, 612, 47, 350, 166, 212, 975, 259, 552, 678, 721, 726, 570, 1123, 955, 512, 143, 365, 273, 108, 1116, 670, 170, 491, 56, 622, 865, 501, 503, 875, 429, 142, 1138, 372, 445, 148, 1173, 656, 197, 1126, 438, 346]. Proceedings [1297, 1298, 1058, 1130, 1145, 1184, 1221, 1295, 1273, 1293, 1177, 1284, 1278, 740, 1288, 1292, 1277, 1130, 1280, 1285, 1282, 1287]. Process [1201, 734, 792, 650]. Processable [969]. processes [25, 49, 52]. Processing [199, 179, 270, 200, 462, 1140, 843, 875]. processor [752, 45, 677, 274]. processors [850, 408]. PROCOL [784, 693]. Producing [823]. Product [498, 372]. products [507, 729]. prof [1293]. Professor [1100, 1022]. profil [1229]. profile [1229]. profiles [1085]. profiling [828]. Program [1155, 199, 1025, 537, 714, 1076, 206, 980, 870, 808, 1048, 960, 989, 991, 333, 669, 840, 1198, 683, 1065, 910, 807, 1067, 729, 580, 1085, 654]. Programacão [574]. Programmacão [575]. Programme [961, 849, 553, 271, 272, 702]. Programmed [716, 596, 990]. Programmen [405, 307, 271, 272, 751]. Programmentwicklung [991]. programmer [649]. programmeren [13, 483]. programmers [874]. Programmieren [728]. Programmiersprache [424]. Programmiersprachen [494, 909]. Programmierung [99, 872, 627]. Programmierungssprachen [785]. Programmierungstechnik [99]. Programming [1291, 1299, 1300, 1302, 1246, 1072, 80, 23, 81, 109, 783, 784, 248, 1042, 1090, 1158, 1303, 965, 712, 754, 83, 935, 480, 830, 902, 1062, 119, 1188, 612, 543, 254, 124, 292, 801, 168, 169, 974, 1002, 259, 401, 1224, 221, 621, 1064, 1162, 176, 1141, 179, 567, 630, 634, 635, 1095, 133, 183, 184, 185, 1124, 1236, 1260, 425, 1170, 1191, 644, 1211, 145, 1085, 959, 1018, 1053, 270, 317, 734, 1251, 239, 279, 522, 658, 703, 1086, 245, 924, 1297, 1298, 743, 745, 22, 110, 154, 785, 1186, 286, 114, 899, 934, 1257, 993]. programming [478, 204, 604, 968, 13, 483, 715, 758, 973, 48, 905, 99, 494, 174, 498, 861, 907, 979, 624, 868, 1082, 769, 631, 633, 1011, 1150, 1190, 7, 872, 874, 429, 1110, 921, 693, 1281, 1286, 267, 430, 646, 1017, 1070, 374, 325, 1172, 708, 486, 128, 728, 316, 707, 967, 161, 345, 127, 171, 193, 1026]. Programming-ALGOL [486]. programmirovaniëiìa [1061]. programmy [985]. programok [868]. programozasi [868]. Programs [1234, 824, 1187, 207, 208, 904, 1107, 946, 1031, 912, 235, 271, 272, 1071, 706, 29, 992, 828, 752, 711, 89, 941, 720, 853, 553, 868, 409, 355, 561, 947, 418, 310, 419, 265, 639, 1274, 580, 1038, 581, 1069, 517, 106, 1056, 1057, 849, 702, 608, 609]. Project [1250, 328, 1098, 1260, 1102]. project-based [1098]. Projections [668, 686]. Projet [328]. prolog [1261]. Proof [1219, 115, 1000, 1235]. Proofs [350]. proper [862, 908]. Properties [784, 459, 831, 832, 398, 693, 1281, 1286, 517]. property [458, 1036]. Proposal [451, 178, 1083, 373, 244, 255, 293, 297, 260, 238, 780, 275, 318]. Proposals [234, 71]. Proposed [2, 6]. prover [713]. Prozeduren [379, 670, 289]. Prozedurschachtelungen [949]. PS [1024, 1108, 1125, 1136]. PS-Algol [1108, 1125, 1024, 1136]. Pseudo [451]. Pseudo-Random [451]. PSG [1090]. publication [195, 149]. Publications [161]. Published [346, 961]. Purdue [1286]. purpose [197]. purposes [983].

QMIPS [1177]. Quadratic [397, 1179]. Quadrature [673, 906, 637, 645]. Quantification [1112, 945]. Quantifying [853]. Quantiles [493, 496, 1004, 936]. Quantitative [1177]. Quarterly [182, 232]. queens [1091]. Quelques [473]. quindiagonal [666]. Quintic [678, 721].

R [750, 707, 708, 250, 796, 171, 175, 855, 621, 1138, 961, 663, 828, 539, 540, 904, 524, 659, 660, 739]. R. [487, 439]. R2 [231, 431]. R65 [218]. R65-26 [218]. R67 [320]. R67-15 [320]. R68 [373]. R68-24 [373]. Ractliffe [171]. Radar [891]. Random [451, 1137]. Range [47]. ranges [799, 354]. Rapport [463, 464, 603, 328]. rather [619]. Ratio [397, 312]. ratios [164]. Raulefs [970, 971]. Rayleigh [353]. Re [28]. reactions [581, 1069]. reader [1109]. Real [692, 273, 1157, 598, 392, 184, 185, 693, 1281, 1017, 278, 36, 303]. real-time [693, 1281, 1017]. Realisierung [896, 336, 324]. Realization [929, 148]. Realzeitrechner [961]. Realzeitrechner-Modelle [961]. Rechenanlagen [99]. Rechenmaschinen [311]. Rechenmaschinen-Systeme [311]. Recognition [532, 354]. Recognizers [213]. recollection [1169]. Recommendation [351]. Recommendations [3]. Reconstruction [668, 686]. Record [748, 749, 219, 1300, 1302]. Record-Classes [748, 749]. rectangular [531]. Recursion [933, 664, 704, 789, 872]. recursion-oriented [872]. Recursive [1043, 49, 228, 143, 777, 25, 494, 56, 142, 925, 590, 591]. Recursively [126, 777, 886, 315, 925]. REDUCE [760]. redundancy [106]. reele [1157]. Reeves [221]. Ref [621]. Reference [756, 231, 582, 648, 453, 761, 431]. referencing [391, 862, 908]. Refinement [286]. refinements [334]. Refining [206]. reflections [1254]. Regular [481, 407]. rekenen [1193, 445]. Rekeninstituut [797]. rekursionsorientierte [872]. rekursive [494]. Related [247, 97]. Relation [353, 249, 693, 1281]. Relational [1168]. relations [600]. Relationships [988]. Relative [213]. Relays [1243]. Reliability [1129]. Reliable [970, 971, 970]. rellen [598]. Remaining [306]. Remark [29, 709, 481, 36, 612, 163, 488, 1003, 1004, 1081, 860, 724, 686, 356, 312, 313, 431, 433, 776, 74, 700, 701, 798, 936, 332, 407, 114]. Remarks [5, 303, 222, 436, 78, 880, 446, 647]. Remarques [364]. René [28]. repeatable [1028]. reply [76]. Report [9, 10, 11, 786, 603, 794, 4, 254, 256, 294, 801, 166, 167, 100, 351, 1107, 182, 232, 17, 103, 138, 1194, 1, 7, 314, 316, 244, 658, 703, 370, 442, 443, 444, 520, 588, 778, 817, 818, 1127, 328, 112, 113, 463, 464, 606, 839, 91, 611, 848, 358, 415, 236, 105, 238, 152, 367, 368, 369, 889, 926, 111, 605, 258, 623, 102, 652, 484]. Reports [436, 1144]. Representation [578, 1232, 156, 125, 400, 98, 676, 720, 848, 412, 144, 192, 194]. Representations [378, 498, 682, 767]. Reprint [1047]. Requirements [783, 801, 571, 1037]. requiring [1043]. research [758]. Resetting [1080]. resheniia [985]. resistance [333]. responses [110]. Restricted [684, 813, 664, 704]. restrictions [12, 719]. restructuring [980]. resultaten [1111]. results [12, 1111]. retirement [1293]. Revidierter [926]. Review [594, 705, 750, 966, 250, 967, 1026, 605, 606, 608, 609, 484, 486, 487, 346, 937, 127, 128, 171, 175, 221, 723, 500, 405, 1259, 512, 1084, 1138, 269, 438, 439, 193, 961, 740, 172, 173]. Reviews [345, 707, 708, 970, 971, 528]. Revised [381, 112, 113, 794, 567, 630, 103, 138, 236, 1194, 731, 658, 703, 778, 817, 818, 889, 331, 384, 966, 839, 676, 631, 606, 926]. revisited [642, 1019]. rewriting [1220]. Ricatti [477]. Riemann [230]. Right [819]. ring [462, 1151]. Ritchie [1258]. Roe [1111]. role [944, 1022, 963]. Rolle [963]. Romberg [169, 168]. ROME [146]. Roots [153, 598, 75]. Round [62]. Round-off [62]. Rounding [327]. Routh [996]. routine [63]. routines [453, 1030]. RS [1084, 1055]. Rt [953]. RTL [784, 693]. RTL/2 [693]. ruimten [885]. rule [1000]. Rules [1092, 840, 945]. Run [828, 904]. Run-time [828, 904]. Russia [1304]. Russian [198, 236, 1251, 653, 929, 444, 520]. Rutishauser [528, 605, 606, 258, 1245].

S [594, 705, 750, 528, 966, 258, 1061, 1034, 1035]. S13 [674]. S14 [340, 679, 680, 565, 440, 936]. S15 [798, 332, 396, 303, 229, 361]. S16 [392, 210]. S20 [181]. S22 [481, 407]. Safe [1118]. SAIL [613, 774, 812, 815, 699]. Samelson [250, 605, 606, 258, 175]. Sampling [673]. San [1297, 1302]. SATHE [964]. Satzstruktur [427, 428]. Satzstruktur-Grammatik [427, 428]. saving [810]. scale [646]. SCAMP [828]. Scan [189]. scanner [253]. scheduling [60]. schema [1111]. schemata [614]. Scheme [535, 1188, 1149, 200, 476, 58, 1104, 1188]. Schemes [1039, 1111]. schnellen [730]. Schrankenzahlen [375, 376]. Schultz [707, 708]. Schweizer [1245]. Science [1040, 1117, 716, 802, 1280, 1259, 1071, 276, 746, 1156, 989, 596, 990, 814, 1041, 1059, 1075, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 971, 970]. scientific [12]. Scientists [521]. Scope [532, 1092, 939, 940]. scopes [1046]. Scoping [1134, 1044]. Second [477, 982, 1259, 1178, 1221, 937, 975, 913, 1282, 1287, 1192, 99]. section [888]. sections [1179, 581, 1069]. securities [1203]. Segment [460, 824]. Segmentation [946, 235]. seiten [961]. Selected [1215]. Selecting [814, 1079, 1105, 1120, 1097]. selection [1144]. Semantic [955, 19, 928, 819, 958]. Semantics [2, 466, 499, 556, 954, 1201, 1196, 1072, 475, 534, 1140, 1226, 861, 907, 1009, 1122, 1142, 651]. seminar [1058, 1115, 1144, 1145, 1184]. SEN [1214, 1216]. sensitive [458]. sentence [427, 428]. sentence-structure [427, 428]. separate [948, 1123]. September [1287, 381]. Sequence [640, 641, 300]. Sequential [1234, 18, 1082, 8]. Sequentielle [8]. series [1074, 698]. Set [338, 273, 929, 315]. Sets [727]. Seven [1038]. Several [965]. Shaded [804, 805]. shallow [1228]. Shape [898]. Shaped [1208]. shapes [1179]. SHARE [3]. sharing [1196, 268]. Sharp [707, 708]. Shift [617, 353]. Short [162, 277]. Shorter [1148]. should [129, 509]. SID [537]. side [918, 920, 995, 1119]. side-effects [1119]. SIEVE [79]. SIGACT [1300, 1302]. Sign [753]. Significance [691]. SIGPLAN [1291, 1292, 1299, 1299, 1300, 1302, 1176, 905, 921]. SIGPLAN-SIGACT [1300, 1302]. similar [462]. Simple [743, 930, 545, 546]. simplex [176]. Simplifying [612, 259]. SIMULA [783, 251, 801, 729]. Simulation [863, 775, 251, 545, 546, 877, 891, 545, 546]. simultaneous [557]. Sine [84, 85]. Single [189, 799]. Single-Scan [189]. Singular [478, 296, 1263]. sistema [1061]. Size [460, 924, 976]. Sizes [824]. Skew [682]. SMALGOL [26]. SMALGOL-61 [26]. Small [529, 1052]. Smith [292]. Smoothing [726]. SNOBOL [829]. SOAP [580]. soda [1228]. Software [1073, 710, 1289, 1296, 970, 971, 1141, 419, 1209, 1071, 837, 943]. SOL [177]. solids [675]. Solution [247, 452, 80, 557, 228, 245, 666, 1177, 1091, 835, 544, 975, 856, 576, 698]. Solutions [286, 334, 1066, 165]. solver [836, 838]. Solving [902, 1063, 830, 165, 619, 571]. Some [450, 964, 459, 790, 537, 126, 549, 719, 211, 941, 563, 1051, 873, 642, 71, 880, 777, 446, 1041, 347, 1006, 417, 426, 643, 925, 986, 473, 398, 1111]. Soobshchenie [444, 520]. Sort [602, 709, 590, 591]. Sorties [382]. sorts [1038]. SoRuCom [1304]. Sound [72]. Source [870, 1220]. Source-to-Source [870]. Sources [1296]. South [1300]. Southern [1279]. Soviet [1304]. SP1 [1225]. space [611, 683, 1065, 810]. space-saving [810]. spaced [615]. spaces [885]. Sparse [1039]. Special [1239]. Specialization [1187]. specific [517]. Specification [203, 955, 1087]. speculative [1230]. speed [662, 654]. spelen [1233]. Spezifikation [1087]. Spherical [772]. Spline [708, 552, 678, 721, 759, 803, 707]. Splines [726, 803]. Spots [306, 448]. Sprache [111, 991, 258, 553, 623, 102, 314, 956, 652, 926, 872]. Sprachen [99]. Sprachkonzepte [872]. Springer [528]. Springer-Verlag [528]. Squared [301, 1081, 798, 361]. Squares [286, 296, 334, 836, 838, 165]. sravnenie [1034, 1035]. sredstva [1061]. St. [1287]. Stable [570]. stack [472]. Stage [353, 427, 428]. Stages [1102]. standard [848, 491, 304, 503, 846]. Standardfunktionen [491, 503]. standards [793]. standpoint [633]. Stanford [1244]. STATAL [1116, 1126, 1138]. State [1232, 747, 851, 265, 1196]. State-dependent [1232]. Statement [610, 809, 16, 1125, 821]. Statements [14, 499, 54, 556, 558, 625, 133]. Static [1140, 831, 832, 928, 1044]. Stationary [397]. Statistical [753, 772, 773, 691, 640, 641, 1041, 1116, 619, 1066, 1138, 1126]. Statistics [772, 941]. statistische [1041]. steady [1121]. Steger [1111]. step [903]. step-until [903]. Steve [1218]. Stichworten [99]. Stiefel [963]. Storage [826, 535, 395, 1032, 506, 1039, 380, 30, 389, 476, 58, 1007, 426, 68]. Storing [898]. Story [919]. Strathclyde [1288]. Stratified [673]. String [765, 359, 401, 857, 107, 677]. strings [145]. stromingen [1111]. strong [998]. strongly [932, 1118]. Structural [475, 212, 386]. Structure [82, 288, 474, 533, 610, 50, 1029, 677, 1273, 946, 467, 115, 339, 94, 48, 225, 61, 950, 1150, 427, 428, 1015, 1087, 593, 474, 912]. structure-based [1150]. Structured [1092, 968, 994, 758, 802, 980, 948, 874, 702, 741]. Structures [199, 620, 1058, 1145, 1184, 479, 938, 555, 1006, 1007, 770, 1094, 1133, 893, 885]. structuring [930]. Struktur [1087]. Student [183, 492, 493, 495, 496, 680, 1003, 1004, 936]. Students [716, 596, 990, 793]. Studiedag [1139]. studies [1284]. Study [712, 1107, 1221, 624, 1235, 1139]. Style [1247, 133, 576]. Styles [1246]. Sub [4]. Sub-Committee [4]. Subcommittee [256, 294]. Subject [397]. sublanguage [854, 866]. Sublanguages [548, 563, 983]. Subrata [1259]. subroutines [507]. Subset [1178, 842, 1192, 878, 167]. substrata [1228]. Successor [260, 244, 158, 238]. Suche [670]. suggested [145]. Suggestions [146, 217]. Suit [562]. suites [1269]. Summanden [730]. summary [905, 921]. Summation [688, 421, 511]. Summen [730]. Summierungsverfahren [421, 422, 511]. Sums [730]. super [846]. super-language [846]. Superimposed [434]. superposées [434]. Supplement [111, 152, 794, 839]. Support [1219]. Supports [1050]. surfaces [849]. Survey [1158, 618, 669, 735, 736]. Susan [1084]. suspicions [1200]. Sven [1218]. Swiss [1245]. syllabus [1175, 1115, 1144]. Symbol [532, 5, 201, 465, 87, 939, 940, 64, 107]. symbolbruken [87]. Symbolic [995, 1067, 639, 576, 807]. symbols [214, 215, 192]. Symmetric [616, 617, 498, 682, 692, 779, 544, 184, 185]. symmetrical [643]. Symmetrie [1237]. symmetrischen [779]. Symmetry [1237]. Symposia [1273, 1240]. Symposium [1292, 1300, 1302, 1279, 1295, 1293, 1298, 1130, 541, 1284]. Synchronous [264]. Syntactic [791, 95, 123, 131, 19, 892]. syntactical [73]. Syntax [2, 526, 203, 57, 1047, 222, 979, 409, 577, 458, 459, 530, 468, 601, 391, 202, 347, 551, 497, 180, 362, 731, 315, 447, 374, 449]. syntax-checking [180]. syntax-directed [530]. Synthex [69]. systeemtheorie [1146]. System [1219, 894, 1090, 896, 390, 1077, 1078, 543, 547, 760, 974, 1002, 1224, 855, 977, 806, 1032, 689, 186, 235, 360, 268, 586, 323, 1114, 924, 987, 249, 755, 252, 253, 339, 611, 715, 295, 130, 174, 619, 1070, 324]. Systeme [311, 1087]. systèmes [434]. Systems [452, 935, 264, 644, 734, 1217, 376, 1129, 1146, 454, 666, 1177, 544, 856, 1230, 571, 434, 735, 736, 1021, 1055, 1087, 245, 1286, 1084]. szamitogepes [868].

T [512]. T. [316]. Tabellen [961]. Table [532, 252, 465, 834, 265, 877]. Tables [335, 1005, 560, 201, 939, 940, 978, 1036, 279]. TACPOL [783, 801]. tailed [643]. Talk [1210]. tape [30]. Tarjan [795]. Tauber [1101]. Tauberian [1101]. Taylor [698]. Tchebycheff [46]. teaching [983, 72]. team [919]. Tears [629, 833]. Technical [981, 1128, 162, 99, 277]. Technique [931, 542, 560, 810, 64]. Techniques [211, 189, 54]. Technology [716, 1304, 596, 990, 99]. Tel [1233]. TELCOMP [419]. teleprinter [125, 144]. Telescope [33, 34]. temperatures [333]. term [544]. Termination [1031, 1000, 947]. Terms [198, 730, 99, 733]. tervezese [868]. Test [753, 853, 1269]. testbed [943]. Testing [1252, 722, 1066, 1141, 691, 241, 820, 834, 1198]. Tests [162]. Texas [1279]. Text [781]. Th [311, 75]. Their [480, 765, 1240, 350, 960, 1153]. theorem [789, 713]. theorem-prover [713]. Theoretic [1180, 1015]. theorie [1145]. theories [1058, 1184]. Theory [1219, 104, 1071, 1146, 790, 1027, 717, 1101]. thermometers [333]. Think [1249]. Third [1231, 1304, 1177, 1284]. Third-order [1231]. those [960]. thoughts [986]. threaded [45]. Three [353, 570, 544, 212]. Three-Stage [353]. Thursday [1285]. Time [1029, 571, 268, 828, 904, 693, 1281, 1017]. Time-sharing [268]. TINMAN [783, 801]. Titan [1151]. toeval [1193]. token [1262]. too [1255]. Tool [1219, 27, 729]. Tools [1073, 814]. Torgil [345]. Torino [1177]. TORRIX [924, 885]. tot [1156]. total [581, 1069]. tracing [883]. Tracking [891]. Traduction [504, 328]. Tranal [807, 1067]. Transform [365, 366, 432, 433]. Transformation [286, 870, 1220]. Transformations [1048, 334]. Transitivity [727]. translate [419]. Translating [1107]. Translation [1073, 528, 298, 564, 189, 18, 792, 40, 49, 807, 1067, 410, 873, 1274, 8, 515, 328, 504]. translations [89, 581, 1069]. Translator [117, 206, 999, 50, 100, 321, 30, 40, 41, 118, 120, 252, 403, 309, 188]. translëiìator [1061]. transonic [1111]. transportable [852]. Transportation [720, 845, 687]. transput [453, 900, 888, 962]. transsone [1111]. treating [482]. Treatment [1180, 1113, 1128, 862, 908]. Tree [690, 1027, 590, 591]. trees [456, 600, 843]. tribute [1293]. Tridiagonal [617, 692, 856]. trigonometric [787]. trigonometrischen [787]. Triplex [390, 800, 421, 422, 511, 424, 732, 956, 323, 326, 797, 406, 311, 421, 423, 511, 322, 375]. Triplex-ALGOL [511, 390, 421, 422, 424, 732, 323, 511, 797, 406, 421, 423, 375]. Triplex-ALGOL-60 [326]. Triplex-Algol-Z23 [311]. Triplex-Langzahl-Arithmetik [800]. Triplex-Zahlen [322]. Triplexrechnungen [289]. Trouble [306, 448]. Turanski [111, 111]. Turing [545, 546]. Turingmaschine [545, 546]. Tutorial [404, 816, 881, 105, 815]. twelve [1019]. Twentieth [1300]. Two [714, 117, 671, 1106, 97, 502, 955, 755, 670, 305, 427, 428, 1104]. two-dimensional [1104]. Two-Level [955]. two-stage [427, 428]. Type [932, 1029, 628, 831, 832, 1028, 877, 581, 1069]. typechecking [1060]. typed [932, 1118]. Types [1106, 233, 986]. Typesetting [319]. typing [998]. Tzschach [608, 609, 405].

U [528, 258]. Übers [258]. Übersetzer [41]. ubersichtlicher [909]. Übungen [439]. UDC [621]. uit [1233]. uitg [483]. UK [1288, 1089]. Umwandlung [271, 272]. Unambiguous [224]. Uncertainty [640]. UNCOL [32]. undergraduate [802]. Understanding [643, 241]. Undirected [667]. unequally [615]. Unfolding [1262]. Unified [1241]. unintentional [597]. Union [1304]. unions [831, 832]. Univac [732, 800]. Univac-Algol-60-Compiler [800]. Universal [43, 543, 1250, 545, 546]. Universality [1247]. universelle [545, 546]. University [1279, 1295, 1280, 1285, 882, 746, 747, 1074, 1244, 851]. Unlimited [273]. Unnecessary [841, 892]. unproven [1200]. Unraveling [1236]. Unsolvable [126]. Unterprogrammen [896]. Untersuchungen [1087]. until [903, 1285]. Unwinding [1256]. Updating [478]. upper [635]. USA [1291]. Usage [1048]. Use [30, 82, 288, 802, 618, 143, 456, 1043, 462, 249, 87, 45, 98, 170, 101, 1274, 142, 105, 877, 145, 741]. Used [211]. User [1276, 699]. Users [750, 1084, 659, 739, 660, 1055]. Using [186, 955, 875, 1182, 664, 704, 1043, 828, 347, 125, 675, 803, 619, 1108, 770, 576, 144, 646, 1085, 887, 891, 1087, 698]. USSR [1160, 1270, 1251]. ustoichivogo [985]. Utrecht [1074].

Vakantiecursus [1059, 1075, 1117, 1131, 1146, 1157, 1179, 1185, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1233, 1237]. Validation [690]. validity [16]. valuation [517]. Value [478, 296, 241, 755, 1263]. valued [670]. Values [753, 296, 397, 510, 810, 1255, 887]. Variable [933, 353, 1048, 924, 672, 557]. Variable-Shift [353]. Variables [1181, 1143, 684, 1122, 1168]. Variants [526]. variatierekening [1075]. variations [1075]. varying [1028]. Vauquois [605, 606, 258]. vector [992, 239, 197, 780, 216]. Vectorcardiographic [96]. Vectorization [1104]. Vectors [924]. verbess [99]. Verfahren [754, 670, 730]. Verifiable [1088]. Verification [1234, 858, 1141, 995]. Verkettungsmatrix [627]. Verlag [258, 961, 528]. vermoedens [1200]. Verschwörung [1238, 1242]. Version [756, 1205]. Versus [783, 1247, 859]. Very [518]. vi [258]. vielen [730]. view [958]. Virtual [1076, 454, 706]. Visionaries [1100]. Visual [910]. Vol [528, 605, 606]. volume [1111, 1273, 924]. voor [1041, 1059, 1075, 1117, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 372, 445, 885]. vormen [1179]. vs [801]. VW [1001]. VW-grammar [1001]. VWF [446].

W [707, 708, 605, 606, 258, 512, 329, 330, 331, 383, 384, 599, 719, 981, 877, 648, 337]. W-grammar [337]. W-grammars [719]. W.G [256, 294]. Wählbarer [896]. Walke [961]. Walker [1026, 967]. walkthrough [718]. Warehouses [1217]. Warming [1111]. Was [1249]. Wave [481, 299, 700, 407]. Wegstein [605, 606, 258]. Weight [378]. Wells [221]. West [1282]. WG [1109]. Where [1254]. Which [935, 779, 580]. White [568]. Who [1208]. Wichmann [723]. Wiener [1101]. Wijngaaden [500]. Wijngaarden [605, 606, 1293, 1103, 457, 484, 636, 1012, 1169, 315, 886, 1022]. Wijngarden [258]. William [111, 111]. winnings [1233]. Winnipeg [1280]. winst [1233]. Wirth [1195]. Wiskunde [1175, 1041, 1059, 1075, 1117, 1131, 1146, 1157, 1193, 1197, 1200, 1206, 1212, 1222, 1229, 1130, 1233]. Wissenschaften [528]. Within [235, 273, 186]. without [833]. Woodger [605, 606, 258]. Woodward [750, 1084]. Wooldridge [171]. Word [39]. Words [1195, 99]. work [12, 1245]. worked [1066]. Working [540, 1278, 1282, 1287, 371, 472, 740]. Workshop [1289, 971, 1286, 1177, 970]. world [1226, 1142]. Wörterbuch [99]. Wörterverzeichnis [99]. Writing [710, 763, 764, 865, 868]. wrong [989].

X [528, 41, 256, 294, 303, 41, 305, 417, 642, 275]. X1 [1213, 1214, 117, 118, 1216, 40]. X8 [311, 289, 1235]. XI [882]. XII [528, 1273].

Yazyk [653]. yazyke [444, 520]. Year [88, 1058, 1115]. Years [1160, 1270, 1144, 1261, 1019]. yezik [585]. York [528].

Z [686, 668, 216]. Z23 [311]. zadach [985]. Zahlen [322]. Zehnder [608, 609, 405]. Zeichnen [849]. Zero [536]. Zeros [353, 317, 960, 36]. Zeta [230]. Zielke [314]. Zonneveld [1213, 1214, 1216]. zum [487, 849, 572]. Zur [111, 896, 1245, 987, 452, 598, 754, 670, 730, 787, 949, 427, 428, 1087]. Zurich [2, 12]. zwei [670, 311]. zweistufigen [427, 428]. Zyklen [670].

References

Perlis:1958:PRI

[1] Alan J. Perlis and K. Samelson. Preliminary report — International Algebraic Language. Communications of the ACM, 1(12):8–22, December 1958. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Backus:1959:SSP

[2] J. W. Backus. The syntax and semantics of the proposed International Algebraic Language of the Zurich ACM-GAMM Conference. In ????, editor, Information Processing: Proceedings of the International Conference on Information Processing, pages 125–131. R. Oldenbourg, München, Germany, 1959.

Bratman:1959:RSA

[3] Harvy Bratman, Julien Green, John Stockman, and Albert R. Watson. Recommendations of the SHARE ALGOL Committee. Communications of the ACM, 2(10):25–26, October 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1959:ASC

[4] Edsger W. Dijkstra, W. Heise, Alan J. Perlis, and K. Samelson. ALGOL sub-committee report — extensions. Communications of the ACM, 2(9): 24–??, ???? 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Green:1959:RAS

[5] Julien Green. Remarks on ALGOL and symbol manipulation. Communications of the ACM, 2(9):25–27, September 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irons:1959:PIA

[6] Edgar T. Irons and F. S. Acton. A proposed interpretation in ALGOL. Communications of the ACM, 2(12):14–15, December 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Perlis:1959:RAL

[7] A. J. Perlis and K. Samelson. Report on the algorithmic language AL-GOL by the ACM committee on programming languages and the GAMM committee on programming. Numerische Mathematik, 1:41–60, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Samelson:1959:SFG

[8] K. Samelson and F. L. Bauer. Sequentielle Formelübersetzung. (German) [Sequential formula translation]. Elektronische Rechenanlagen, 1 (??):176–182, 1959. CODEN ELRAA4. ISSN 0013-5720. See also [18].

Arden:1960:LPR

[9] B. Arden, B. Galler, and R. Graham. Letters: Preliminary report (February 4, 1960) of Peter Naur. Communications of the ACM, 3(6):A13, June 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [17].

Backus:1960:RALa

[10] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Numerische Mathematik, 2(1):106–136, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL https://link.springer.com/article/10.1007/BF01386216.

Backus:1960:RALb

[11] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Acta Polytechnica Scandinavica: Mathematics and Computing Machinery Series, (5):284–??, ???? 1960. CODEN APSMAT. ISSN 0355-2713.

Bauer:1960:PCL

[12] F. L. Bauer and K. Samelson. The problem of a common language, especially for scientific numerical work (motives, restrictions, aims and results of the Zurich Conference on ALGOL). In Anonymous, editor, Information Processing 1959: Proceedings of the International Conference on Information Processing, UNESCO, Paris, 15–20 June 1959, pages 120–125. UNESCO, Paris, France, 1960.

Dijkstra:1960:CPA

[13] Edsger W. Dijkstra. Cursus programmeren in ALGOL 60 (Dutch) [Course in programming in ALGOL 60]. Report ??, Mathematisch Centrum, Amsterdam, Amsterdam, The Netherlands, 1960. ?? pp.

Floyd:1960:ADA

[14] Robert W. Floyd. An algorithm defining ALGOL assignment statements. Communications of the ACM, 3(3):170–171, March 1960. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huskey:1960:NDA

[15] Harry D. Huskey, M. H. Halstead, and R. McArthur. NELIAC—a dialect of ALGOL. Communications of the ACM, 3(8):463–468, August 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McIsaac:1960:CAS

[16] Paul McIsaac. Combining ALGOL statement analysis with validity checking. Communications of the ACM, 3(7):418–419, July 1960. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:1960:RALa

[17] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the ACM, 3(5):299–314, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [9]. Dedicated to the Memory of William Turanski [American group member who was killed by an automobile just prior to the January 1960 Conference].

Samelson:1960:SFT

[18] K. Samelson and Friedrich L. Bauer. Sequential formula translation. Communications of the ACM, 3(2):76–83; addendum, 351, February 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1960:SSA

[19] Joseph W. Smith. Syntactic and semantic augments to ALGOL. Communications of the ACM, 3(4):211–213, April 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegstein:1960:AA

[20] J. H. Wegstein. Algorithms: Announcement. Communications of the ACM, 3(2):73, February 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woodger:1960:IA

[21] M. Woodger. An introduction to Algol 60. The Computer Journal, 3(2): 67–75, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/3/
2/67.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/030067.sgm.abs.html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/67.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/68.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/71.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/72.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/75.tif.

Anonymous:1961:APa

[22] Anonymous. ALGOL programming. Nordisk tidskrift for informationsbehandling, 1(1):64, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=1&issue=1&spage=64.

Anonymous:1961:APb

[23] Anonymous. ALGOL programming. Nordisk tidskrift for informationsbehandling, 1(4):290, December 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=1&issue=4&spage=290.

Arden:1961:LEC

[24] Bruce Arden, Bernard A. Galler, and Robert M. Graham. Letter to the Editor: criticisms of ALGOL 60. Communications of the ACM, 4(7): 309, July 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bachelor:1961:LEG

[25] Gilbert A. Bachelor. Letter to the Editor: on Grau’s recursive processes and on compiling expressions in ALGOL. Communications of the ACM, 4(6):252–253, June 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bachelor:1961:S

[26] G. A. Bachelor, J. R. H. Dempster, Donald E. Knuth, and J. Speroni. SMALGOL-61. Communications of the ACM, 4(11):499–502, November 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Batchelder:1961:CBM

[27] B. A. Batchelder. CAIN—a bombing mission planning tool. Communications of the ACM, 4(3):139–140, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bemer:1961:LER

[28] R. W. Bemer. Letter to the Editor: Re: René De La Briandais’ letter on FORTRAN. Communications of the ACM, 4(3):A12–A13, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [38].

Borsch-Supan:1961:RFO

[29] W. Börsch-Supan. Remark: On frequently occurring errors in Algol-60 programs. Communications of the ACM, 4(4):181, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bottenbruch:1961:UMT

[30] H. Bottenbruch. Use of magnetic tape for data storage in the ORACLE– ALGOL translator. Communications of the ACM, 4(1):15–19, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bratman:1961:AC

[31] Harvey Bratman. Automatic coding. Communications of the ACM, 4(3): 140, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bratman:1961:AFU

[32] Harvey Bratman. A alternate form of the “UNCOL diagram”. Communications of the ACM, 4(3):142, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brons:1961:ATa

[33] K. A. Brons. Algorithm 37: Telescope 1. Communications of the ACM, 4 (3):151, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brons:1961:ATb

[34] K. A. Brons. Algorithm 38: Telescope 2. Communications of the ACM, 4(3):151–152, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1961:MC

[35] Charles E. Clark. Monte Carlo. Communications of the ACM, 4(3):141, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Collinge:1961:RAR

[36] Robert M. Collinge. Remark on Algorithm 25: Real zeros of an arbitrary function. Communications of the ACM, 4(3):154, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conway:1961:LEA

[37] Melvin E. Conway. Letter to the Editor: ALGOL 60 comment. Communications of the ACM, 4(10):465, October 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeLaBriandais:1961:LEO

[38] Rene De La Briandais. Letter to the Editor: an open letter to Bob Bemer. Communications of the ACM, 4(3):A12, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See response [28].

DesJardins:1961:LEW

[39] Paul E. Des Jardins. Letters to the Editor: Word inversion. Communications of the ACM, 4(3):A13, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1961:ATA

[40] Edsger W. Dijkstra. Algol 60 translation: an Algol 60 translator for the x1 and Making a translator for Algol 60. Report 35, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1961. ?? pp. URL http://www.cs.utexas.edu/

users/EWD/MCReps/MR35.PDF.

Dijkstra:1961:AUX

[41] E. W. Dijkstra. Ein ALGOL-60-Übersetzer für die X 1. II. (German) [an ALGOL-60 translator for the X 1. II]. Math.-Tech.-Wirtschaft, 8: 115–119, 1961.

Dijkstra:1961:LED

[42] E. W. Dijkstra. Letter to the Editor: Defense of ALGOL 60. Communications of the ACM, 4(11):502–503, November 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dobrusky:1961:UCO

[43] W. B. Dobrusky and T. B. Steel. Universal computer-oriented language. Communications of the ACM, 4(3):138, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duncan:1961:LEC

[44] F. G. Duncan. Letter to the Editor: On the calculation of interest. Communications of the ACM, 4(3):A13, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evans:1961:UTL

[45] A. Evans, Jr., A. J. Perlis, and H. Van Zoeren. The use of threaded lists in constructing a combined ALGOL and machine-like assembly processor. Communications of the ACM, 4(1):36–41, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gianni:1961:AT

[46] A. J. Gianni. Algorithm 36: Tchebycheff. Communications of the ACM, 4(3):151, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gibb:1961:APRa

[47] A. Gibb. Algol procedures for range arithmetic. Technical Report ??, Applied Mathematics and Statistics Laboratories, Stanford University, Stanford, California, 1961. ?? pp.

Goldfinger:1961:POP

[48] Roy Goldfinger. Problem-oriented programming language structure. Communications of the ACM, 4(3):138, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grau:1961:RPA

[49] A. A. Grau. Recursive processes and ALGOL translation. Communications of the ACM, 4(1):10–15, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grau:1961:SAT

[50] A. A. Grau and P. Naur. The structure of an ALGOL translator. Report ORNL-3054, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1961. ?? pp.

Herndon:1961:ABF

[51] John R. Herndon. Algorithm 57: Ber or Bei function. Communications of the ACM, 4(4):181, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hormann:1961:CLH

[52] Aiko Hormann. Computer languages for heuristic processes. Communications of the ACM, 4(3):138, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huskey:1961:BCA

[53] H. D. Huskey and W. H. Wattenburg. A basic compiler for arithmetic expressions. Communications of the ACM, 4(1):3–9, January 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huskey:1961:CTB

[54] H. D. Huskey and W. H. Wattenburg. Compiling techniques for Boolean expressions and conditional statements in ALGOL 60. Communications of the ACM, 4(1):70–75, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huskey:1961:PAb

[55] Harry D. Huskey. From the President of ACM. Communications of the ACM, 4(3):136, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irons:1961:CIR

[56] E. T. Irons and W. Feurzeig. Comments on the implementation of recursive procedures and blocks in ALGOL 60. Communications of the ACM, 4(1):65–69, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irons:1961:SDC

[57] Edgar T. Irons. A syntax directed compiler for ALGOL 60. Communications of the ACM, 4(1):51–55, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jensen:1961:SAS

[58] J. Jensen, P. Mondrup, and P. Naur. A storage allocation scheme for ALGOL 60. Communications of the ACM, 4(10):441–445, October 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knuth:1961:AC

[59] Donald E. Knuth and Jack N. Merner. ALGOL 60 confidential. Communications of the ACM, 4(6):268–272, June 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [134].

Leavenworth:1961:ACP

[60] B. Leavenworth. Algorithm 40: Critical path scheduling. Communications of the ACM, 4(3):152–153, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levine:1961:LSC

[61] Syndey P. Levine. Logical structure of compilers. Communications of the ACM, 4(3):139, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piper:1961:LER

[62] C. A. Piper. Letter to the Editor: Round-off. Communications of the ACM, 4(3):A13, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rader:1961:CAB

[63] Patty Jane Rader. Certification of Algorithm 4: Bisection routine. Communications of the ACM, 4(3):153, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:1961:GTS

[64] Douglas T. Ross. A generalized technique for symbol manipulation and numerical calculation. Communications of the ACM, 4(3):147–150, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rutishauser:1961:IAP

[65] H. Rutishauser. Interference with an ALGOL-procedure. Annual Review in Automatic Programming, 2:67–75, January 1961. CODEN ARAPDX. ISSN 0066-4138 (print), 1878-545X (electronic).

Sanders:1961:LEN

[66] Norman Sanders. Letter to the Editor: Northwest Computing Association. Communications of the ACM, 4(3):A13, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sassouni:1961:ACC

[67] Papken Sassouni. Algorithm 39: Correlation coefficients with matrix multiplication. Communications of the ACM, 4(3):152, March 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sattley:1961:ASA

[68] Kirk Sattley. Allocation of storage for arrays in ALGOL 60. Communications of the ACM, 4(1):60–65, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simmons:1961:S

[69] R. F. Simmons. Synthex. Communications of the ACM, 4(3):140, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Strachey:1961:BO

[70] C. Strachey. Bitwise operations. Communications of the ACM, 4(3):146, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Strachey:1961:SPI

[71] Christopher Strachey and Maurce V. Wilkes. Some proposals for improving the efficiency of ALGOL 60. Communications of the ACM, 4(11): 488–491, November 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Swets:1961:SLA

[72] John A. Swets. Sound learning: an application of a computer to automated teaching. Communications of the ACM, 4(3):139, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taylor:1961:SCA

[73] Warren Taylor, Lloyd Turner, and Richard Waychoff. A syntactical chart of ALGOL 60. Communications of the ACM, 4(9):393, September 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1961:RAC

[74] Henry C. Thacher, Jr. Remark on Algorithm 16: CROUT with pivoting. Communications of the ACM, 4(3):154, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Traub:1961:CIM

[75] J. F. Traub. Comparison of iterative methods for the calculation of n-th roots. Communications of the ACM, 4(3):143–145, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Turnep:1961:LEA

[76] Lloyd D. Turnep, David Dahm, Warren Taylor, and Richard E. Waychoff. Letter to the Editor: ALGOL 60 reply. Communications of the ACM, 4 (9):365, September 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinberg:1961:LEA

[77] G. M. Weinberg. Letter to the Editor: ALGOL: pleasure through pain. Communications of the ACM, 4(6):252, June 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkinson:1961:RA

[78] J. H. Wilkinson. Remarks on Algorithms 2 and 3. Communications of the ACM, 4(3):153, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wood:1961:AS

[79] T. C. Wood. Algorithm 35: SIEVE. Communications of the ACM, 4(3): 151, March 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andersen:1962:APC

[80] Christian Andersen. Algol programming: Contribution no. 4: Solution of linear equations. Nordisk tidskrift for informationsbehandling, 2(4): 256–257, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0006-3835&volume=2&issue=4&spage=

256. See remark [114].

Anonymous:1962:APC

[81] Anonymous. Algol programming: Contribution no. 3: Calculation of complementary Fresnel integrals. Nordisk tidskrift for informationsbehandling, 2(3):192–194, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=2&issue=3&spage=192.

Bottenbruch:1962:SUA

[82] H. Bottenbruch. Structure and use of ALGOL 60. Journal of the ACM, 9(2):161–221, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Christiansen:1962:APC

[83] Søren Christiansen. Algol programming: Contribution no. 3: Calculation of complementary Fresnel integrals. Nordisk tidskrift for informationsbehandling, 2(3):192–194, 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cundiff:1962:AEA

[84] John L. Cundiff. Algorithm 88: Evaluation of asymptotic expression for the Fresnel sine and cosine integrals. Communications of the ACM, 5(5): 280, May 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cundiff:1962:AEFa

[85] John L. Cundiff. Algorithm 89: Evaluation of the Fresnel sine integral. Communications of the ACM, 5(5):280, May 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cundiff:1962:AEFb

[86] John L. Cundiff. Algorithm 90: Evaluation of the Fresnel cosine integral. Communications of the ACM, 5(5):281, May 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dahl:1962:LOS

[87] Ole-Johan Dahl. Litt om symbolbruken i Algol. (Norwegian) [A little about symbol use in Algol]. Nordisk tidskrift for informationsbehandling, 2(1):7–8, March 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=2&issue=1&spage=7.

Dahlstrand:1962:HYE

[88] Ingemar Dahlstrand. A half year’s experience with the Facit-Algol 1 compiler. Nordisk tidskrift for informationsbehandling, 2(3):137–142, September 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=2&issue=3&spage=137.

DeMaine:1962:DCP

[89] Paul A. D. De Maine and Robert D. Seawright. Digital computer programs for physical chemistry ALGOL 60 translations of the FORTRAN programs. Macmillan Publishing Company, New York, NY, USA, 1962. 2 v pp.

Dijkstra:1962:OEA

[90] E. W. Dijkstra. Operating experience with ALGOL 60. The Computer Journal, 5(2):125–127, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/2/125.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/050125.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/125.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/126.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
127.tif.

Dijkstra:1962:PAR

[91] Edsger W. Dijkstra. A primer of ALGOL 60: report on the algorithmic language ALGOL 60, volume 2 of APIC studies in data processing. Academic Press, New York, USA, 1962. ISBN 0-12-216250-1. ISSN 0067-2483. xi + 114 pp. LCCN QA76.73.A24.

Duncan:1962:EIA

[92] F. G. Duncan. Errata: “Implementation of ALGOL 60 for the English Electric KDF9”. The Computer Journal, 5(3):176, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/176.tif. See [93].

Duncan:1962:IAE

[93] F. G. Duncan. Implementation of Algol 60 for the English Electric KDF9. The Computer Journal, 5(2):130–131, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/2/130.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/050130.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/131.tif. See er-

rata [92].

Floyd:1962:NPS

[94] Robert W. Floyd. On the nonexistence of a phrase structure grammar for ALGOL 60. Communications of the ACM, 5(9):483–484, September 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Floyd:1962:SAO

[95] Robert W. Floyd. On syntactic analysis and operator precedence. Report CA-62-2, Computer Associates, Inc., Woburn, MA, USA, August

1962. 31 pp. URL https://archives.stanford.edu/catalog/sc0625_
aspace_ref138_9bj.

Forsythe:1962:VDA

[96] G. E. Forsythe, J. von der Groeben, and J. G. Toole. Vectorcardiographic diagnosis with the aid of ALGOL. Communications of the ACM, 5(2): 118–122, February 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ginsburg:1962:TFL

[97] Seymour Ginsburg and H. Gordon Rice. Two families of languages related to ALGOL. Journal of the ACM, 9(3):350–371, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grau:1962:FNR

[98] A. A. Grau. On a floating-point number representation for use with algorithmic languages. Communications of the ACM, 5(3):160–161, March 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heinhold:1962:FPW

[99] J. Heinhold and F. L. Bauer. Fachbegriffe der Programmierungstechnik. Wörterverzeichnis für die Programmierung von Digital-Rechenanlagen mit Stichworten in fünf Sprachen. 2. verbess. und durch ein ALGOL-Wörterbuch erweiterte Auflage. (German) [Technical terms of programming technology. List of words for programming digital computers with key words in five languages. Second improved and extended edition with an ALGOL dictionary]. R. Oldenbourg, München, Germany, 1962. 52 pp. LCCN QA76.5 .G4 1962.

Hoare:1962:REA

[100] C. A. R. Hoare. Report on the Elliott Algol translator. The Computer Journal, 5(2):127–129, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/2/127.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/050127.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/127.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/128.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
129.tif.

Hockney:1962:AAE

[101] R. W. Hockney. ABS12 ALGOL: an extension to Algol 60 for industrial use. The Computer Journal, 4(4):292–300, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/4/4/292.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
04/040292.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/292.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_04/tiff/293.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
294.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/295.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/296.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/297.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/298.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/299.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_04/tiff/300.tif.

Naur:1962:BAS

[102] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, A. J. Perlis, S. Ruti shauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die algorithmische Sprache ALGOL 60. (German) [Report on the ALGOL 60 language]. Elektronisches Rechnen und Regeln. Sonderband 1. Akademie-Verlag, Berlin, Germany, 1962. xi + 49 pp. LCCN QA76.5 .N3515. Zur Erinnerung an William Turanski (In memory of William Turanski).

Naur:1962:RRA

[103] Peter Naur. Revised report on the algorithmic language ALGOL 60. Report ??, Regnecentralen,, Copenhagen, Denmark, 1962. ?? pp.

Ross:1962:ATL

[104] Douglas T. Ross. An algorithmic theory of language. Technical Memorandum ESL-TM-156 (AD 296 998), Electronic Systems Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, November 1962. ix + 68 + 9 pp. URL https://web.archive.org/

web/20130626171225/http://www.dtic.mil/cgi-bin/GetTRDoc?AD=
AD0296998.

Schwarz:1962:IAT

[105] H. R. Schwarz. An introduction to ALGOL: a tutorial paper on ALGOL with explanations and examples to make the use of the ALGOL report more familiar. Communications of the ACM, 5(2):82–95, February 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1962:RCA

[106] Henry C. Thacher, Jr. A redundancy check for ALGOL programs. Communications of the ACM, 5(6):337–342, June 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegstein:1962:SLS

[107] J. H. Wegstein and W. W. Youden. A string language for symbol manipulation based on ALGOL 60. Communications of the ACM, 5(1):54–61, January 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wynn:1962:AAP

[108] P. Wynn. An arsenal of ALGOL procedures for complex arithmetic. Nordisk tidskrift for informationsbehandling, 2(4):232–255, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=2&issue=4&spage=232.

Anonymous:1963:APC

[109] Anonymous. Algol programming: Contribution no. 5: Computation of the Fermi function. Nordisk tidskrift for informationsbehandling, 3 (2):141–142, June 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0006-3835&volume=3&issue=2&spage=
141.

Anonymous:1963:OAR

[110] Anonymous. Official actions and responses to ALGOL as a programming language. Communications of the ACM, 6(4):159–160, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Backus:1963:BAS

[111] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Bericht über die Algorithmische Sprache ALGOL 60 (ALGOL Bulletin Supplement No. 3). Zur Erinnerung an William Turanski. Herausgegeben von P. Naur. (German) [Report on the algorithmic language ALGOL 60 (ALGOL Bulletin supplement no. 3). In memory of William Turanski. Edited by P. Naur]. Elektronische Datenverarbeitung, 2(??):1–17, 1963. ISSN 0374-3012.

Backus:1963:RRAa

[112] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. The Computer Journal, 5 (4):349–367, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/5/4/349.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050349.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/349.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/350.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/351.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/352.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/353.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/tiff/355.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
356.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/357.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/358.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/359.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/tiff/362.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
363.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/365.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/367.tif;   https://standardpascaline.org/Algol60-
RevisedReport.pdf.

Backus:1963:RRAb

[113] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, 4:420– 453, December 1963. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL https://link.springer.com/article/

10.1007/BF01386340. Edited by Peter Naur. Dedicated to the memory of William Turanski.

Bohman:1963:APR

[114] Jan Bohman. Algol programming: Remark on algorithm no. 4, LL-GAUSS (Chr. Andersen, BIT 2, 4 (1962)). Nordisk tidskrift for informationsbehandling, 3(1):60, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=3&issue=1&spage=60. See [80].

Brown:1963:NPN

[115] Peter J. Brown. Note on the proof of the non-existence of a phrase structure grammar for ALGOL 60. Communications of the ACM, 6(3): 105, March 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1963:A

[116] E. W. Dijkstra. ,,Algol 60“. Statistica Neerlandica, 17(2):187–188, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00317.x.

Dijkstra:1963:ATP

[117] E. W. Dijkstra. Appendix — two papers on an ALGOL translator for the X1. Annual Review in Automatic Programming, 3:327–??, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/B9780080097633500175;   http://www.sciencedirect.com/science/
article/pii/S0066413863800142.

Dijkstra:1963:ATX

[118] Edsger W. Dijkstra. An ALGOL-60 translator for the X1. Annual Review in Automatic Programming, 3(??):329–345, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Dijkstra:1963:DMI

[119] E. W. Dijkstra. On the design of machine independent programming languages. Annual Review in Automatic Programming, 3(0):27–42, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/B9780080097633500072;   http://www.sciencedirect.com/science/

article/pii/S0066413863800038. Originally issued as Report M.R.34 of the Mathematics Centre, Amsterdam.

Dijkstra:1963:MTA

[120] Edsger W. Dijkstra. Making a translator for ALGOL-60. Annual Review in Automatic Programming, 3(??):347–356, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Duncan:1963:IOA

[121] F. G. Duncan. Input and output for ALGOL 60 on KDF 9. The Computer Journal, 5(4):341–344, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/4/341.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/050341.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/tiff/342.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
343.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/344.tif.

Fettis:1963:AMH

[122] Henry E. Fettis. Algorithm 163: Modified Hankel function. Communications of the ACM, 6(4):161–162, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Floyd:1963:SAO

[123] Robert W. Floyd. Syntactic analysis and operator precedence. Journal of the ACM, 10:316–333, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Froberg:1963:APC

[124] Carl-Erik Fröberg. Algol programming: Contribution no. 5: Computation of the Fermi function. Nordisk tidskrift for informationsbehandling, 3 (2):141–142, 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gerard:1963:HRA

[125] J. M. Gerard and A. Sambles. A hardware representation for AL-GOL 60 using Creed teleprinter equipment. The Computer Journal, 5(4):338–340, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/5/4/338.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050338.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/tiff/338.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/339.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/340.tif.

Ginsburg:1963:SRU

[126] Seymour Ginsburg and Gene F. Rose. Some recursively unsolvable problems in ALGOL-like languages. Journal of the ACM, 10(1):29–47, January 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Golub:1963:RBG

[127] Gene H. Golub. Review: A Guide to ALGOL Programming, by Daniel D. McCracken. J. Amer. Statist. Assoc., 58(304):1202, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283384.

Golub:1963:RGA

[128] Gene H. Golub. Review: A Guide to ALGOL Programming by Daniel D. McCracken. J. Amer. Statist. Assoc., 58(304):1202, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283384.

Higman:1963:WES

[129] B. Higman. What EVERYBODY should know about ALGOL. The Computer Journal, 6(1):50–56, April 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/6/1/50.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/060050.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/51.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
52.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/55.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
56.tif.

Hoare:1963:EAI

[130] C. A. R. Hoare. The Elliott ALGOL input/output system. The Computer Journal, 5(4):345–348, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/4/345.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/050345.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_05/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/tiff/346.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
347.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/348.tif.

Huskey:1963:SDB

[131] H. D. Huskey, Ralph Love, and Niklaus Wirth. A syntactic description of BC NELIAC. Communications of the ACM, 6(7):367–375, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaupe:1963:NDE

[132] Arthur F. Kaupe, Jr. A note on the dangling else ALGOL 60. Communications of the ACM, 6(8):460, August 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:1963:APG

[133] Peter Naur. Algol programming: Goto statements and good Algol style. Nordisk tidskrift for informationsbehandling, 3(3):204–208, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=3&issue=3&spage=204.

Naur:1963:CAC

[134] P. Naur. Certification of algorithm for calculating primes by means of GPS. Communications of the ACM, 6(1):39, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [59].

Naur:1963:DGAa

[135] Peter Naur. The design of the GIER ALGOL compiler. Part I. Nordisk tidskrift for informationsbehandling, 3(2):124–140, June 1963. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=3&issue=2&spage=124.

Naur:1963:DGAb

[136] Peter Naur. The design of the GIER ALGOL compiler. Part II. Nordisk tidskrift for informationsbehandling, 3(3):145–166, September 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=3&issue=3&spage=145.

Naur:1963:DPA

[137] Peter Naur. Documentation problems: ALGOL 60. Communications of the ACM, 6(3):77–79, March 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:1963:RRA

[138] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, and P. Nauer. Revised report on the algorithmic language ALGOL 60. Communications of the ACM, 6 (1):1–17, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Dedicated to the memory of William Turanski.

Relph:1963:ACH

[139] A. P. Relph. Algorithm 192: Confluent hypergeometric. Communications of the ACM, 6(7):388, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Relph:1963:AH

[140] A. P. Relph. Algorithm 191: Hypergeometric. Communications of the ACM, 6(7):388, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [724].

Rosin:1963:LEB

[141] Robert F. Rosin. Letter to the Editor: BCD data in ALGOL 58. Communications of the ACM, 6(2):51, February 1963.

Rutishauser:1963:URPa

[142] H. Rutishauser. The use of recursive procedures in ALGOL 60. Annual Review in Automatic Programming, 3:43–51, January 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Rutishauser:1963:URPb

[143] H. Rutishauser. The use of recursive procedures in ALGOL 60. In International Tracts in Computer Science and Technology and Their Application, pages 43–51. Elsevier, Amsterdam, The Netherlands, 1963.

Samet:1963:CHR

[144] Paul A. Samet. Correspondence: A hardware representation for ALGOL 60 using Creed Teleprinter equipment. The Computer Journal, 6(3):286, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/286.tif. See correspondence [156, 194].

Shoffner:1963:PLS

[145] Miriam G. Shoffner and Peter J. Brown. Programming languages: A suggested method of making fuller use of strings in ALGOL 60. Communications of the ACM, 6(4):169–171, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Utman:1963:SAR

[146] R. E. Utman. Suggestions on ALGOL 60 (ROME) issues. Communications of the ACM, 6(1):20–23, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanWijngaarden:1963:GA

[147] A. van Wijngaarden. Generalized ALGOL. Annual Review in Automatic Programming, 3:17–26, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Watt:1963:RAP

[148] J. M. Watt. The realization of ALGOL procedures and designational expressions. The Computer Journal, 5(4):332–337, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/4/332.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/050332.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/tiff/333.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
334.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/336.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/337.
tif.

Wells:1963:LEA

[149] Mark B. Wells. Letter to the Editor: on ALGOL 60 as a publication language. Communications of the ACM, 6(8):421, August 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1963:GAa

[150] Niklaus Emil Wirth. A Generalization of Algol. Ph.D. thesis, University of California, Berkeley, Berkeley, CA, 1963. 107 pp. URL https://www.

proquest.com/pqdtglobal/docview/302115469.

Wirth:1963:GAb

[151] Niklaus Wirth. A generalization of ALGOL. Communications of the ACM, 6(9):547–554, September 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woodger:1963:SAR

[152] M. Woodger. Supplement to the ALGOL 60 report. Communications of the ACM, 6(1):18–20, January 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andersen:1964:AMF

[153] Christian Andersen. The QD-algorithm as a method for finding the roots of a polynomial equation when all roots are positive. Technical Report STAN-CS-64-9 (AD604012), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1964. 74 pp. URL http:/

/i.stanford.edu/TR/CS-TR-64-9.html.

Anonymous:1964:APd

[154] Anonymous. Algol programming. Nordisk tidskrift for informationsbehandling, 4(4):250–260, December 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=4&issue=4&spage=250.

Anonymous:1964:NGA

[155] Anonymous. Notes: The GIER ALGOL III compiler. Nordisk tidskrift for informationsbehandling, 4(2):135–136, June 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&issue=2&spage=135.

Barron:1964:CRA

[156] D. W. Barron. Correspondence: On the representation of ALGOL. The Computer Journal, 6(4):335, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_06/Issue_04/tiff/335.tif. See

[144, 194].

Baumann:1964:IA

[157] R. Baumann, M. Feliciano, F. L. Bauer, and K. Samelson. Introduction to Algol. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1964. x + 142 pp. LCCN QA76.5 .B33.

Cheatham:1964:ENS

[158] T. E. Cheatham, Jr. Editor’s note: A successor to ALGOL? Communications of the ACM, 7(7):422, July 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cyvin:1964:AGF

[159] S. J. Cyvin and B. N. Cyvin. Algorithm 225: Gamma function with controlled accuracy. Communications of the ACM, 7(5):295, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cyvin:1964:AND

[160] S. J. Cyvin. Algorithm 226: Normal distribution function. Communications of the ACM, 7(5):295, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farkas:1964:RPPb

[161] I. Farkas. Recent publications and presentations: A Guide to AL-GOL Programming, by Daniel D. McCracken. American Mathematical Monthly, 71(2):232, February 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forsythe:1964:TPA

[162] George E. Forsythe. Tests of Parlett’s ALGOL eigenvalue procedure Eig3 (in Technical Notes and Short Papers). Mathematics of Computation, 18(87):486–487, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.ams.org/journals/

mcom/1964-18-087/S0025-5718-1964-0165669-4/S0025-5718-1964-
0165669-4.pdf;   https://www.math.utah.edu/pub/tex/bib/mathcomp1960.
bib.

Garwick:1964:RFG

[163] Jan V. Garwick. Remark on further generalization of ALGOL. Communications of the ACM, 7(7):422–423, July 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:AIB

[164] Walter Gautschi. Algorithm 222: Incomplete beta functions ratios. Communications of the ACM, 7(3):143, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golub:1964:NMS

[165] Gene H. Golub and Peter A. Businger. Numerical methods for solving linear least squares problems (by G. Golub); An ALGOL procedure for finding linear least squares solutions (by Peter Businger). Technical Report STAN-CS-64-12 (AD608292), Stanford University, Department of Computer Science, Stanford, CA, USA, August 28, 1964. ii + 27 pp.

URL http://i.stanford.edu/TR/CS-TR-64-12.html.

Gorn:1964:RIO

[166] S. Gorn. Report on input-output procedures for ALGOL 60. Communications of the ACM, 7(10):628–630, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gorn:1964:RSA

[167] S. Gorn. Report on SUBSET ALGOL 60 (IFIP). Communications of the ACM, 7(10):626–628, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gram:1964:APC

[168] Christian Gram. Algol programming: Contribution no. 8: Definite integrals by Romberg’s method. Nordisk tidskrift for informationsbehandling, 4(1):54–60, March 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=4&issue=1&spage=54.

Gram:1964:APN

[169] Christian Gram. Algol programming: Note on “Definite Integral by Romberg’s Method” [BIT 4(1), 1964, pp. 54–60]. Nordisk tidskrift for informationsbehandling, 4(2):118–119, June 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&issue=2&spage=118.

Healy:1964:ANN

[170] M. J. R. Healy. ALGOL note: Note on the use of procedures. Communications of the ACM, 7(12):735, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Healy:1964:BRBb

[171] M. J. R. Healy. Book review: An Introduction to ALGOL Programming, by R. Wooldridge; J. F. Ractliffe. Journal of the Royal Statistical Society. Series A (General), 127(2):300, ???? 1964. CODEN JSSAEF. ISSN

0035-9238. URL https://www.jstor.org/stable/2344021.

Hoare:1964:BRA

[172] C. A. R. Hoare. Book review: Algol on the KDF9. The Computer Journal, 7(2):154, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/7/2/154.full.pdf+html.

Hoare:1964:BRB

[173] C. A. R. Hoare. Book review: Algol on the KDF9. The Computer Journal, 7(2):154, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/7/2/154.full.pdf+html.

Hoare:1964:EAP

[174] C. A. R. Hoare. The Elliott ALGOL programming system. In P. Wegner, editor, Introduction to System Programming, pages 156–166. Academic

Press, New York, USA, 1964. URL https://archive.org/details/
introductiontosy00wegn/page/2.

Hohn:1964:BRI

[175] Franz E. Hohn. Book review: Introduction to ALGOL by R. Baumann, M. Feliciano, F. L. Bauer, K. Samelson. Pi Mu Epsilon Journal, 4(1): 31, Fall 1964. CODEN PMEJBR. ISSN 0031-952X. URL https://www.

jstor.org/stable/24339270.

Josefsen:1964:APC

[176] Harold Josefsen. Algol programming: Contribution no. 10: Linear programming by the modified simplex method. Nordisk tidskrift for informationsbehandling, 4(3):189–196, September 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&issue=3&spage=189. See certifi-

cation [248, 634, 635].

Knuth:1964:FDS

[177] Donald E. Knuth and J. L. McNeley. A formal definition of SOL. IEEE Transactions on Electronic Computers, EC-13(4):409–414, August 1964. CODEN IEECA8. ISSN 0367-7508. URL http://

ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=4038205;
https://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=4038205.

Knuth:1964:PIO

[178] D. E. Knuth, L. L. Bumgarner, D. E. Hamilton, P. Z. Ingerman, M. P. Lietzke, J. N. Merner, and D. T. Ross. A proposal for input-output conventions in ALGOL 60. Communications of the ACM, 7(5):273–283, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Russian translation by M. I. Ageev in Sovremennoe Programmirovanie 1 (Moscow: Soviet Radio, 1966), 73–107.

Langefors:1964:AGP

[179] Börje Langefors. ALGOL-GENIUS, a programming language for general data processing. Nordisk tidskrift for informationsbehandling, 4(3): 162–176, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0006-3835&volume=4&issue=3&spage=
162.

Lietzke:1964:MSC

[180] Marjorie P. Lietzke. A method of syntax-checking ALGOL 60. Communications of the ACM, 7(8):475–478, August 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lotsch:1964:AFI

[181] Helmut Lotsch and Malcolm Gray. Algorithm 244: Fresnel integrals [S20]. Communications of the ACM, 7(11):660–661, November 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Metropolis:1964:IQRb

[182] N. C. Metropolis. ICR quarterly report. Technical Report TID-21413, Institute for Computer Research, University of Chicago, Chicago, IL, USA, August 1, 1964. 231 pp. URL http://www.osti.gov/

energycitations/product.biblio.jsp?osti_id=4659841&query_id=
0.

Naur:1964:AGS

[183] Peter Naur. Automatic grading of students’ ALGOL programming. Nordisk tidskrift for informationsbehandling, 4(3):177–188, September 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=4&issue=3&spage=177.

Naur:1964:APC

[184] Peter Naur. Algol programming: Contribution no. 9: Eigenvalues and eigenvectors of real symmetric matrices. Nordisk tidskrift for informationsbehandling, 4(2):120–130, 1964. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See erratum [185].

Naur:1964:EAP

[185] Peter Naur. Erratum: “Algol Programming: Contribution no. 9: Eigenvalues and eigenvectors of real symmetric matrices” [BIT 4(2), 1964, p. 127]. Nordisk tidskrift for informationsbehandling, 4(3):204, 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [184].

Naur:1964:UMC

[186] Peter Naur. Using machine-code within an ALGOL system. Nordisk tidskrift for informationsbehandling, 4(2):115–117, June 1964. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=4&issue=2&spage=115.

Perlis:1964:EAM

[187] A. J. Perlis and Renato Iturriaga. An extension to ALGOL for manipulating formulae. Communications of the ACM, 7(2):127–130, February 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pullin:1964:FAT

[188] D. Pullin. A Fortran to Algol translator. The Computer Journal, 7(1):24– 27, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/7/
1/24.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/070024.sgm.abs.html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/24.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/25.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/26.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/27.tif.

Randell:1964:SST

[189] B. Randell and L. J. Russell. Single-scan techniques for the translation of arithmetic expressions in ALGOL 60. Journal of the ACM, 11(2):159– 167, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rose:1964:EAL

[190] Gene F. Rose. An extension of ALGOL-like languages. Communications of the ACM, 7(2):52–61, February 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ryder:1964:NAC

[191] K. L. Ryder. Note on an Algol 60 compiler for Pegasus I. The Computer Journal, 6(4):336–338, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/6/4/336.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
04/060336.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/336.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_04/tiff/337.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
338.tif.

Samet:1964:CAG

[192] P. A. Samet and P. J. Taylor. Comments on the ALCOR group representation of ALGOL symbols. Communications of the ACM, 7(7):424, July 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Szameitat:1964:BRG

[193] K. E. J. W. Szameitat. Book review: A Guide to Algol Programming by D. D. McCracken. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 32(3):327–328, 1964. CODEN ISTRDP. ISSN 0373-1138 (print), 2212-1846 (electronic). URL

https://www.jstor.org/stable/1401880.

Taylor:1964:CRA

[194] P. Taylor. Correspondence: On the representation of ALGOL. The Computer Journal, 6(4):335–335, 367–367, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/335.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
04/tiff/367.tif. See [144, 156].

Voss:1964:DAP

[195] J. R. Voss. In defense of ALGOL 60 as a publication language. Communications of the ACM, 7(1):2, January 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woodger:1964:A

[196] Michael Woodger. Algol. IEEE Transactions on Electronic Computers, EC-13(4):377–381, August 1964. CODEN IEECA8. ISSN 0367-

7508. URL https://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038200.

Wynn:1964:GPV

[197] P. Wynn. General purpose vector epsilon algorithm ALGOL procedures. Numerische Mathematik, 6(1):22–36, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ageev:1965:EEM

[198] M. I. Ageev and J. Maclaren. English equivalents of metalinguistic terms of Russian ALGOL. Technical Report STAN-CS-65-19 (AD0612686, TT-65-61839), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1965. 115 pp. URL https://apps.dtic.mil/

sti/citations/tr/AD0612686.

Anderson:1965:PSP

[199] James P. Anderson. Program structures for parallel processing. Communications of the ACM, 8(12):786–788, December 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barnes:1965:KAL

[200] J. G. P. Barnes. A KDF9 Algol list-processing scheme. The Computer Journal, 8(2):113–119, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/8/2/113.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/080113.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
02/tiff/113.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/114.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/115.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/116.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/118.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/tiff/119.tif.

Batson:1965:OST

[201] Alan Batson. The organization of symbol tables. Communications of the ACM, 8(2):111–112, February 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burkhardt:1965:LES

[202] Walter H. Burkhardt. Letter to the Editor: syntax and generalization of ALGOL 60. Communications of the ACM, 8(5):261, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Letter to the editor.

Burkhardt:1965:MSS

[203] Walter H. Burkhardt. Metalanguage and syntax specification. Communications of the ACM, 8(5):304–305, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Christiansen:1965:APE

[204] S. Christiansen. Algol programming: Error integral with complex argument. Nordisk tidskrift for informationsbehandling, 5(4):287–293, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=5&issue=4&spage=287.

Chu:1965:ALC

[205] Yaohan Chu. An ALGOL-like computer design language. Communications of the ACM, 8(10):607–615, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duncan:1965:PRO

[206] F. G. Duncan. Possibilities for refining an object program compiled with an ALGOL translator. Nordisk tidskrift for informationsbehandling, 5(2): 85–95, June 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=5&issue=2&spage=85.

Forsythe:1965:AGPa

[207] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Technical Report CS-TR-65-17, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1965. ii + 17 pp.

URL http://i.stanford.edu/pub/cstr/reports/cs/tr/65/17/CS-
TR-65-17.pdf;   http://i.stanford.edu/TR/CS-TR-65-17.html.

Forsythe:1965:AGPb

[208] George E. Forsythe and Niklaus Wirth. Automatic grading programs. Communications of the ACM, 8(5):275–278, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Galler:1965:IE

[209] Bernard A. Galler and Michael J. Fischer. The iteration element. Communications of the ACM, 8(6):349, June 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1965:ALF

[210] W. Gautschi. Algorithm 259: Legendre functions for arguments larger than one [S16]. Communications of the ACM, 8(8):488–492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [860].

Gries:1965:STU

[211] D. Gries, M. Paul, and H. R. Wiehle. Some techniques used in the ALCOR ILLINOIS 7090. Communications of the ACM, 8(8):496–500, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Griffiths:1965:PCS

[212] T. V. Griffiths. On procedures for constructing structural descriptions for three parsing algorithms. Communications of the ACM, 8(10):594, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Letter to the editor.

Griffiths:1965:REC

[213] T. V. Griffiths and S. R. Petrick. On the relative efficiencies of contextfree grammar recognizers. Communications of the ACM, 8(5):289–300, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [212].

Gunn:1965:ASa

[214] J. H. Gunn. Algorithm 260: 6-J symbols. Communications of the ACM, 8(8):492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1965:ASb

[215] J. H. Gunn. Algorithm 261: 9-J symbols. Communications of the ACM, 8(8):492–493, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1965:AZV

[216] J. H. Gunn. Algorithm 252 [Z]: Vector coupling or Clebsch–Gordan coefficients. Communications of the ACM, 8(4):217, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Haag:1965:LEM

[217] James N. Haag. Letter to the Editor: more suggestions for ALGOL 60. Communications of the ACM, 8(6):347, June 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Letter to the editor.

Herriot:1965:RA

[218] J. G. Herriot. R65-26 ALGOL. IEEE Transactions on Electronic Computers, EC-14(2):278, April 1965. CODEN IEECA8. ISSN 0367-

7508. URL https://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&
arnumber=4038434.

Hoare:1965:RH

[219] C. A. R. Hoare. Record handling. WG 2.1 Working Paper ??, [to do], [to do], November 1965. ?? pp.

Hockney:1965:MMF

[220] Roger W. Hockney. Minimum multiplication Fourier analysis. Technical Report CS-TR-65-32, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1965. ?? pp. URL http://i.

stanford.edu/TR/CS-TR-65-32.html.

Hopgood:1965:BRC

[221] F. R. A. Hopgood. Book review: A Course on Programming in Algol 60 by C. M. Reeves, M. Wells. Mathematical Gazette, 49(370):484–485, 1965. CODEN MAGAAS. ISSN 0025-5572. URL http://www.jstor.

org/stable/3612239.

Iverson:1965:LER

[222] Kenneth E. Iverson. Letter to the Editor: Remarks on syntax and generalization of ALGOL 60. Communications of the ACM, 8(9):542, September 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Letter to the editor.

Jensen:1965:GMC

[223] Jørn Jensen. Generation of machine code in ALGOL compilers. Nordisk tidskrift for informationsbehandling, 5(4):235–245, December 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=5&issue=4&spage=235.

Johnston:1965:CUC

[224] John B. Johnston. A class of unambiguous computer languages. Communications of the ACM, 8(3):147–149, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kanner:1965:SYA

[225] H. Kanner, P. Kosinski, and C. L. Robinson. The structure of yet another ALGOL compiler. Communications of the ACM, 8(7):427–438, July 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landin:1965:CBAa

[226] P. J. Landin. A correspondence between ALGOL 60 and Church’s lambda-notation: Part I. Communications of the ACM, 8(2):89–101, February 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landin:1965:CBAb

[227] P. J. Landin. A correspondence between ALGOL 60 and Church’s lambda-notation: Part II. Communications of the ACM, 8(3):158–165, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lynch:1965:RSC

[228] W. C. Lynch. Recursive solution of a class of combinatorial problems: An example. Communications of the ACM, 8(10):617–620, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

MacLaren:1965:APN

[229] M. D. MacLaren. Algorithm 272: Procedure for the normal distribution functions [S15]. Communications of the ACM, 8(12):789–790, December 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [303, 356].

Markman:1965:RZF

[230] B. Markman. The Riemann zeta function. Nordisk tidskrift for informationsbehandling, 5(2):138–141, 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

McKeeman:1965:AAR

[231] W. M. McKeeman. Algorithm 268: ALGOL 60 reference language editor [R2]. Communications of the ACM, 8(11):667–668, November 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Metropolis:1965:IQR

[232] N. C. Metropolis. ICR quarterly report, number 4. Technical Report COO-614-31, Institute for Computer Research, University of Chicago, Chicago, IL, USA, February 1, 1965. 89 pp.

URL http://www.osti.gov/energycitations/product.biblio.jsp?
osti_id=4632649&query_id=0.

Naur:1965:COT

[233] Peter Naur. Checking of operand types in ALGOL compilers. Nordisk tidskrift for informationsbehandling, 5(3):151–163, September 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=5&issue=3&spage=151.

Naur:1965:PIA

[234] Peter Naur. Proposals for introduction on aims. WG 2.1 Working Paper ??, [to do], [to do], October 1965. ?? pp.

Naur:1965:PSA

[235] Peter Naur. The performance of a system for automatic segmentation of programs within an ALGOL compiler (GIER ALGOL). Communications of the ACM, 8(11):671–676, November 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:1965:RRA

[236] P. Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, Heinz Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL-60. (Russian), volume Moskau, Russia. Verlag “Mir”, 1965. 80 pp.

Samet:1965:EAB

[237] P. A. Samet. The efficient administration of blocks in Algol. The Computer Journal, 8(1):21–23, April 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/8/1/21.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/080021.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/22.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
23.tif.

Seegmueller:1965:PBR

[238] G. Seegmueller. A proposal for a basis for a report on a successor to ALGOL 60. Report ??, Bavarian Academy of Science, Munich, West Germany, October 1965. ?? pp.

Thorelli:1965:APC

[239] Lars Erik Thorelli. Algol programming: Contribution no. 13: Evaluation of a function expression described by a numerical vector. Nordisk tidskrift for informationsbehandling, 5(2):133–141, June 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=5&issue=2&spage=133.

vanWijngaarden:1965:ODD

[240] A. van Wijngaarden. Orthogonal design and description of a formal language. Report MR 76, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, October 1965. ?? pp.

Weil:1965:TUD

[241] Roman L. Weil, Jr. Testing the understanding of the difference between call by name and call by value in ALGOL 60. Communications of the ACM, 8(6):378, June 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1965:CDA

[242] Niklaus Wirth and C. A. R. Hoare. A contribution to the development of ALGOL. Report 35, Stanford University, Department of Computer Science, Stanford, CA, USA, December 1965. ?? pp. Revised February 1966. Published in [281].

Wirth:1965:EGA

[243] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL, and its formal definition. Technical Report CS-TR-65-20, Stanford University, Department of Computer Science, Stanford, CA, USA, April 27,

1965. xv + 115 pp. URL http://i.stanford.edu/pub/cstr/reports/
cs/tr/65/20/CS-TR-65-20.pdf;   http://i.stanford.edu/TR/CS-
TR-65-20.html.

Wirth:1965:PRS

[244] Niklaus Wirth. A proposal for a report on a successor of ALGOL 60. Report MR 75, Mathematisch Centrum, Amsterdam, The Netherlands, August 1965. ii + 36 pp. URL https://ir.cwi.nl/pub/9209. IFIP WG 2.1 Working Document.

Zachariassen:1965:APC

[245] J. Zachariassen. Algol programming: Contribution no. 12: Solution of systems of linear equations by Gaussian elimination. Nordisk tidskrift for informationsbehandling, 5(1):61–67, March 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=5&issue=1&spage=61.

Zaremba:1965:AC

[246] W. A. Zaremba. On ALGOL I/O conventions. Communications of the ACM, 8(3):167–169, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abrahams:1966:FSD

[247] Paul W. Abrahams. A final solution to the dangling else of ALGOL 60 and related languages. Communications of the ACM, 9(9):679–682, September 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Asker:1966:APC

[248] Bengt Asker. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(3), 1964, pp. 189–196]. Nordisk tidskrift for informationsbehandling, 6(1):82, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=6&issue=1&spage=82. See [176, 634, 635].

Boussard:1966:ACC

[249] J. C. Boussard. An ALGOL compiler: construction and use in relation to an elaborate operating system. Communications of the ACM, 9(3):179– 182, March 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carr:1966:BRI

[250] John W. Carr III. Book review: Introduction to ALGOL (R. Baumann, M. Feliciano, F. L. Bauer and K. Samelson). SIAM Review, 8(3):406– 407, July 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200

(electronic). URL https://www.jstor.org/stable/2028240.

Dahl:1966:SAB

[251] Ole-Johan Dahl and Kristen Nygaard. SIMULA, an ALGOL-based simulation language. Communications of the ACM, 9(9):671–678, September 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1966:OSAd

[252] Edsger W. Dijkstra. THE operating system — table of the ALGOL trans-

lator. Web archive, 1966. URL http://archive.computerhistory.
org/resources/text/Knuth_Don_X4100/PDF_index/k-9-pdf/k-9-c1063-
4-THEOS-Translator-Table.pdf.

Dijkstra:1966:OSAe

[253] Edsger W. Dijkstra. THE operating system — ALGOL lex scan-

ner. Web archive, 1966. URL http://archive.computerhistory.org/
resources/text/Knuth_Don_X4100/PDF_index/KnuthDigitalArchive-
Index.html#C1063-5.

Fischer:1966:BRK

[254] K. Fischer. Book report: K. Nicot, Elementary Programming and Algol. Computing, 1(4):376, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Galler:1966:PDA

[255] B. A. Galler and A. J. Perlis. A proposal for definitions in ALGOL. Communications of the ACM, 9(7):481–482, July 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garwick:1966:RAX

[256] J. V. Garwick, J. M. Merner, P. Z. Ingerman, and M. Paul. Report on the ALGOL X I-O Subcommittee W.G. 2.1. Working paper ??, [to do], [to do], July 1966. ?? pp.

Gautschi:1966:AD

[257] Walter Gautschi. Algorithm 282: Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 9(4):272, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [488].

Geppert:1966:BBJ

[258] M. G. Geppert. Buchbesprechung: Backus, J. W., F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. Van Wijngarden und M. Woodger: Bericht über die algorithmische Sprache Algol 60. (Report on the algorithmic language Algol 60. Übers. Kerner J. O. u. Mitarb.). Akademie-Verlag Berlin 1962. vi + 49 S., Preis MDN 10,50. Biometrische Zeitschrift, 8 (1-2):121, 1966. CODEN BIZEB3. ISSN 0006-3452.

Hellgren:1966:APC

[259] Bo Hellgren. Algol programming: Contribution no. 18: Procedures for simplifying Boolean expressions. Nordisk tidskrift for informationsbehandling, 6(3):260–269, September 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=6&issue=3&spage=260.

Hoare:1966:PSA

[260] C. A. R. Hoare and N. Wirth. A proposal for a successor to ALGOL 60. Communications of the ACM, 9(7):??, July 1966. URL ????.

Kurtz:1966:DA

[261] Shulom Kurtz. In defense of ALGOL. Communications of the ACM, 9(3): 135, March 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKeeman:1966:ACL

[262] W. M. McKeeman. An approach to computer language design. Technical Report STAN-CS-66-48 (AD0639166, SS639-166), Stanford University, Department of Computer Science, Stanford, CA, USA, August

31, 1966. iv + 124 pp. URL https://apps.dtic.mil/sti/citations/
AD0639166.

Parnas:1966:FPM

[263] David L. Parnas. On facilitating parallel and multiprocessing in ALGOL. Communications of the ACM, 9(4):257, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1966:LDF

[264] David L. Parnas. A language for describing the functions of synchronous systems. Communications of the ACM, 9(2):72–76, February 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). SFD-ALGOL.

Parnas:1966:STA

[265] David L. Parnas. State table analysis of programs in an ALGOL-like language. Communications of the ACM, 9(7):481, July 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pike:1966:ALG

[266] M. C. Pike and I. D. Hill. Algorithm 291: Logarithm of gamma function. Communications of the ACM, 9(9):684, September 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ruhe:1966:APE

[267] Axel Ruhe. Algol programming: Eigenvalues of a complex matrix by the QR-method. Nordisk tidskrift for informationsbehandling, 6(4):350– 358, July 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=6&issue=4&spage=350.

Ryan:1966:CSI

[268] J. L. Ryan, R. L. Crandall, and M. C. Medwedeff. A conversational system for incremental compilation and execution & in a time-sharing environment. fjcc, pages 1–21, 1966.

Stetter:1966:BRJb

[269] H. J. Stetter. Book review: J. Hawgood, Numerical Methods in ALGOL. Computing, 1(4):373, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Svejgaard:1966:APC

[270] Bj. Svejgaard. Algol programming: Contribution no. 17: List processing. Nordisk tidskrift for informationsbehandling, 6(2):164–175, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=6&issue=2&spage=164.

Theissen:1966:AUBa

[271] E. Theissen. Automatische Umwandlung von BELL-Programmen in ALGOL-Programme. (German) [Automatic conversion from BELL programs to ALGOL programs]. Computing, 1(4):354–357, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Theissen:1966:AUBb

[272] E. Theissen. Algorithmus 3: Umwandlung von BELL-Programmen in ALGOL-Programme. (German) [Algorithm 3: Conversion from BELL programs to ALGOL programs]. Computing, 1(4):358, December 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Tienari:1966:SPM

[273] M. Tienari and V. Suokonautio. A set of procedures making real arithmetic of unlimited accuracy possible within Algol 60. Nordisk tidskrift for informationsbehandling, 6(4):332–338, July 1966. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=6&issue=4&spage=332.

Trundle:1966:LAL

[274] R. W. L. Trundle. LITHP — an ALGOL list processor. The Computer Journal, 9(2):167–172, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/9/2/167.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
02/090167.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/tiff/168.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
169.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/170.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/171.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/172.
tif.

vanWijngaarden:1966:DPA

[275] A. van Wijngaarden and B. J. Mailloux. A draft proposal for the algorithmic language ALGOL X. WG 2.1 Working Paper 47 (Warsaw 2), [to do], [to do], October 1966. ?? pp.

vanWijngaarden:1966:NAI

[276] Adriaan van Wijngaarden. Numerical analysis as an independent science. Nordisk tidskrift for informationsbehandling, 6(1):66–81, March 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=6&issue=1&spage=66.

Varah:1966:CPA

[277] J. M. Varah. Certification of Parlett’s ALGOL eigenvalue procedure Eig 3 (in Technical Notes and Short Papers). Mathematics of Computation, 20(95):437–438, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://www.math.utah.edu/pub/tex/
bib/mathcomp1960.bib.

Varah:1966:ERM

[278] James M. Varah. Eigenvectors of a real matrix by inverse iteration. Technical Report CS-TR-66-34, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1966. ?? pp. URL http:/

/i.stanford.edu/TR/CS-TR-66-34.html.

Veinott:1966:PDT

[279] Cyril G. Veinott. Programming decision tables in FORTRAN, COBOL or ALGOL. Communications of the ACM, 9(1):31–35, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1966:CDAa

[280] N. Wirth and C. Hoare. A contribution to the development of ALGOL. Technical Report STAN-CS-66-35 (PB176758), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1966. 64 pp.

Wirth:1966:CDAb

[281] Niklaus Wirth and C. A. R. Hoare. A contribution to the development of ALGOL. Communications of the ACM, 9(6):413–432, June 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [282]. This is the defining document for the Algol W language; see

Wirth:1966:ECD

[282] Niklaus Wirth and C. A. R. Hoare. Errata: “A Contribution to the Development of ALGOL”. Communications of the ACM, 9(12):878, December 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [281].

Wirth:1966:EEG

[283] Niklaus Wirth and Helmut Weber. Errata: “EULER: a generalization of ALGOL and it formal definition: Part 1”. Communications of the ACM, 9(12):878, December 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [284].

Wirth:1966:EGAa

[284] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL and its formal definition: Part 1. Communications of the ACM, 9(1): 13–25, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [283].

Wirth:1966:EGAb

[285] Niklaus Wirth and Helmut Weber. EULER: a generalization of ALGOL, and its formal definition: Part II. Communications of the ACM, 9(2): 89–99, February 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bjorck:1967:API

[286] Åke Björck and Gene H. Golub. Algol programming: Iterative refinement of linear least squares solutions by Householder transformation. BIT (Nordisk tidskrift for informationsbehandling), 7(4):322–337, December 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=7&issue=4&spage=322.

Bond:1967:AAF

[287] Gillian Bond and M. L. V. Pitteway. Algorithm 301: Airy function. Communications of the ACM, 10(5):291–292, May 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bottenbruch:1967:SUA

[288] H. Bottenbruch. 1.6. structure and use of ALGOL 60. In Klerer and Korn [1276], pages 1–181–1–238. LCCN QA76.5 .K524.

Christ:1967:PFT

[289] H. Christ. Prozeduren Für Triplexrechnungen In Algol X8. Report ??, Rechenzentrum, Universität Karlsruhe, Karlsruhe, West Germany, 1967. ?? pp.

Duncan:1967:ABNa

[290] F. G. Duncan. ALGOL bulletin no. 25. ACM SIGPLAN Notices, 2 (5):1–30, May 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.org/doi/

10.1145/1139522.1139526.

Duncan:1967:ABNb

[291] F. G. Duncan. ALGOL Bulletin no. 26. ACM SIGPLAN Notices, 2 (11):1–49, November 1967. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.org/

doi/10.1145/1139498.1139500.

Froberg:1967:APC

[292] Carl-Erik Fröberg and Arne Sundström. Algol programming: Contribution no. 20: Smith’s normal form. BIT (Nordisk tidskrift for informationsbehandling), 7(2):163–169, June 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=7&issue=2&spage=163.

Galler:1967:PDA

[293] B. A. Galler and A. J. Perlis. A proposal for definitions in ALGOL. Communications of the ACM, 10(4):204–219, April 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garwick:1967:RAX

[294] J. V. Garwick, J. M. Merner, P. Z. Ingerman, and M. Paul. Report on the ALGOL X I-O Subcommittee W.G. 2.1. Working paper ??, [to do], [to do], July 1967. ?? pp.

Golub:1967:CAC

[295] Gene H. Golub and Lyle B. Smith. Chebyshev approximation of continuous functions by a Chebyshev system of functions. Technical Report STAN-CS-67-72 (PB175581), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1967. 54 pp. URL http://i.

stanford.edu/TR/CS-TR-67-72.html.

Golub:1967:LSS

[296] Gene H. Golub and Peter Businger. Least squares, singular values and matrix approximations (and an ALGOL procedure for computing the singular value decomposition). Technical Report STAN-CS-67-73 (AD662883), Stanford University, Department of Computer Science, Stanford, CA, USA, July 31, 1967. i + 22 pp. URL https://apps.

dtic.mil/sti/html/tr/AD0662883.

Goos:1967:APA

[297] G. Goos, H. Scheidig, G. Seegmüller, and H. Walther. Another proposal for ALGOL 67. Report 6704, Bavarian Academy of Science, Munich, West Germany, May 1967. 111 pp. URL https://

softwarepreservation.computerhistory.org/ALGOL/report/Goos-
et_al-Another_Proposal_for_ALGOL67.pdf.

Grau:1967:TA

[298] A. A. Grau, U. Hill, and H. Langmaack. Translation of ALGOL 60, volume 137 of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-86937-8 (e-book), 3-642-86939-4 (print). ISSN 0072-7830 (print), 2196-9701 (electronic). ix + 403 pp. LCCN QA1-939. URL

http://www.springerlink.com/content/978-3-642-86937-2.

Gunn:1967:ACW

[299] J. H. Gunn. Algorithm 300: Coulomb wave functions. Communications of the ACM, 10(4):244–245, April 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [700].

Hansen:1967:AAF

[300] Erik Hansen. Alfons (Algol Fortran nesting sequence). Report Risø-M-564, Atomenergikommissionens Bibliotek, Risø, Denmark, 1967. 10 pp.

Hill:1967:ACS

[301] I. D. Hill and M. C. Pike. Algorithm 299: Chi-squared integral. Communications of the ACM, 10(4):243–244, April 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [1081, 798].

Hill:1967:ANCa

[302] I. D. Hill and S. A. Joyce. Algorithm 304: Normal curve integral. Communications of the ACM, 10(6):374–375, June 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [303, 332].

Hill:1967:RAS

[303] I. D. Hill and S. A. Joyce. Remarks on Algorithm 123 [S15]: Real error function, ERF(x); Algorithm 180 [S15]: Error function — large X; Algorithm 181 [S15]: Complementary error function — large X; Algorithm 209 [S15]: Gauss; Algorithm 226 [S15]: Normal distribution function; Algorithm 272 [S15]: Procedure for the normal distribution functions; Algorithm 304 [S15]: Normal curve integral. Communications of the ACM, 10(6):377–378, June 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [160, 229, 302].

Hopgood:1967:AAP

[304] F. R. A. Hopgood and A. G. Bell. The Atlas ALGOL preprocessor for non-standard dialects. The Computer Journal, 9(4): 360–364, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/9/4/360.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090360.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/tiff/360.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/362.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/364.tif.

Jorrand:1967:IDL

[305] P. Jorrand. Intersection de deux langages “context-free”. application à la grammaire du langage ALGOL X. (French) [Intersection of two “contextfree” languages. Application to the grammar of the language ALGOL X]. Report ??, Mathematiques Appliquées, Université de Grenoble, Grenoble, France, October 1967. ?? pp.

Knuth:1967:RTS

[306] Donald E. Knuth. The remaining trouble spots in ALGOL 60. Communications of the ACM, 10(10):611–618, October 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [448]. Reprinted in E. Horowitz, Programming Languages: A Grand Tour (Computer Science Press, 1982), 61–68.

Kunzi:1967:NMM

[307] H. P. Künzi, H. G. Tzschach, and C. A. Zehnder. Numerische Methoden der mathematischen Optimierung, mit ALGOL- und FORTRAN-Programmen. (German), volume Band 8 of Leitfäden der angewandten Mathematik und Mechanik. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1967. 151 pp.

Mailloux:1967:IA

[308] Barry James Mailloux. On the implementation of Algol 68. Technical report, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1967. ?? pp.

McEwan:1967:AAA

[309] A. T. McEwan. An Atlas Autocode to ALGOL 60 translator. The Computer Journal, 9(4):353–359, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/9/4/353.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/090353.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/353.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/tiff/354.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
355.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/356.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/357.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/358.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/359.tif.

Meyer:1967:CLP

[310] Albert R. Meyer and Dennis M. Ritchie. The complexity of loop programs. In Solomon Rosenthal, editor, ACM ’67: Proceedings of the 1967 22nd national conference, pages 465–569. ACM Press, New York, NY 10036, USA, January 1967. ISBN 1-4503-7494-8. URL https://dl.

acm.org/doi/abs/10.1145/800196.806014. Cited in [1258].

Nickel:1967:ZNR

[311] K. Nickel. Zwei Neue Rechenmaschinen-Systeme An Der Th Karlsruhe: Hydra-X8 Und Triplex-Algol-Z23. Umschau, 67:525–526, 1967.

Pike:1967:RAI

[312] M. C. Pike and I. D. Hill. Remark on Algorithm 179: Incomplete Beta ratio. Communications of the ACM, 10(6):375–376, June 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pitteway:1967:RAA

[313] M. L. V. Pitteway. Remark on Algorithm 301: Airy function. Communications of the ACM, 10(7):453, July 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmid:1967:BRK

[314] E. Schmid. Book report: I. O. Kerner und G. Zielke, Einführung in die algorithmische Sprache ALGOL. Computing, 2(2):172, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Sintzoff:1967:EVW

[315] Michel Sintzoff. Existence of a van Wijngaarden syntax for every recursively enumerable set. Annales de la Société Scientifique de Bruxelles, 81(2):115–118, ???? 1967. ISSN 0037-959X.

Stetter:1967:BRE

[316] H. J. Stetter. Book report: T. Ekman und C. E. Fröberg, Introduction to ALGOL-Programming. Computing, 2(3):294, September 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Svejgaard:1967:APC

[317] Bj. Svejgaard. Algol programming: Contribution no. 21: Zeros of polynomials. BIT (Nordisk tidskrift for informationsbehandling), 7(3):240–246, September 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=7&issue=3&spage=240.

vanWijngaarden:1967:DPA

[318] A. van Wijngaarden, B. J. Mailloux, and J. E. L. Peck. A draft proposal for the algorithmic language ALGOL 67. Report MR 88, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, May 1967. ?? pp.

vonSydow:1967:CTA

[319] L. von Sydow. Computer typesetting of ALGOL. Communications of the ACM, 10(3):172–174, March 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watts:1967:RCD

[320] D. G. Watts and H. S. Stone. R67-15: A contribution to the development of ALGOL. IEEE Transactions on Electronic Computers, EC-16(1):112– 113, February 1967. CODEN IEECA8. ISSN 0367-7508. URL https:/

/ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=4039011.

Wilner:1967:AFT

[321] Wayne T. Wilner. ALGOL to FORTRAN translator. Report CGTM 15 (CS239), Stanford University, Stanford, CA, USA, Spring 1967.

?? pp. URL https://www.slac.stanford.edu/vault/collvault/
greylit/cgtm/CGTM15.pdf.

Wippermann:1967:ACT

[322] H.-W. Wippermann. Ein Algol-60 Compiler Mit Triplex-Zahlen. Zeitschrift für Angewandte Mathematik und Mechanik, 47:T76–T79, 1967. CODEN ZAMMAX. ISSN 0044-2267.

Wippermann:1967:MFS

[323] H.-W. Wippermann. Manual Für Das System Triplex-Algol Karlsruhe. Report ??, Institut F. Angewandte Mathematik — Rechenzentrum, Universität Karlsruhe, Karlsruhe, West Germany, 1967. ?? pp.

Wippermann:1967:RIE

[324] H.-W. Wippermann. Realisierung Einer Intervall-Arithmetik In Einem Algol 60-System. Elektronische Rechenanlagen, 9:224–233, 1967. CO-DEN ELRAA4. ISSN 0013-5720.

Wirth:1967:CBC

[325] Niklaus Wirth. On certain basic concepts of programming languages. Technical Report CS-TR-67-65, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1, 1967. ii + 30 pp.

URL http://i.stanford.edu/pub/cstr/reports/cs/tr/67/65/CS-
TR-67-65.pdf;   http://i.stanford.edu/TR/CS-TR-67-65.html.

Apostolatos:1968:ALT

[326] N. Apostolatos, U. Kulisch, R. Krawczyk, B. Lortz, K. Nickel, and H.-W. W. Wippermann. The algorithmic language Triplex-ALGOL-60. Numerische Mathematik, 11(2):175–180, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). From the first page: “triplex numbers, previously called Fehlerschrankenzahlen by K. Nickel and interval numbers by R. E. Moore. … Triplex numbers are intervals with a main value enclosed, which corresponds to conventional numbers of real type. The first number following [ is the lower bound of the interval associated with a triplex number, the next number is the previous mentioned main value, and the last number followed by ] is the upper bound of the interval. The following relation has to be fulfilled: lower bound ≤ main value ≤ upper bound.”.

Apostolatos:1968:RCA

[327] N. Apostolatos, H. Christ, H. Santo, and H.-W. Wippermann. Rounding control and the algorithmic language Algol 68. Report 6824, Rechenzentrum Der Universität Karlsruhe, Karlsruhe, West Germany, 1968. ?? pp.

Arnal:1968:PTD

[328] P. Arnal, J. Buffet, A. Quere, et al. Projet de traduction du rapport ALGOL 68. (French) [Translation project of the ALGOL 68 report ]. Report ??, Faculté des Sciences de Lille, Laboratoire de Calcul, Lille, France, ???? 1968. ?? pp.

Bauer:1968:AW

[329] Henry R. Bauer, Sheldon I. Becker, and Susan L. Graham. ALGOL W. Technical Report CS-TR-68-86, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1968. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-68-86.html.

Bauer:1968:AWI

[330] Henry R. Bauer, Sheldon I. Becker, and Susan L. Graham. ALGOL W implementation. Technical Report CS-TR-68-98, Stanford University, Department of Computer Science, Stanford, CA, USA, May 20, 1968.

vi + 147 pp. URL http://i.stanford.edu/TR/CS-TR-68-98.html;
 http://infolab.stanford.edu/TR/CS-TR-68-98.html. From the in-

troduction: “The project was initiated and directed by Professor Niklaus Wirth, who proposed many of the ideas incorporated in the compiler and suggested ways to bring them about. Joseph W. Wells, Jr. and Edwin H. Satterthwaite, Jr. wrote the PL360 System in which the compiler is embedded, the linkages to the compiler, and the loader.”.

Bauer:1968:AWR

[331] Henry R. Bauer, Sheldon I. Becker, Susan L. Graham, George E. Forsythe, and Edwin H. Satterthwaite. ALGOL W (revised). Technical Report CS-TR-68-89, Stanford University, Department of Computer Science, Stanford, CA, USA, March 1968. i + 124 pp. URL http://i.

stanford.edu/TR/CS-TR-68-89.html.

Bergson:1968:ACR

[332] A. Bergson. Certification of and remark on Algorithm 304 [S15]: Normal curve integral. Communications of the ACM, 11(4):271, April 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [302, 303].

Bertman:1968:APC

[333] B. Bertman. An Algol program for calibrating resistance thermometers at low temperatures. Journal of Computational Physics, 3(1):157– 158, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900120.

Bjorck:1968:IRL

[334] Åke Björck and Gene H. Golub. Iterative refinements of linear least squares solutions by Householder transformations. Technical Report STAN-CS-68-83 (AD664237), Stanford University, Department of Computer Science, Stanford, CA, USA, January 19, 1968. i + 28 pp. URL

http://i.stanford.edu/TR/CS-TR-68-83.html.

Bjork:1968:DTA

[335] Harry Björk. Decision tables in ALGOL 60. BIT (Nordisk tidskrift for informationsbehandling), 8(3):147–153, September 1968. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=8&issue=3&spage=147.

Christ:1968:RMB

[336] H. Christ. Realisierung Einer Maschinenintervallarithmetik Auf Beliebigen Algol-60 Compilern. Elektronische Rechenanlagen, 10:217–222, 1968. CODEN ELRAA4. ISSN 0013-5720.

deChastellier:1968:WG

[337] Gue de Chastellier and Alain Colmerauer. W-grammar. In ACM, editor, Proceedings of the 24th National Conference of the ACM, 1969, pages 511–518. ACM Press, New York, NY 10036, USA, 1968.

DeVogelaere:1968:ASB

[338] R. De Vogelaere. Algorithm 335: A set of basic input-output procedures. Communications of the ACM, 11(8):567–573, August 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1968:SMSb

[339] Edsger W. Dijkstra. The structure of the “THE”-multiprogramming system. Communications of the ACM, 11(5):341–346, May 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dorrer:1968:ADS

[340] Egon Dorrer. Algorithm 322: F -distribution [S14]. Communications of the ACM, 11(2):116–117, February 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duncan:1968:ABNa

[341] F. G. Duncan. ALGOL bulletin no. 27. ACM SIGPLAN Notices, 3 (9):1–65, September 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.org/

doi/10.1145/1139560.1139562.

Duncan:1968:ABNb

[342] F. G. Duncan. ALGOL bulletin no. 28. ACM SIGPLAN Notices, 3 (11):1–83, November 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.org/

doi/10.1145/1139536.1139538.

Duncan:1968:ABNc

[343] F. G. Duncan. ALGOL bulletin no. 29. ACM SIGPLAN Notices, 3 (12):1–69, December 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.org/

doi/10.1145/1139540.1139542.

Enea:1968:M

[344] Horace J. Enea. MLISP. Technical Report CS-TR-68-92 (PB178078), Stanford University, Department of Computer Science, Stanford, CA,

USA, March 1968. ?? pp. URL http://i.stanford.edu/TR/CS-TR-
68-92.html.

Farkas:1968:RIA

[345] I. Farkas. Reviews: Introduction to Algol Programming, by Torgil Ekman and Carl-Erik Fröberg. American Mathematical Monthly, 75(5): 566, May 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fox:1968:BRP

[346] L. Fox. Book review: Procédures Algol en Analyse Numérique, 324 pages. (Published by Min. de l’Éducation Nationale). The Computer Journal, 11(1):30, January 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/11/1/30.full.pdf+html.

Foxley:1968:ISA

[347] Eric Foxley and Peter King. The implementation of syntax analysis using ALGOL, and some mathematical applications. The Computer Journal, 10(4):325–335, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/10/4/325.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/100325.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/325.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_04/tiff/326.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
327.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/329.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/330.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/331.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_04/tiff/333.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
334.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/335.tif.

Geurts:1968:PAD

[348] L. J. M. Geurts, C. H. A. Koster, and L. G. L. Th. Meertens. Praktijkcursus Algol 60: 3e dr. Report MR 85/68, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Nether-

lands, 1968. 53 pp. URL https://ir.cwi.nl/pub/21513.

Ghinea:1968:OMC

[349] Monique Ghinea. Organisation de la mémoire centrale dans la compilation du langage ALGOL 60. (French) [Organization of central memory in compilation of the ALGOL 60 language]. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie, 12 (60)(1):87–92, 1968. ISSN 0373-2908. URL http://www.

jstor.org/stable/43679563.

Good:1968:IAB

[350] D. I. Good and R. L. London. Interval arithmetic for the Burroughs B-5500 four Algol-procedures and proofs of their correctness. Report ??, University of Wisconsin, Madison, Madison, WI, USA, 1968. ?? pp.

Hoare:1968:RDR

[351] C. A. R. Hoare. Recommendation on draft report on ALGOL 68. Unknown, ???? 1968.

Igarashi:1968:A

[352] S. Igarashi, T. Iwamura, K. Sakuma, T. Simauti, T. Simizu, S. Takasu, E. Wada, and N. Yoneda. ALGOL N. Report ??, Research Institute for Mathematical Sciences, Kyoto University, Kyoto, Japan, December 14,

1968. 9 pp. URL http://hdl.handle.net/2433/100407;   https:/
/repository.kulib.kyoto-u.ac.jp/server/api/core/bitstreams/
f79390b7-b7ae-44a9-b8e0-0e8c9bb051c0/content.

Jenkins:1968:TSV

[353] M. A. Jenkins and J. F. Traub. A three-stage variable-shift iteration for polynomial zeros and its relation to generalized Rayleigh iteration. Technical Report STAN-CS-68-107 (AD668558, AD0673673), Stanford University, Department of Computer Science, Stanford, CA, USA, August

26, 1968. viii + 40 + 2 pp. URL http://i.stanford.edu/TR/CS-TR-
68-107.html;   https://apps.dtic.mil/sti/html/tr/AD0673673/
index.html.

Koch:1968:RRA

[354] F. Koch. The recognition of ranges in ALGOL 68. Report ??, University of Calgary, Calgary, AB, Canada, September 1968. ?? pp.

Kunzi:1968:NMM

[355] Hans Paul Kunzi, H. G. Tzschach, and C. A. Zehnder. Numerical methods of mathematical optimization with ALGOL and FORTRAN programs. Computer science and applied mathematics; a series of monographs and textbooks. Academic Press, New York, USA, June 1968. ISBN 0-12-428850-2. x + 171 pp. LCCN ???? URL http://www.

cbooks.com/sqlnut/SP/search/gtsumt?source=&isbn=0124288502.

MacLaren:1968:RAP

[356] M. D. MacLaren. Remark on Algorithm 272: Procedure for the normal distribution functions. Communications of the ACM, 11(7):498, July 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [229].

Mailloux:1968:IA

[357] Barry James Mailloux. On the implementation of Algol 68. PhD thesis, University of Amsterdam, Amsterdam, The Netherlands, June 12, 1968.

36 pp. URL https://ir.cwi.nl/pub/18491/18491A.pdf;   https:/
/lib.uva.nl/discovery/fulldisplay/alma990027594610205131/31UKB_
UAM1_INST:UVA.

Mailloux:1968:PDR

[358] B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Penultimate draft report on the algorithmic language Algol 68, 1: chapters 1–9. Report ??, Mathematisch Centrum, University of Amsterdam, Amsterdam, The Netherlands, 1968. ?? pp. URL https://ir.cwi.nl/pub/9180.

Milner:1968:SHA

[359] R. Milner. String handling in Algol. The Computer Journal, 10(4):321–324, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/4/321.full.pdf+html;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/100321.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/321.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/322.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/323.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/324.tif.

Naur:1968:FGA

[360] Peter Naur. Features of the GIER ALGOL 4 system. BIT (Nordisk tidskrift for informationsbehandling), 8(1):36–42, March 1968. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=8&issue=1&spage=36.

OBrien:1968:CAC

[361] William M. O’Brien and Joan Wood. Certification of Algorithm 299 [S15]: Chi-squared integral. Communications of the ACM, 11(4):271, April 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peck:1968:SA

[362] J. E. L. Peck. The syntax of ALGOL 68. Report ??, Department of Mathematics, University of Calgary, Calgary, AB, Canada, March 1968. ?? pp.

Pham:1968:CBP

[363] Daniel Pham and Monique Ghinea. Compilation des boucles Poue dans le langage ALGOL 60. (French) [Compilation of Poue loops in the AL-GOL 60 language]. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie, 12 (60)(2): 131–141, 1968. ISSN 0373-2908. URL http://www.jstor.org/stable/

43681553.

Pham:1968:RCD

[364] Daniel Pham. Remarques sur la compilation du langage Algol 60. Bulletin mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumanie, 12 (60)(1):109–121, 1968. ISSN 0373-

2908. URL http://www.jstor.org/stable/43679566.

Singleton:1968:AAPa

[365] Richard C. Singleton. Algorithm 338: Algol procedures for the Fast Fourier Transform [C6]. Communications of the ACM, 11(11):773–776, November 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singleton:1968:AAPb

[366] Richard C. Singleton. Algorithm 339: An Algol procedure for the Fast Fourier Transform with arbitrary factors [C6]. Communications of the ACM, 11(11):776–779, November 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanWijngaarden:1968:DRA

[367] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Draft report on the algorithmic language ALGOL 68. Report MR 93, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, January 1968. ?? pp.

vanWijngaarden:1968:FDR

[368] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Final draft report on the algorithmic language ALGOL 68. Report MR 100, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, December 1968. ?? pp.

vanWijngaarden:1968:PDR

[369] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Penultimate draft report on the algorithmic language ALGOL 68. Report MR 99, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, October 1968. ?? pp.

vanWijngaarden:1968:RAL

[370] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report on the algorithmic language ALGOL 68. Report MR 101, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, October 1968. ?? pp.

vanWijngaarden:1968:WDA

[371] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Working document on the algorithmic language ALGOL 68. Report MR 95, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, July 1968. ?? pp.

Veltkamp:1968:APV

[372] G. W. Veltkamp. ALGOL procedures voor het berekenen van een inwendig product in dubbele precisie. (Dutch) [ALGOL procedures for calculating an inner product in double precision]. Technical report 22, RC-Informatie, Technische Hogeschool Eindhoven, Eindhoven, The Netherlands, 1968. ?? pp.

Waite:1968:RPD

[373] W. M. Waite. R68-24 A proposal for definitions in ALGOL. IEEE Transactions on Computers, C-17(5):522–523, May 1968. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=1687389.

Wilkes:1968:OIS

[374] M. V. Wilkes. The outer and inner syntax of a programming language. The Computer Journal, 11(3):260–263, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/11/3/260.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/110260.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/260.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_03/tiff/261.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
262.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/263.tif.

Wippermann:1968:DST

[375] H.-W. Wippermann. Definition von Schrankenzahlen in Triplex-Algol. (German) [Definition of number bounds in Triplex-Algol]. Computing, 3(2):99–109, June 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wippermann:1968:IAS

[376] H.-W. Wippermann. Implementierung Eines Algol-Systems Mit Schrankenzahlen. Angew. Informatik (Elektron. Datenverarbeitung), 10: 189–194, 1968.

Woodward:1968:NPA

[377] P. M. Woodward. A narrative preview of ALGOL 68. RRE Memorandum 2499, Royal Radar Establishment, Malvern, Worcestershire, UK, October 1968. ?? pp.

Agrawala:1969:WDL

[378] Vishnu K. Agrawala and Johan G. Belinfante. Weight diagrams for Lie group representations: A computer implementation of Freudenthal’s algorithm in ALGOL and FORTRAN. BIT (Nordisk tidskrift for informationsbehandling), 9(4):301–314, December 1969. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=9&issue=4&spage=301.

Albrecht:1969:AAP

[379] R. Albrecht and E. Visotschnig. Algorithmus 9. ALGOL-Prozeduren zu den modifizierten Algorithmen nach Minty und Moore. (German) [Algorithm 9. ALGOL-procedures to the modified algorithms according to Minty and Moore]. Computing, 4(1):76–81, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Andrews:1969:PCS

[380] M. P. Andrews. Practical considerations in the storage of modes. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, University of British Columbia, August 1969, pages 78–84. University of British Columbia, Vancouver, BC, Canada, 1969.

Anonymous:1969:APR

[381] Anonymous. Algorithms policy, revised September, 1969 (includes AL-GOL, FORTRAN, and PL/I). Communications of the ACM, 12(9):513, September 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Assabgui:1969:ESA

[382] M. Assabgui and L. Trilling. Entrées–sorties ALGOL 68. (French) [Input–outputs ALGOL 68]. Report 13, Département d’Informatique, Université de Montréal, Montréal, QC, Canada, November 1969. ?? pp.

Bauer:1969:AWI

[383] Henry Bauer, Sheldon Becker, Susan L. Graham, Edwin Satterthwaite, George E. Forsythe, and Robert W. Floyd. Algol W language description. Report CS 110, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1969. ii + 103

pp. URL https://bitsavers.org/pdf/stanford/cs_techReports/
CS110_ALGOL_W_Revised_Sep69.pdf.

Bauer:1969:AWR

[384] Henry R. Bauer, Sheldon I. Becker, Susan L. Graham, Robert W. Floyd, George E. Forsythe, and Edwin H. Satterthwaite. AL-GOL W (revised). Technical Report CS-TR-68-110, Stanford University, Department of Computer Science, Stanford, CA, USA,

September 1969. ?? pp. URL http://i.stanford.edu/TR/
CS-TR-68-110.html;   https://bitsavers.org/pdf/stanford/cs_
techReports/CS110_ALGOL_W_Revised_Sep69.pdf.

Boussard:1969:IA

[385] J. E. Boussard and C. Pair. Introduction — ALGOL 68. Revue française d’Informatique et de Recherche Opérationnelle, (B-3):17–52, December 1969.

Bowlden:1969:ASF

[386] H. J. Bowlden. ALGOL 68 structural flowchart. Report 69-1C4COMPS-R2, Westinghouse Research Laboratories, Churchill, PA, USA, October 1969. ?? pp.

Bowlden:1969:CIA

[387] H. J. Bowlden. A comparative introduction to ALGOL 68. Report 69-1C4-COMPS-P2, Westinghouse Research Laboratories, Churchill, PA, USA, 1969. ?? pp.

Bowlden:1969:EFC

[388] H. J. Bowlden. Environmental factors in computer language design and implementation. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, University of British Columbia, August 1969, pages 97–109. ????, ????, August 1969.

Branquart:1969:OLS

[389] P. Branquart and J. Lewi. On object language and storage allocation in ALGOL 68 compilers. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, University of British Columbia, August 1969, pages 25–34. University of British Columbia, Vancouver, BC, Canada, 1969.

Brockhaus:1969:BFT

[390] M. Brockhaus, B. Rothmaier, and P. Schroth. Benutzeranleitung Für Triplex-Algol Im System Hydra 2. Interner Bericht 69/11, Inst. F. Informatik, Universität Karlsruhe, Karlsruhe, West Germany, 1969. ?? pp.

Brown:1969:CRA

[391] W. E. Brown. The cross-referencing of ALGOL 68 syntax. Report ??, University of Calgary, Calgary, AB, Canada, June 1969. ?? pp.

Cobb:1969:CAS

[392] S. M. Cobb. Certification of Algorithm 47 [S16]: Associated Legendre functions of the first kind for real or imaginary arguments. Communications of the ACM, 12(11):635–636, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:1969:ABA

[393] Jerome A. Feldman and Paul D. Rovner. An Algol-based associative language. Communications of the ACM, 12(8):439–449, August 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Finch:1969:DAO

[394] P. M. Finch. Defining and applied occurrences of identifiers. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 110–117. University of British Columbia, Vancouver, BC, Canada, August 1969.

Fites:1969:SOG

[395] P. E. Fites. Storage organization and garbage collection in ALGOL 68. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 85–96. University of British Columbia, Vancouver, BC, Canada, August 1969.

Gautschi:1969:ACE

[396] Walter Gautschi. Algorithm 363: Complex error function [S15]. Communications of the ACM, 12(11):635, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [626].

Golub:1969:SVR

[397] Gene H. Golub and Richard R. Underwood. Stationary values of the ratio of quadratic forms subject to linear constraints. Technical Report STAN-CS-69-142, Stanford University, Department of Computer Science, Stanford, CA, USA, November 1969. 22 pp. URL http://

i.stanford.edu/TR/CS-TR-69-142.html. Cited in Åke Björck’s bib-

liography on least squares, which is available by anonymous ftp from

math.liu.se in pub/references.

Goos:1969:EEA

[398] G. Goos. Einige Eigenschaften von ALGOL 68. (German) [Some properties of ALGOL 68]. Elektronische Datenverarbeitung, 11(??):??, September 1969.

Goos:1969:IA

[399] G. Goos. Eine implementierung von ALGOL 68. (German) [An implementation of ALGOL 68]. Report 6906, Computing Centre, Technische Hochschule München, Munich, West Germay, ???? 1969. ?? pp.

Goos:1969:MRO

[400] G. Goos, H. Scheidig, and H. Wössner. Mode representation and operator identification in ALGOL 68. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, August 1969, pages 36–41. ????, ????, 1969.

Hellgren:1969:APC

[401] Bo Hellgren. Algol programming: Contribution no. 24: Algorithms for the interpolation of a function defined by a character string. BIT (Nordisk tidskrift for informationsbehandling), 9(3):283–293, September 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=9&issue=3&spage=283. See

remark [612].

Herriot:1969:ADA

[402] John G. Herriot. An ambiguity in the description of ALGOL 60. Communications of the ACM, 12(10):581, October 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1969:FAT

[403] Edward Burlingame Hill. A FORTRAN to ALGOL translator. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1969. 271 pp.

Jorrand:1969:TA

[404] P. Jorrand. Tutorial on ALGOL 68. In ????, editor, Proceedings of the third annual Princeton conference on information sciences and systems, March 1969, pages 403–407. ????, ????, 1969.

Junginger:1969:BRH

[405] W. Junginger. Book review: H. P. Künzi. H. G. Tzschach und C. A. Zehnder, Numerische Methoden der mathematischen Optimierung mit ALGOL- und FORTRAN-Programmen (Leitfäden der angewandten Mathematik und Maechanik, Bd. 8). Computing, 4(1):87, March 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kaldenbach:1969:LGT

[406] P. Kaldenbach. Lineare Gleichungssysteme Und Triplex-Algol. Diplomarbeit, Inst. F. Informatik, University Karlsruhe, 1969. ?? pp.

Kolbig:1969:CASa

[407] K. S. Kölbig. Certification of Algorithm 292 [S22]: Regular Coulomb wave functions and of remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 12(5):278–279, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Korenjak:1969:PMC

[408] A. J. Korenjak. A practical method for constructing LR (k) processors. Communications of the ACM, 12(11):613–623, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koster:1969:SDP

[409] C. H. A. Koster. Syntax directed parsing of ALGOL 68 programs. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, August 1969, pages 61–69. University of British Columbia, Vancouver, BC, Canada, 1969.

Landelle:1969:DTA

[410] A. Landelle and J. Pleyber. A definition of the translation of ALGOL 60 to ALGOL 68. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 49–60. University of British Columbia, Vancouver, BC, Canada, 1969.

Lewi:1969:ILN

[411] J. Lewi and P. Branquart. Implementation of local names in ALGOL 68. Report R121, MBLE Research Laboratory, Brussels, Belgium, November 1969. ?? pp.

Lindsey:1969:ICR

[412] C. H. Lindsey. An ISO-code representation for ALGOL 68. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 1–24. University of British Columbia, Vancouver, BC, Canada, August 1969.

Mailloux:1969:AEL

[413] Barry J. Mailloux and John E. L. Peck. Algol 68 as an extensible language. ACM SIGPLAN Notices, 4(8):9–13, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL https://dl.acm.org/doi/10.1145/1115858.1115861.

Marshall:1969:AGC

[414] S. Marshall. An ALGOL 68 garbage collector. Technical Report TM 0111, Dartmouth College, Hanover, NH, USA, 1969. ?? pp.

Marshall:1969:PRA

[415] S. Marshall. Preliminary report on ALGOL 68 implementation. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 42–48. University of British Columbia, Vancouver, BC, Canada, 1969.

Mazurkiewicz:1969:NEG

[416] A. W. Mazurkiewicz. A note on enumerable grammars. Information and Control, 14(??):555–558, ???? 1969. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

Meek:1969:AXS

[417] B. L. Meek. ALGOL X, some comments. The Computer Bulletin, ?? (??):298, August 1969. CODEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

Meertens:1969:GAP

[418] L. Meertens. On the generation of ALGOL 68 programs involving infinite modes. ACM SIGPLAN Notices, 4(3):90–92, March 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL https://dl.acm.org/doi/10.1145/1316865.1316869.

Mulholland:1969:STT

[419] K. A. Mulholland. Software to translate TELCOMP programs into KDF9ALGOL. The Computer Journal, 12(3):221–224, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/221.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/120221.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/221.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/224.tif.

Nadrchal:1969:GLA

[420] Jaroslav Nadrchal. Guide to the language ALGOL 68. Report ??, Mathematics Institut, Czechoslovakia Academy of Sciences, Praha, Czechoslovakia, ???? 1969 (??). ?? pp.

Nickel:1969:KBS

[421] K. Nickel. Das Kahan-Babuska’sche Summierungsverfahren In Triplex-Algol 60. (German) [The Kahan-Babuška summation method in Triplex-Algol 60]. Interner Bericht Des Inst. F. Informatik 69/3, Universität Karlsruhe, Karlsruhe, Germany, 1969. ?? pp. Published in [511].

Nickel:1969:KST

[422] K. Nickel. Das Kahan-Babuska’sche Summierungsverfahren In Triplex-Algol 60. Interner Bericht Des Inst. F. Informatik 69/3, Universität Karlsruhe, Karlsruhe, West Germany, 1969. ?? pp. Published In: Z. Angew. Math. Mech. 50, 369–373, 1970.

Nickel:1969:TA

[423] K. Nickel. Triplex-Algol and applications. In E. Hansen, editor, Topics In Interval Analysis, pages 10–24. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1969.

Oswald:1969:BPT

[424] W. Oswald. Beschreibung Der Programmiersprache Triplex-Algol Und Erstellung Eines Compilers In Algol 60. Diplomarbeit, Inst. F. Informatik, University Karlsruhe, 1969. ?? pp.

Overholt:1969:APP

[425] K. J. Overholt. Algol programming: Partition functions (modulo d). BIT (Nordisk tidskrift for informationsbehandling), 9(1):83–86, March 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=9&issue=1&spage=83.

Peck:1969:SMS

[426] J. E. L. Peck. On storage of modes and some context conditions. In ????, editor, Proceedings of the Informal Conference on ALGOL 68 Implementation, pages 70–77. University of British Columbia, Vancouver, BC, Canada, 1969.

Peter:1969:ZSGa

[427] R. Peter. Zur zweistufigen Satzstruktur-Grammatik. (German) [On twostage sentence-structure grammar]. Studia Sci. Math. Hung., 2(??):455– 456, ???? 1969 (??).

Peter:1969:ZSGb

[428] R. Peter. Zur zweistufigen Satzstruktur-Grammatik. (German) [On twostage sentence-structure grammar]. Studia Sci. Math. Hung., 3(??):181– 194, ???? 1969 (??).

Roberts:1969:ABL

[429] John B. Roberts. An Algol-based language for defining collateral execution of procedures in general and hybrid programming in particular. ACM SIGPLAN Notices, 4(3):17–37, March 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

https://dl.acm.org/doi/10.1145/1316865.1316866.

Satterthwaite:1969:MEP

[430] Edwin H. Satterthwaite. MUTANT 0.5: an experimental programming language. Technical Report CS-TR-69-120, Stanford University, Department of Computer Science, Stanford, CA, USA, February 1969. ?? pp.

URL http://i.stanford.edu/TR/CS-TR-69-120.html.

Sauer:1969:RAR

[431] G. Sauer. Remark on Algorithm 268 [R2]: ALGOL 60 reference language editor. Communications of the ACM, 12(7):407, July 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singleton:1969:AAC

[432] Richard C. Singleton. Algorithm 345: An Algol convolution procedure based on the Fast Fourier Transform [C6]. Communications of the ACM, 12(3):179–184, March 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singleton:1969:RAC

[433] Richard C. Singleton. Remark on Algorithm 339 [C6]: An Algol procedure for the Fast Fourier Transform with arbitrary factors. Communications of the ACM, 12(3):187, March 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sintzoff:1969:GSA

[434] Michel Sintzoff. Grammaires superposées et autres systèmes formels. (French) [Superimposed grammars and other formal systems]. Report ??, Journées d’Étude sur l’Analyse Syntaxique, Centre d’Automatique Fontainebleau, Fontainebleau, France, 1969. ?? pp.

Sintzoff:1969:IDD

[435] Michel Sintzoff. Introduction — la description d’ALGOL 68. Revue française d’Informatique et de Recherche Opérationnelle, B-3(??):3–16, ???? 1969.

Sintzoff:1969:RDR

[436] Michel Sintzoff, P. Branquart, J. Lewi, and P. L. Wodon. Remarks on the draft reports on ALGOL 68. Report R96, MBLE Research Laboratory, Brussels, Belgium, January 1969. ?? pp.

Sipala:1969:FDA

[437] P. Sipala. Formatting and display of Algol expressions. The Computer Journal, 12(4):365–367, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/12/4/365.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/120365.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/365.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_04/tiff/366.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
367.tif.

Stetter:1969:BRP

[438] H. J. Stetter. Book review: Procédures ALGOL en Analyse Numérique. Computing, 4(4):386, December 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stetter:1969:BRR

[439] H. J. Stetter. Book review: R. Herschel, ALGOL: – Übungen. Computing, 4(2):184, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

TadeudeMedeiros:1969:APF

[440] Adilson Tadeu de Medeiros and Georges Schwachheim. Algorithm 349: Polygamma functions with arbitrary precision [S14]. Communications of the ACM, 12(4):213–214, April 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [771].

vanderMeulen:1969:IIA

[441] S. G. van der Meulen and C. H. Lindsey. Informal introduction to AL-GOL 68: 4th draft. Report ??, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1969.

301 pp. URL https://ir.cwi.nl/pub/6762.

vanWijngaarden:1969:RALa

[442] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report on the algorithmic language ALGOL 68. Numerische Mathematik, 14(2):79–218, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.fh-jena.de/

~kleine/history/languages/Algol68-Report.pdf;   http://www.
fh-jena.de/~kleine/history/languages/Algol68-RevisedReport.
pdf;   https://web.archive.org/web/20080406061108/http://www.
fh-jena.de/~kleine/history/languages/Algol68-Report.pdf.

vanWijngaarden:1969:RALb

[443] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report of Algorithmic Language ALGOL 68. Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. ISBN 3-662-38647-X (print), 3-662-39504-5 (e-book). 79–218 pp. LCCN QA76.9.A43. URL http://

www.springerlink.com/content/978-3-662-39504-2.

vanWijngaarden:1969:SOA

[444] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Soobshchenie ob algoritmicheskom yazyke ALGOL 68. (Russian) [report on the algorithmic language ALGOL 68]. Kibernetika, 6(??):????, 1969. CODEN KBRNA5. ISSN 0023-1274. Text in Russian and English. Translation to Russian by A. A. Bährs, A. P. Ershov, L. L. Zmievskaya and A. F. Rar.

Veltkamp:1969:APV

[445] G. W. Veltkamp. ALGOL procedures voor het rekenen in dubbele lengte. (Dutch) [ALGOL procedures for double-length arithmetic]. Technical report 21, RC-Informatie, Technische Hogeschool Eindhoven, Eindhoven, The Netherlands, 1969. ?? pp.

Wegner:1969:SRV

[446] P. Wegner. Some remarks on VWF notation. Report 69-12, Center for Computing and Information Sciences, Brown University, Providence, RI, USA, December 1969. ?? pp.

Westland:1969:ASP

[447] J. Westland. An ALGOL 68 syntax and parser. Report ??, University of Calgary, Calgary, AB, Canada, September 1969. ?? pp.

Wood:1969:LEF

[448] Derick Wood. Letter to the Editor: A few more trouble spots in ALGOL 60. Communications of the ACM, 12(5):247–248, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [306].

Zaremba:1969:SAI

[449] W. A. Zaremba. A syntax for Algol input/output formats. The Computer Journal, 12(4):342–348, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/12/4/342.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/120342.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/342.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_04/tiff/343.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
344.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/346.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/348.tif.

Abramson:1970:POS

[450] H. D. Abramson. A partial ordering of some of the elements of Algol. BIT (Nordisk tidskrift for informationsbehandling), 10(1):6–9, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=10&issue=1&spage=6.

Ahrens:1970:PRN

[451] J. H. Ahrens, Ulrich Dieter, and A. Grube. Pseudo-random numbers. A new proposal for the choice of multiplicators. Computing, 6(1–2):121– 138, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alefeld:1970:AAA

[452] G. Alefeld and J. Herzberger. Algol-60 Algorithmen zur Auflösung linearer Gleichungssysteme mit Fehlererfassung. (German) [Algol-60 algorithms for solution of linear systems of equation with error control]. Computing, 6(1–2):28–34, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Andrews:1970:CRA

[453] M. P. Andrews and J. E. L. Peck. Cross reference of the ALGOL 68 transput routines. Report ??, University of British Columbia, Vancouver, BC, Canada, March 1970. ?? pp.

Baecker:1970:GCV

[454] H. D. Baecker. Garbage collection for virtual memory systems. Report ??, University of Calgary, Calgary, AB, Canada, November 1970. ?? pp.

Baecker:1970:IAH

[455] H. D. Baecker. Implementing the ALGOL 68 heap. BIT (Nordisk tidskrift for informationsbehandling), 10(4):405–414, December 1970. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=10&issue=4&spage=405.

Baecker:1970:UAT

[456] H. D. Baecker. The use of Algol 68 for trees. The Computer Journal, 13(1):25–27, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/13/1/25.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/130025.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/25.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/26.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/27.tif.

Baker:1970:AVW

[457] J. L. Baker. Acceptors from van Wijngaarden grammars. Technical Report 70-02-10, Computer Science Group, University of Washington, Seattle, WA, USA, February 1970. ?? pp.

Baker:1970:SAP

[458] J. L. Baker. The syntax of ALGOL 68, property grammars, and contextsensitive languages. Report ??, University of Calgary, Calgary, AB, Canada, September 1970. ?? pp.

Baker:1970:SFP

[459] J. L. Baker. Some formal properties of the syntax of ALGOL 68. Technical Report ??, Computer Science Group, University of Washington, Seattle, WA, USA, May 1970. ?? pp.

Batson:1970:MSS

[460] A. Batson, Shy-Ming M. Ju, and D. C. Wood. Measurements of segment size. Communications of the ACM, 13(3):155–159, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berry:1970:IIM

[461] Daniel M. Berry. The importance of implementation models in ALGOL 68: or how to discover the concept of necessary environment. ACM SIG-PLAN Notices, 5(9):14–24, September 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blake:1970:RPP

[462] L. F. Blake, R. E. Lawson, and I. M. Yuille. A ring processing package for use with FORTRAN or a similar high-level language. The Computer Journal, 13(1):40–47, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/13/1/40.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130040.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/41.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/
42.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/45.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/47.tif.

Boussard:1970:REAa

[463] J. C. Boussard and J. J. Duby. Rapport d’évaluation d’ALGOL 68. (French) [ALGOL 68 evaluation report]. Report ??, IMAG and Centre Scientifique IBM France, Grenoble, France, July 1970. ?? pp.

Boussard:1970:REAb

[464] J. C. Boussard and J. J. Duby. Rapport d’évaluation d’ALGOL 68. (French) [ALGOL 68 evaluation report]. Revue française d’Informatique et de Recherche Opérationnelle, B-1(??):??, July 1970.

Bowlden:1970:STB

[465] H. J. Bowlden. A symbol table for the B-6500. Report 70-1K4-COMPS-R1, Westinghouse Research Laboratories, Churchill, PA, USA, April 1970. ?? pp.

Boyle:1970:ASA

[466] J. M. Boyle and A. A. Grau. An algorithmic semantics for ALGOL 60 identifier denotation. Journal of the ACM, 17(2):361–382, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Branquart:1970:APSa

[467] P. Branquart and J. Lewi. Analysis of the parenthesis structure of AL-GOL 68. Report R130, MBLE Research Laboratory, Brussels, Belgium, April 1970. ?? pp.

Branquart:1970:CFS

[468] P. Branquart, J. Lewi, and J. P. Cardinael. A context-free syntax of ALGOL 68. Technical Note N66, MBLE Research Laboratory, Brussels, Belgium, August 1970. ?? pp.

Branquart:1970:GPA

[469] P. Branquart and J. Lewi. General principles of an ALGOL 68 garbage collector. Technical Note N60, MBLE Research Laboratory, Brussels, Belgium, January 1970. ?? pp.

Branquart:1970:ICAa

[470] P. Branquart and L. Lewi. On the implementation of coercions in AL-GOL 68. Report R123, MBLE Research Laboratory, Brussels, Belgium, January 1970. ?? pp.

Branquart:1970:ICAb

[471] P. Branquart and L. Lewi. On the implementation of coercions in AL-GOL 68. In ????, editor, Proceedings of the International Computing Symposium, Bonn, 1970, page ?? ????, ????, 1970.

Branquart:1970:LGA

[472] P. Branquart, J. Lewi, and J. P. Cardinael. Local generators and the ALGOL 68 working stack. Technical Note N62, MBLE Research Laboratory, Brussels, Belgium, September 1970. ?? pp.

Branquart:1970:QAI

[473] P. Branquart and J. Lewi. Quelques aspects de l’implementation d’ALGOL 68. (French) [Some aspects of the implementation of ALGOL 68]. Report ??, Séminaire de Programmation, IMAG 1968–69, Grenoble, France, 1970. ?? pp.

Branquart:1970:SCA

[474] P. Branquart and J. Lewi. Structure d’un compilateur d’ALGOL 68. (French) [Structure of an ALGOL 68 compiler]. Report R130, MBLE Research Laboratory, Brussels, Belgium, April 1970. ?? pp.

Branquart:1970:SCS

[475] P. Branquart, J. Lewi, Michel Sintzoff, and P. L. Wodon. Structural composition of semantics in ALGOL 68. Report R125, MBLE Research Laboratory, Brussels, Belgium, April 1970. ?? pp.

Branquart:1970:SSA

[476] P. Branquart and J. Lewi. A scheme of storage allocation and garbage collection for ALGOL 68. Report R133, MBLE Research Laboratory, Brussels, Belgium, July 1970. ?? pp.

Bray:1970:CAR

[477] T. Bray. Certification of Algorithm 22, Ricatti–Bessel Functions of First and Second Kind. Communications of the ACM, 13(7):448, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Businger:1970:APU

[478] P. Businger. Algol programming: Updating a singular value decomposition. BIT (Nordisk tidskrift for informationsbehandling), 10(3):376– 385, September 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0006-3835&volume=10&issue=3&
spage=376.

Charmonman:1970:CSA

[479] S. Charmonman. A comparison of the structures of APL, FORTRAN, ALGOL and PL/I. ACM SIGPLAN Notices, 5(3):22–24, March 1970.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cocke:1970:PLT

[480] John Cocke and J. T. Schwartz. Programming languages and their compilers: Preliminary notes. Report ??, Courant Institute of Mathematical Sciences New York University, 251 Mercer Street, New York, NY, USA,

April 1970. 771 pp. URL https://www.softwarepreservation.org/
projects/FORTRAN/CockeSchwartz_ProgLangCompilers.pdf. Second

Revised Version.

Cody:1970:RAC

[481] W. J. Cody and Kathleen A. Paciorek. Remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 13(9): 573, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crespi-Reghizzi:1970:LTG

[482] S. Crespi-Reghizzi and R. Morpurgo. A language for treating graphs. Communications of the ACM, 13(5):319–323, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1970:CPA

[483] Edsger W. Dijkstra. Cursus programmeren in Algol 60: 11e uitg. (Dutch) [Course in programming in Algol 60: 11th edition]. Report ??, Mathematisch Centrum, Amsterdam, Amsterdam, The Netherlands, 1970.

?? pp. URL http://www.cs.utexas.edu/users/EWD/MCReps/CR1970-
013.PDF.

Douglas:1970:BRB

[484] A. S. Douglas. Book review: Report on the Algorithmic Language, Algol 68, by A. van Wijngaarden. Journal of the Royal Statistical Society. Series A (General), 133(3):486–487, ???? 1970. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2343572.

Fites:1970:ECA

[485] P. E. Fites. On error classes in ALGOL 68. Report ??, University of Alberta, Edmonton, AB, Canada, April 1970. ?? pp.

Florian:1970:BRD

[486] H. Florian. Book review: D. J. Malcolme-Lawes, Programming-ALGOL. Computing, 6(3–4):376, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRR

[487] H. Florian. Book review: R. Herschel, Anleitung zum praktischen Gebrauch von ALGOL 60, 3. Aufl. Computing, 6(3–4):375, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gautschi:1970:RAD

[488] Walter Gautschi and Bruce J. Klein. Remark on Algorithm 282, Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 13 (1):53–54, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [257].

Goos:1970:IAF

[489] G. Goos and H. Scheidig. Une implementation d’ALGOL 68. (French) [An implementation of ALGOL 68]. Report ??, Séminaire de Programmation, IMAG 1968–69, Grenoble, France, 1970. ?? pp.

Goyer:1970:GCI

[490] Pierre Goyer. A garbage collector to be implemented on a CDC 3100. Report 34, Département d’Informatique, Université de Montréal, Montréal, QC, Canada, April 1970. ?? pp.

Herzberger:1970:IAS

[491] J. Herzberger. Intervallmäßige Auswertung von Standardfunktionen in Algol-60. (German) [ALGOL-60 procedures evaluating standard functions in interval analysis]. Computing, 5(4):377–384, December 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hill:1970:AASa

[492] G. W. Hill. ACM Algorithm 395: Student’s t-distribution. Communications of the ACM, 13(10):617–619, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [936, 1003].

Hill:1970:AASb

[493] G. W. Hill. ACM Algorithm 396: Student’s t-quantiles. Communications of the ACM, 13(10):619–620, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [1003, 1004, 936].

Hill:1970:ARA

[494] V. Hill. Automatische rekursive Adressenberechnung für höhere Programmiersprachen, insbesondere für ALGOL 68. (German) [Automatic recursive address calculation for higher programming languages, especially for ALGOL 68]. Report ??, Technische Hochschule München, Munich, West Germay, February 1970. ?? pp.

Hill:1970:SD

[495] G. W. Hill. ACM Algorithm 395: Student’s t-distribution. Communications of the ACM, 13(10):617–619, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [936, 1003].

Hill:1970:SQ

[496] G. W. Hill. ACM Algorithm 396: Student’s t-quantiles. Communications of the ACM, 13(10):619–620, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [1003, 1004, 936].

Hodgson:1970:AES

[497] G. S. Hodgson. ALGOL 68 extended syntax. Report ??, Department of Computer Science, University of Manchester, Manchester, UK, March 1970. ?? pp.

Hunter:1970:APA

[498] D. B. Hunter. ALGOL programming: Analysis of the outer product of symmetric group representations. BIT (Nordisk tidskrift for informationsbehandling), 10(1):106–114, March 1970. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=10&issue=1&spage=106.

Igarashi:1970:SAL

[499] Shigeru Igarashi. Semantics of ALGOL-like statements. Technical Report STAN-CS-70-167 (AIM-129, AD712460), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1970. iii + 95 pp.

Innis:1970:BRA

[500] George Innis. Book review: Algol 68, by A. Van Wijngaaden, B. M. Mailloux, J. E. L. Peck, C. H. A. Koster. Biometrics, 26(4):871–872, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528749.

Knight:1970:ACP

[501] K. R. Knight. An Algol construction for procedures as parameters of procedures. Communications of the ACM, 13(4):266, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koster:1970:TLG

[502] C. H. A. Koster. Two level grammars. Report ??, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, May 1970. ?? pp.

Krawczyk:1970:IAS

[503] R. Krawczyk. Intervallmäßerung Auswertung von Standardfunktionen in ALGOL-60. (German) [ALGOL-60 procedures evaluating standard functions in interval analysis]. Computing, 5(3):200, September 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Landelle:1970:TAA

[504] A. Landelle and J. Pleyber. Traduction d’ALGOL 60 en ALGOL 68. (French) [Translation of ALGOL 60 into ALGOL 68]. Report ??, Séminaire de Programmation, IMAG 1968–69, Grenoble, France, 1970. ?? pp.

Lewi:1970:ILN

[505] J. Lewi and P. Branquart. Implementation of local names in ALGOL 68. In International Computing Symposium, Bonn, 1970, page ?? ????, ????, 1970.

Mailloux:1970:SOG

[506] B. J. Mailloux and P. E. Fites. Storage organization and garbage collection for ALGOL 68 implementation. Report ??, University of Alberta, Edmonton, AB, Canada, February 1970. ?? pp.

Malcolm:1970:DCS

[507] Michael A. Malcolm. A description and comparison of subroutines for computing Euclidean inner products on the IBM 360. Technical Report STAN-CS-70-175 (AD713842), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1970. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-70-175.html.

Marshall:1970:AI

[508] Sidney Marshall. ALGOL 68 implementation. In J. E. L. Peck, editor, Proceedings of the IFIP Working Conference on ALGOL 68 Implementation, Munich, 20–24 July 1970, pages 239–243. North-Holland, Amsterdam, The Netherlands, 1970.

Moses:1970:FFL

[509] Joel Moses. The function of FUNCTION in LISP or why the FUNARG problem should be called the environment problem. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(15):13–27, July 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Nickel:1970:FNP

[510] K. Nickel. Fehlerschranken zu Näherungswerten von Polynomwurzeln. (German) [Error bounds for approximate values of polynomials]. Computing, 6(1–2):9–27, March 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nickel:1970:KBS

[511] K. Nickel. Das Kahan–Babuškasche Summierungsverfahren in Triplex-ALGOL 60. (German) [The Kahan–Babuška summation method in Triplex-ALGOL 60]. Zeitschrift für Angewandte Mathematik und Mechanik, 50:369–373, 1970. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Parlett:1970:BRA

[512] Beresford N. Parlett. Book review: ALGOL 60 Procedures in Numerical Algebra, Part II by T. J. Dekker, W. Hoffman. Mathematics of Computation, 24(110):483–485, 1970. CODEN MCMPAF. ISSN 0025-5718,

1088-6842. URL http://www.jstor.org/stable/2004503.

Scheidig:1970:AAG

[513] H. Scheidig. Anpassungsoperationen in ALGOL 68. (German) [Adjustment operations in ALGOL 68]. Report ??, Technische Hochschule München, Munich, West Germay, February 1970. ?? pp.

Schneider:1970:OPA

[514] V. B. Schneider. A one-pass algorithm for compiling ALGOL 68 declarations. Report ??, Purdue University, West Lafayette, IN, USA, ???? 1970. ?? pp.

Schneider:1970:TGA

[515] V. B. Schneider. A translation grammar for ALGOL 68. In ????, editor, AFIPS Spring Joint Computer Conference of the 1970 and International Computing Symposium, Bonn, 1970, page ?? AFIPS Press, Montvale, NJ, USA, 1970.

Simonet:1970:GCF

[516] M. Simonet. Une grammaire context-free d’ALGOL 68. (French) [A context-free grammar of ALGOL 68]. In ????, editor, Congrès d’Informatique AFCET, Paris, 1970, pages 5.3.119–5.3.135. ????, ????, 1970.

Sintzoff:1970:CPP

[517] Michel Sintzoff. Calculating the properties of programs by valuation on specific models. Report N64, MBLE Research Laboratory, Brussels, Belgium, May 1970. ?? pp.

vanderMeulen:1970:IIA

[518] S. G. van der Meulen and C. H. Lindsey. Informal introduction to AL-GOL 68: Very informal introduction to ALGOL 68. Mathematisch Centrum, Amsterdam, The Netherlands, 1970. [unpaged] pp. LCCN

QA76.5 .M46. URL https://ir.cwi.nl/pub/9831;   https://ir.
cwi.nl/pub/9832.

vanWijngaarden:1970:BBN

[519] A. van Wijngaarden. On the boundary between natural and artificial languages. In Linguaggi nella societa e nella tecnica. (Italian) [Languages in Society and Technology], pages 165–175. Edizioni di Comunita, Milano, Italy, 1970.

vanWijngaarden:1970:SOA

[520] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Soobshchenie ob algoritmicheskom yazyke ALGOL 68. (Russian) [report on the algorithmic language ALGOL 68]. Kibernetika, 7(??):????, 1970. CODEN KBRNA5. ISSN 0023-1274. Text in Russian and English. Translation to Russian by A. A. Bährs, A. P. Ershov, L. L. Zmievskaya and A. F. Rar.

Wagner:1970:FAI

[521] Robert A. Wagner. Finiteness assumptions and intellectual isolation of computer scientists. Communications of the ACM, 13(12):759–760, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1970:PLP

[522] Niklaus Wirth. The programming language PASCAL. Report 1, ETH Zürich, Zürich, Switzerland, November 1970. v + 58 pp. URL http://

pascal.hansotten.com/uploads/wirth/The_Programming_Language_
Pascal_1970%20001.pdf;   https://www.research-collection.ethz.
ch/bitstream/handle/20.500.11850/68712/eth-3290-01.pdf.

Wodon:1970:MGC

[523] P. L. Wodon. Methods of garbage collection for ALGOL 68. Report R123, MBLE Research Laboratory, Brussels, Belgium, April 1970. ?? pp.

Woodward:1970:PAR

[524] P. M. Woodward. A primer of ALGOL 68-R. RRE Memorandum 2601, Royal Radar Establishment, Malvern, Worcestershire, UK, February 1970. ?? pp.

Anonymous:1971:BA

[525] Anonymous. Bibliography of Algol 68. In Peck [1278], pages 369– 375. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970. URL https://

softwarepreservation.computerhistory.org/ALGOL/paper/ALGOL_
68-Implementation.pdf.

Bahrs:1971:DSA

[526] A. A. Bährs, Andrei P. Ershov, and A. F. Rar. On description of syntax of Algol 68 and its national variants. In Peck [1278], pages 163–177. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Bauer:1971:PDI

[527] Friedrich L. Bauer, Henry J. Bowlden, Paul Branquart, Ian F. Currie, Andrei P. Ershov, Cornelis H. A. Koster, Barry J. Mailloux, and Manfred Paul. Panel discussion — implementation. In Peck [1278], pages 321–343. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Becker:1971:BRH

[528] Otwin Becker. Besprungen: Rutishauser, H.: Handbook for Automatic Computation, Vol. I / Part a: Description of ALGOL 60, Die Grundlehren der mathematischen Wissenschaften, 135. Band. XII, 323 S. DM 58,-. Grau, A. A., U. Hill, H. Langmaack: Handbook for Automatic Computation, Vol. 1 / Part b: Translation of ALGOL 60, Die Grundlehren der mathematischen Wissenschaften, 137. Band. X, 397 S. DM 64,-. Berlin–Heidelberg–New York 1967, Springer-Verlag. (German) [Reviews]. Zeitschrift für die gesamte Staatswissenschaft / Journal of Institutional and Theoretical Economics, 127(2):365–366, May 1971.

CODEN ???? ISSN 0044-2550. URL https://www.jstor.org/stable/
40749460.

Bell:1971:ALA

[529] D. A. Bell and B. A. Wichmann. An ALGOL-like assembly language for a small computer. Software—Practice and Experience, 1(1):61–72, January/March 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bloom:1971:SFI

[530] Howard M. Bloom. A syntax-directed FORTRAN interpreter for AL-GOL 60. Report ??, H. Diamond Laboratories, Washington, DC, USA, 1971. 159 pp. Reproduced by the National Technical Information Service, Springfield, VA.

Bourgeois:1971:EMA

[531] François Bourgeois and Jean-Claude Lassalle. An extension of the Munkres algorithm for the assignment problem to rectangular matrices. Communications of the ACM, 14(12):802–804, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bowlden:1971:STS

[532] Henry J. Bowlden. A symbol table with scope recognition for the B-6500. In Peck [1278], pages 1–19. ISBN 0-7204-2045-8. LCCN QA76.5

.I28 1970. URL https://softwarepreservation.computerhistory.
org/ALGOL/paper/ALGOL_68-Implementation.pdf.

Branquart:1971:APSb

[533] Paul Branquart, Johan Lewi, and Jean-Pierre Cardinael. Analysis of the parenthesis structure of Algol 68. In Peck [1278], pages 37– 76. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970. URL https://

softwarepreservation.computerhistory.org/ALGOL/paper/ALGOL_
68-Implementation.pdf.

Branquart:1971:CSA

[534] P. Branquart, J. Lewi, Michel Sintzoff, and P. L. Wodon. The composition of semantics in Algol 68. Communications of the ACM, 14(11): 697–708, November 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Branquart:1971:SSA

[535] Paul Branquart and Johan Lewi. A scheme of storage allocation and garbage collection for Algol 68. In Peck [1278], pages 199– 238. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970. URL https://

softwarepreservation.computerhistory.org/ALGOL/paper/ALGOL_
68-Implementation.pdf.

Brent:1971:AGC

[536] R. P. Brent. An algorithm with guaranteed convergence for finding a zero of a function. The Computer Journal, 14(4):422–425, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/14/4/422.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_14/
Issue_04/140422.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/422.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/423.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_14/
Issue_04/tiff/424.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/425.tif.

Brown:1971:SCP

[537] F. D. Brown, Valerie J. Calderbank, and M. D. Poole. Some comments on the portability of a large ALGOL program — the implementation of SID on KDF9. Software—Practice and Experience, 1(4):367–371, October/ December 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cermak:1971:AAG

[538] J. Čermák. ACM Algorithm 412: Graph plotter [J6]. Communications of the ACM, 14(7):492–493, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [701].

Currie:1971:AR

[539] Ian F. Currie, Susan G. Bond, and J. D. Morison. Algol 68-R. In Peck [1278], pages 21–34. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Currie:1971:WDA

[540] I. F. Currie. Working description of ALGOL 68-R. RRE Memorandum 2660, Royal Radar Establishment, Malvern, Worcestershire, UK, 1971 (??). ?? pp.

deBakker:1971:MIS

[541] Jaco de Bakker, G. A. Blaauw, A. J. W. Duijvestijn, Edsger Dijkstra, Piet van der Houwen, G. A. M. Kamsteeg-Kemper, F. E. J. Kruseman Aretz, W. L. van der Poel, J. P. Schaap-Kruseman, M. V. Wilkes, and Guus Zoutendijk. MC-25 informatica symposium: [symposium on the occasion of the 25th anniversary of the Mathematical Centre, Amsterdam, 06-07.01.1972]. C, January 1971. ?? pp.

Dekker:1971:FPT

[542] Theodorus J. Dekker. A floating-point technique for extending the available precision. Numerische Mathematik, 18(3):224–242, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://www-gdz.sub.uni-goettingen.de/cgi-bin/
digbib.cgi?PPN362160546_0018.

Ershov:1971:MPS

[543] Andrei P. Ershov. A multilanguage programming system oriented to languages description and universal optimization algorithms. In Peck [1278], pages 143–162. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Evans:1971:AAS

[544] D. J. Evans. Algorithm 69: An algorithm for the solution of symmetric general three term linear systems. The Computer Journal, 14(4): 444–445, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/444.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/445.
tif.

Feldmann:1971:EEU

[545] H. Feldmann, H. Oberquelle, and C.-P. Ortlieb. Eine einfache universelle Turingmaschine in ALGOL 60 Simulation. (German) [A simple universal Turing machine in ALGOL 60 simulation]. Computing, 8(3–4):241–249, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Feldmann:1971:EUT

[546] H. Feldmann, H. Oberquelle, and C.-P. Ortlieb. Eine einfache universelle Turingmaschine in ALGOL 60 Simulation. (German) [A simple universal Turing machine in ALGOL 60 simulation]. Computing, 8(3–4):241–249, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Golub:1971:AAC

[547] G. H. Golub and L. B. Smith. ACM Algorithm 414: Chebyshev approximation of continuous functions by a Chebyshev system of functions [E2].

Communications of the ACM, 14(11):737–746, November 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goos:1971:PDA

[548] Gerhard Goos, Henry J. Bowlden, Paul Branquart, Barry J. Mailloux, John E. L. Peck, and Philip M. Woodward. Panel discussion — Algol 68 sublanguages. In Peck [1278], pages 265– 281, 289–300. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Goos:1971:SPC

[549] Gerhard Goos. Some problems in compiling Algol 68. In Peck [1278], pages 179–196. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Goyer:1971:GCI

[550] Pierre Goyer. A garbage collector to be implemented on a CDC 3100. In Peck [1278], pages 303–317. ISBN 0-7204-2045-8. LCCN QA76.5 .I28

1970. URL https://softwarepreservation.computerhistory.org/
ALGOL/paper/ALGOL_68-Implementation.pdf.

Haddon:1971:ALS

[551] E. W. Haddon and L. G. Proll. An Algol line-syntax checker. The Computer Journal, 14(2):128–132, May 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/2/128.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/140128.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/tiff/129.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/
130.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/131.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/132.tif.

Herriot:1971:APC

[552] J. G. Herriot and Christian H. Reinsch. ALGOL 60 procedures for the calculation of interpolating natural spline functions. Technical Report STAN-CS-71-200 (PB198416), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1971. ii + 30 pp.

Hieber:1971:HSH

[553] L. Hieber and H. Rzehak. HYGOL, eine Sprache für hybride Programme. (German) [HYGOL, a language for hybrid programs]. Computing, 7 (3–4):311–323, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hill:1971:FFA

[554] I. D. Hill. Faults in functions, in Algol and Fortran. The Computer Journal, 14(3):315–316, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/3/315.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/140315.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/315.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/tiff/316.tif.  See

notes [647].

Housden:1971:PSF

[555] R. J. W. Housden. Phrase structures in Fortran. The Computer Journal, 14(3):224–228, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/3/224.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/140224.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/tiff/225.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/
226.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/228.tif.

Igarashi:1971:SAL

[556] Shigeru Igarashi. Semantics of Algol-like statements. Lecture Notes in Mathematics, 188:117–177, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/chapter/10.
1007/BFb0059697/.

Jennings:1971:ASV

[557] A. Jennings. Algorithm 70: Solution of variable bandwidth positive definite simultaneous equations. The Computer Journal, 14(4):446, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/446.tif.

Knuth:1971:NAG

[558] Donald E. Knuth and R. W. Floyd. Notes on avoiding “go to” statements. Information Processing Letters, 1(1):23–31, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [625]. Reprinted in Writings of the Revolution, E. Yourdon, ed. (New York: Yourdon Press, 1982), 153–162.

Koster:1971:AG

[559] Cornelis H. A. Koster. Affix-grammars. In Peck [1278], pages 95– 109. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970. URL https://

softwarepreservation.computerhistory.org/ALGOL/paper/ALGOL_
68-Implementation.pdf.

Kral:1971:IIT

[560] Jaroslav Král and J. Moudrý. An implementation of identifier tables in a multipass Algol 68 compiler based on a hash-code technique. In Peck [1278], pages 77–81. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Kunzi:1971:NMM

[561] Hans Paul Kunzi, H. G. Tzschach, and C. A. Zehnder. Numerical methods of mathematical optimization with ALGOL and FORTRAN programs. Computer science and applied mathematics. Academic Press, New York, USA, corrected and augmented ed. edition, 1971. ISBN 0-12-428850-2. viii + 219 pp. LCCN QA402.5 .K8131 1971. Translated by Werner C. Rheinboldt and Cornelie J. Rheinboldt.

Lindsey:1971:MHS

[562] Charles H. Lindsey. Making the hardware suit the language. In Peck [1278], pages 347–365. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Lindsey:1971:SAS

[563] Charles H. Lindsey. Some Algol 68 sublanguages. In Peck [1278], pages 283–288. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Lu:1971:TAC

[564] Hu quan Lu. The translation of Algol 68 into Chinese. Report ??, Institute of Mathematics, Academia Sinica, Beijing, China, 1971. 3

pp. URL https://archive.computerhistory.org/resources/text/
algol/ACM_Algol_bulletin/1061739/p33-ru_ian.pdf.

Lucas:1971:AAC

[565] C. W. Lucas, Jr. and C. W. Terrill. ACM Algorithm 404: Complex gamma function [S14]. Communications of the ACM, 14(1):48–49, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lunnon:1971:MDM

[566] W. F. Lunnon. Multi-dimensional map-folding. The Computer Journal, 14(1):75–80, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.oxfordjournals.

org/content/14/1/75.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/140075.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/tiff/75.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/76.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/77.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/tiff/79.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/80.tif.

Malcolm:1971:PRP

[567] Michael A. Malcolm. PL360 (revised) a programming language for the IBM 360. Technical Report STAN-CS-71-215 (AD727115), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1971.

v + 94 pp. URL http://infolab.stanford.edu/pub/cstr/reports/
cs/tr/71/215/CS-TR-71-215.pdf;   https://bitsavers.org/pdf/
stanford/cs_techReports/STAN-CS-71-215_PL360_Rev_May72.pdf.

Manning:1971:AWM

[568] J. R. Manning. Algorithm 68: White to move and mate in n moves. The Computer Journal, 14(2):209–213, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/209.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/210.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_14/Issue_02/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/212.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/213.
tif.

Marshall:1971:AGC

[569] S. Marshall. An Algol 68 garbage collector. In Peck [1278], pages 239–243. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

McVitie:1971:AAT

[570] D. G. McVitie and L. B. Wilson. ACM Algorithm 411: Three procedures for the stable marriage problem [H]. Communications of the ACM, 14 (7):491–492, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgana:1971:TMR

[571] Maria Aurora Morgana. Time and memory requirements for solving linear systems. Technical Report STAN-CS-71-212 (AD727107), Stanford University, Department of Computer Science, Stanford, CA, USA, March

1971. ?? pp. URL http://i.stanford.edu/TR/CS-TR-71-212.html.

Niegel:1971:AAE

[572] W. Niegel. Ein Algorithmus zum aufbau einer endlichen Gruppe. (German) [An algorithm for the generation of a finite group]. Computing, 7(3–4):333–343, September 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Peck:1971:AC

[573] J. E. L. Peck. An ALGOL 68 companion. Technical Report TR-71-01, University of British Columbia, Vancouver, BC, Canada, October 1971.

157 pp. URL https://www.cs.ubc.ca/tr/1971/tr-71-01.

Pereira:1971:PNBa

[574] R. Agonia Pereira and Duarte Costa Cabral. Programacão NEAT, BA-SIC, FORTRAN e ALGOL. Centro de Calculo Cenctifico, Oeiras, Portugal, 1971. 336 pp.

Pereira:1971:PNBb

[575] R. Agonia Pereira and Duarte Costa Cabral. Programmacão NEAT, BASIC, FORTRAN e ALGOL. Centro de Calculo Cientifico, Oeiras, Portugal, 1971. 336 pp.

Roberts:1971:SPD

[576] K. V. Roberts and J. P. Boris. The solution of partial differential equations using a symbolic style of Algol. Journal of Computational Physics, 8(1):83–105, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0021999171900374.

Scheidig:1971:SMC

[577] H. Scheidig. Syntax and mode check in an Algol 68 compiler. In Peck [1278], pages 83–92. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Scheiding:1971:REM

[578] H. Scheiding. Representation and equality of modes. Information Processing Letters, 1(2):61–65, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmitt:1971:AAD

[579] H. Schmitt. ACM Algorithm 409: Discrete Chebychev curve fit [E2]. Communications of the ACM, 14(5):355–356, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scowen:1971:SPW

[580] R. S. Scowen, D. Allin, A. L. Hillman, and M. Shimell. SOAP — a program which documents and edits ALGOL 60 programs. The Computer Journal, 14(2):133–135, May 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/2/133.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/140133.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/133.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/tiff/134.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/
135.tif.

Sheldon:1971:CTD

[581] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bσ) and (a, bσ − σ). (III) Fortran translations of the Algol programs ‘Mandy’ and ‘Barbara’. Computer Physics Communications, 2(5):272–287, August/September 1971. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900038.

Sites:1971:ARM

[582] Richard L. Sites. ALGOL with reference manual. Technical Report STAN-CS-71-230 (PB203601), Stanford University, Department of Computer Science, Stanford, CA, USA, August 1971. 141 pp.

Suzuki:1971:IA

[583] N. Suzuki, T. Fujisaki, K. Kakehi, E. Wada, and T. Simauti. The implementation of ALGOL N. ACM SIGPLAN Notices, 6(12):15–22, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tinhofer:1971:FAN

[584] G. Tinhofer. Flüsse in antisymmetrischen Netzwerken. (German) [Flows in antisymmetric networks]. Computing, 8(1–2):191–199, March 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Toshkov:1971:AYA

[585] D. Toshkov and St. Buchvarov. Algoritmichniyat yezik ALGOL 68. (Bulgarian) [The algorithmic language ALGOL 68]. Nauka i Yzkustvo [Science and Art], Sofia, Bulgaria, 1971. ???? pp.

Trilling:1971:ADE

[586] Laurent Trilling and Jean-Pierre Verjus. An attempted definition of an extensible system. In Peck [1278], pages 119–139. ISBN 0-7204-2045-

8. LCCN QA76.5 .I28 1970. URL https://softwarepreservation.
computerhistory.org/ALGOL/paper/ALGOL_68-Implementation.pdf.

vanderMeulen:1971:IIA

[587] S. G. van der Meulen and C. H. Lindsey. Informal introduction to ALGOL 68. North-Holland Pub. Co., Amsterdam, The Netherlands, 1971. ISBN 0-7204-2048-2. viii + 368 , fold. l. pp. LCCN QA76.5 .M46 1971.

vanWijngaarden:1971:RAL

[588] A. van Wijngaarden, editor. Report on the algorithmic language ALGOL 68. [Centrum Obliczeniowe Polskiej Akademii Nauk], Warszawa, Poland, 1971. 81–218 pp. LCCN QA76.73.A24 R45 1971.

Wodon:1971:MGC

[589] Pierre L. Wodon. Methods of garbage collection for Algol 68. In Peck [1278], pages 245–262. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Woodall:1971:ANR

[590] A. D. Woodall. Algorithm 64: A non-recursive tree sort. The Computer Journal, 14(1):104, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_01/tiff/104.tif.

Woodall:1971:ART

[591] A. D. Woodall. Algorithm 63: A recursive tree sort. The Computer Journal, 14(1):103, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_01/tiff/103.tif.

Wossner:1971:IOA

[592] Hans W”ossner. On identification of operators in Algol 68. In Peck [1278], pages 111–118. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

Yoneda:1971:DSA

[593] Nobuo Yoneda. The description and the structure of ALGOL N. ACM SIGPLAN Notices, 6(12):10–14, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1972:BRI

[594] Anonymous. Book review: Informal Introduction to ALGOL 68, by C. H. Lindsey, S. G. Van Der Muelen. Biometrics, 28(1):250–253, March 1972. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528981.

Baecker:1972:MMA

[595] H. D. Baecker. On a missing mode in ALGOL 68. ACM SIGPLAN Notices, 7(12):20–30, December 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bajpai:1972:FAP

[596] Avinash Chandra Bajpai. Fortran and Algol: a programmed course for students of science and technology. A Series of programmes on mathematics for scientists and technologists. John Wiley, New York, NY, USA, 1972. ISBN 0-471-04371-0 (paperback). 21 + 133 + 87 pp.

Baker:1972:UOA

[597] J. L. Baker. An unintentional omission from ALGOL 68. Information Processing Letters, 1(6):244–245, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1972:ABA

[598] W. Barth. Ein Algorithmus zur Berechnung aller rellen Nullstellen in einem Interval. (German) [An algorithm to complete all real roots in an interval]. Computing, 9(4):327–333, December 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bauer:1972:AWL

[599] Henry Bauer, Sheldon Becker, Susan L. Graham, Edwin Satterthwaite, and Richard L. Sites. Algol W language description. Report ??, [unknown], [unknown], June 1972. 54 pp. URL https://www.algol60.

org/docsW/algolw.pdf.

Bayer:1972:OBT

[600] R. Bayer. Oriented balanced trees and equivalence relations. Information Processing Letters, 1(6):226–228, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Branquart:1972:CFS

[601] P. Branquart, J. P. Cardinael, J. P. Delescaille, and J. Lewi. A contextfree syntax of ALGOL 68. Information Processing Letters, 1(4):141–148, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bron:1972:AAM

[602] Coenraad Bron. ACM Algorithm 426: Merge sort algorithm [M1]. Communications of the ACM, 15(5):357–358, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [709].

Buffet:1972:DDL

[603] J. Buffet, P. Arnal, A. Quéré, and A. van Wijngaarden. Définition du langage algorithmique ALGOL 68: Rapport sur le langage algorithmique ALGOL 68. (French) [Definition of the Algorithmic Language ALGOL 68: Report on the Algorithmic Language ALGOL 68], volume 1354 of Actualités scientifiques et industrielles. Hermann, Paris, France, 1972. ISBN 2-7056-1354-1. vii + 222 pp. LCCN Q111 .A3 no. 1354 QA76.73.A24.

Corlett:1972:PP

[604] Peter Norman Corlett, J. D. Tinsley, and R. A. Court. Practical programming. School Mathematics Project Handbooks. University Press, Cambridge [Eng.], second edition, 1972. ISBN 0-521-08198-X. x + 264 pp. LCCN QA76.6 .C66 1972.

Culik:1972:BRJa

[605] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur (editor), A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 2 (1960), pp. 106–136; also Communications of the Association for Computing Machinery, vol. 3 (1960), pp. 299–314. Journal of Symbolic Logic, 37(3):622–623, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL ttps://www.jstor.org/

stable/2272793.

Culik:1972:BRJb

[606] Karel Čulík. Book review: J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. Numerische Mathematik, vol. 4 (1962–1963), pp. 420–453; also ibid., vol. 6 (1963), pp. 1–17; and The Computer Journal, vol. 5 no. 4 (1963), pp. 349–367. Journal of Symbolic Logic, 37(3):623–624, September 1972. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL https://www.jstor.

org/stable/2272794.

DEC:1972:DML

[607] Digital Equipment Corporation. DECsystem-10 mathematical languages handbook: FORTRAN, BASIC, ALGOL. Digital Equipment Corporation, Maynard, MA, USA, second edition, 1972. 388 pp.

Dinkel:1972:BRB

[608] John Dinkel. Book review: Numerical Methods of Mathematical Optimization with ALGOL and FORTRAN Programs (Hans P. Kunzi, H. G. Tzschach and C. A. Zehnder). SIAM Review, 14(2):359–361, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dinkel:1972:BRN

[609] John Dinkel. Book review: Numerical Methods of Mathematical Optimization with ALGOL and FORTRAN Programs (Hans P. Kunzi, H. G.

Tzschach and C. A. Zehnder). SIAM Review, 14(2):359–361, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Earley:1972:MIC

[610] Jay Earley and Paul Caizergues. Method for incrementally compiling languages with nested statement structure. Communications of the ACM, 15(12):1040–1044, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eastman:1972:PRS

[611] Charles M. Eastman. Preliminary report on a system for general space planning. Communications of the ACM, 15(2):76–87, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Emery:1972:APR

[612] J. Emery and J. Vine. Algol programming: Remark on “Procedures for Simplifying Boolean Expressions” [BIT 6(3), 1966, pp. 260–269, by Bo Hellgren]. BIT (Nordisk tidskrift for informationsbehandling), 12(3): 434, September 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0006-3835&volume=12&issue=3&

spage=434. See [401].

Feldman:1972:RDS

[613] Jerome A. Feldman, James R. Low, Daniel C. Swinehart, and Russell H. Taylor. Recent developments in SAIL, an ALGOL-based language for artificial intelligence. Technical Report STAN-CS-72-308 (AIM-176, AD754109), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1972. ?? pp. URL http://i.stanford.

edu/TR/CS-TR-72-308.html.

Fischer:1972:LCS

[614] Michael J. Fischer. Lambda calculus schemata. ACM SIGACT News, 4 (14):104–109, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gill:1972:AIU

[615] P. E. Gill and G. F. Miller. An algorithm for the integration of unequally spaced data. The Computer Journal, 15(1):80–83, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/1/
80.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/150080.sgm.abs.html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/tiff/81.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/83.tif.

Golub:1972:LAS

[616] Gene H. Golub, Richard R. Underwood, and James H. Wilkinson. The Lanczos algorithm for the symmetric Ax = λBx problem. Technical Report STAN-CS-72-270 (PB209616 SU326P30-16), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1972. iii

+ 24 pp. URL http://i.stanford.edu/TR/CS-TR-72-270.html.

Grad:1972:ALS

[617] J. Grad and E. Zakrajsek. LR algorithm with Laguerre shift for symmetric tridiagonal matrices. The Computer Journal, 15(3):268–270, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/268.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/150268.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/268.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_15/
Issue_03/tiff/270.tif.

Hill:1972:DEE

[618] I. D. Hill. Data editing and estimation: The use of Algol in a bronchitis survey. Applied Statistics, 21(1):37–41, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hogben:1972:OIC

[619] David Hogben. OMNITAB II: a computing system for statistical problem solving using Dutch rather than FORTRAN or ALGOL. Statistica Neerlandica, 26(3):85–99, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00176.x.

Hudson:1972:IPS

[620] R. J. W. Hudson. An implementation of phrase structures in ALGOL. Software—Practice and Experience, 2(3):231–240, July/September 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

ISO:1972:PLA

[621] ISO. Programming language: ALGOL: UDC 681.3.04. Ref. no. ISO/ R 1538-1972 (E), volume R 1538 of ISO recommendation. International Organization for Standardization, Geneva, Switzerland, 1972. ISBN ???? 82 pp. LCCN ????

Jones:1972:AMI

[622] H. S. P. Jones. Algorithm 72: Multiple integer arithmetic procedures in ALGOL. The Computer Journal, 15(3):281–282, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/281.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_03/tiff/282.tif.

Kerner:1972:BAS

[623] I. O. Kerner. Bericht über die Algorithmische Sprache ALGOL 68. (German) [Report on the Algorithmic Language ALGOL 68]. Akademie-Verlag, Berlin, Germany, 1972. ???? pp. German translation by Kerner of [443].

Kildall:1972:AEA

[624] Gary A. Kildall and Alan B. Roberts. ALGOL-E: an experimental approach to the study of programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(1):127–135, March 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knuth:1972:ENA

[625] Donald E. Knuth and R. W. Floyd. Erratum: “Notes on avoiding ‘go to’ statements”. Information Processing Letters, 1(4):177, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [558].

Kolbig:1972:CAC

[626] K. S. Kölbig. Certification of “Algorithm 363: Complex error function”. Communications of the ACM, 15(6):465–466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [396].

Krullis:1972:PVL

[627] S. Krullis. Programmierung Einer Verkettungsmatrix Von Literaturzitaten In Algol 60. Semesterarbeit Am Inst. F. Informatik ??, Universität Stuttgart, Stuttgart, West Germany, 1972. ?? pp.

Ledgard:1972:MTC

[628] Henry F. Ledgard. Model for type checking — with an application to Algol 60. Communications of the ACM, 15(11):956–966, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindsey:1972:AFT

[629] Charles H. Lindsey. Algol 68 with fewer tears. The Computer Journal, 15(2):176–188, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/176.full.pdf+html;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/150176.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_02/tiff/176.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/177.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/178.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_15/
Issue_02/tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/tiff/181.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/
182.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/183.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/184.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/185.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_15/
Issue_02/tiff/186.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/187.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/tiff/188.tif.

Malcolm:1972:PRA

[630] Michael A. Malcolm. PL360 (revised again) a programming language for the IBM 360. Technical Report STAN-CS-71-215 (revised again), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972. iv + 103 pp. URL https://bitsavers.org/

pdf/stanford/cs_techReports/STAN-CS-71-215_PL360_Revised_A_
Programming_Language_For_The_IBM_360_May71.pdf. Original ver-

sion May 1971.

Malcolm:1972:PRP

[631] Michael A. Malcolm. PL360 (revised): a programming language for the IBM 360. Technical Report STAN-CS-71-215 (AD727115), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1972.

vii + 103 pp. URL http://i.stanford.edu/TR/CS-TR-71-215.html.

Marshall:1972:IA

[632] Sidney Marshall. On the implementation of ALGOL 68. Ph.D. thesis, School of Engineering, Dartmouth College, Hanover, NH, USA, June

1972. 547 pp. URL https://bitsavers.trailing-edge.com/pdf/
dartmouth/theses/Marshall_-_On_The_Implementation_of_ALGOL_
68_197206.pdf;   https://www.proquest.com/pqdtglobal/docview/
302689344.

Mathur:1972:BDC

[633] F. P. Mathur. A brief description and comparison of programming languages FORTRAN, ALGOL, COBOL, PL/I, and LISP 1.5 from a critical standpoint. Report ??, Jet Propulsion Laboratory, California Inst. of Technology, Pasadena, CA, USA, 1972. 13. pp.

McKay:1972:APCa

[634] A. C. McKay. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(4), 1964, pp. 189–196]. BIT (Nordisk tidskrift for informationsbehandling), 12(2):267–268, June 1972. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=12&issue=2&spage=267. See [176,

248, 635].

McKay:1972:APCb

[635] A. C. McKay. Algol programming: Certification of contribution no. 10: LPMODSIMPLEX [BIT 4(4), 1964, pp. 189–196] with upper bounds. BIT (Nordisk tidskrift for informationsbehandling), 12(2):268–271, 1972. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See [176, 248, 634].

Morris:1972:BVW

[636] James H. Morris, Jr. A bonus from van Wijngaarden’s device. Communications of the ACM, 15(8):773, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oliver:1972:DAC

[637] J. Oliver. A doubly-adaptive Clenshaw–Curtis quadrature method. The Computer Journal, 15(2):141–147, May 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/2/141.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/150141.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_15/Issue_02/tiff/141.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/tiff/142.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/
143.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/145.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/146.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_15/
Issue_02/tiff/147.tif.

Peck:1972:AC

[638] J. E. L. Peck. An ALGOL 68 companion. Technical Report TR-71-01, Department of Computer Science, University of British Columbia, Vancouver, BC, Canada, March 1972. ii + 155 pp. URL https://www.

cs.ubc.ca/tr/1971/tr-71-01.

Petravic:1972:AOS

[639] M. Petravic, G. Kuo-Petravic, and K. V. Roberts. Automatic optimization of symbolic Algol programs. I. General principles. Journal of Computational Physics, 10(3):503–533, December 1972. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900502.

Phillips:1972:SAAa

[640] J. P. N. Phillips. Statistical algorithms: Algorithm AS 48: Uncertainty function for a binary sequence. Applied Statistics, 21(1):97–99, March 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/apstat/48.

Phillips:1972:SAAb

[641] J. P. N. Phillips. Statistical algorithms: Algorithm AS 49: Autocorrelation function for a binary sequence. Applied Statistics, 21(1):100–103, March 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876

(electronic). URL http://lib.stat.cmu.edu/apstat/49.

Rain:1972:SFA

[642] M. Rain. Some formal aspects of MARY or ALGOL X revisited. Report ??, SINTEF, Technical University of Norway, Trondheim, Norway, 1972. ?? pp.

Rogers:1972:USL

[643] William H. Rogers and John W. Tukey. Understanding some long-tailed symmetrical distributions. Statistica Neerlandica, 26(3):211–226, 1972.

CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). Collection of articles in honour of Professor H. C. Hamaker.

Rosen:1972:PSL

[644] Saul Rosen. Programming systems and languages (1965–1975). Communications of the ACM, 15(7):591–600, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sack:1972:AGQ

[645] R. A. Sack and A. F. Donovan. An algorithm for Gaussian quadrature given modified moments. Numerische Mathematik, 18(5):465–478, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saunders:1972:LSL

[646] Michael A. Saunders. Large-scale linear programming using the Cholesky factorization. Technical Report STAN-CS-72-252 (SU326 P30-14), Stanford University, Department of Computer Science, Stanford, CA, USA,

January 1972. 64 pp. URL http://i.stanford.edu/TR/CS-TR-72-
252.html.

Schlechtendahl:1972:CRF

[647] E. G. Schlechtendahl and U. Schumann. Correspondence: Remarks on “Faults in functions, in ALGOL and FORTRAN”. The Computer Journal, 15(2):147, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/150147.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/

147.tif. See [554].

Sites:1972:AWR

[648] Richard L. Sites. Algol W reference manual. Technical Report STAN-CS-71-230 (PB203601), Stanford University, Department of Computer Science, Stanford, CA, USA, February 1972. i + 141

pp. URL http://i.stanford.edu/TR/CS-TR-71-230.html;   https:
//bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-71-230_
Algol_W_Reference_Manual_Feb72.pdf.

Summers:1972:AAF

[649] Victor Summers and Dennis J. Frailey. The Algol and/or Fortran programmer’s guide to Jovial. Technical report CP-72017, Computer Science/Operations Research Center, Institute of Technology, Southern Methodist University, Dallas, TX, USA, 1972. iv + 20 pp.

Terrenoire:1972:EAI

[650] M. Terrenoire and M. Simonet. An evaluation of ALGOL 68 for interrogation process algorithms. International Journal of Computer and Information Sciences, 1(1):67–73, March 1972. CODEN IJCIAH. ISSN 0091-7036.

vanderPoel:1972:CCS

[651] W. L. van der Poel. Comment on the composition of semantics in Algol 68. Communications of the ACM, 15(8):772, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanWijngaarden:1972:BAS

[652] A. van Wijngaarden, Immo Ottomar Kerner, et al., editors. Bericht über die algorithmische Sprache ALGOL 68. (German) [Report on the algorithmic language ALGOL 68]. Elektronisches Rechnen und Regeln, Sonderband l5. Akademie-Verlag, Berlin, Germany, 1972. 381 pp. LCCN QA76.73.A24 R45 1972.

Vasilev:1972:IA

[653] Vladimir Alekseevich Vasil§ev. Yazyk ALGOL-68: Osnovnye Ponyaiya. (Russian) [The ALGOL-68 Language: Basic Concepts]. Izdatel’stvo Nauka, Moskva, USSR, 1972. 128 pp. LCCN QA76.73.A24 V37.

Wichmann:1972:EES

[654] B. A. Wichmann. Estimating the execution speed of an Algol program. ACM SIGPLAN Notices, 7(8):24–44, August 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wichmann:1972:FAC

[655] B. A. Wichmann. Five Algol compilers. The Computer Journal, 15 (1):8–12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/8.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150008.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/
10.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/12.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/8.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/
9.tif.

Wichmann:1972:NII

[656] B. A. Wichmann. A note on the IFIP input-output procedures for Algol 60. Software—Practice and Experience, 2(4):403–404, October 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wirth:1972:PCG

[657] Niklaus Wirth. On “Pascal”, code generation, and the CDC 6000 computer. Technical Report STAN-CS-72-257 (PB208519), Stanford University, Department of Computer Science, Stanford, CA, USA, February

1972. i + 38 pp. URL http://i.stanford.edu/TR/CS-TR-72-257.
html;   http://infolab.stanford.edu/pub/cstr/reports/cs/tr/
72/257/CS-TR-72-257.pdf;   https://bitsavers.org/pdf/stanford/
cs_techReports/STAN-CS-72-257_6600_PASCAL_Feb72.pdf.

Wirth:1972:PLPb

[658] Niklaus Wirth. The programming language PASCAL (revised report). Report 5, ETH Zürich, Zürich, Switzerland, November 1972. viii

+ 48 pp. URL http://pascal.hansotten.com/uploads/wirth/The_
Programming_Language_Pascal_1972.pdf.

Woodward:1972:ARS

[659] P. M. (Philip Mayne) Woodward and Susan Gillian Bond. ALGOL 68-R Users Guide. Her Majesty’s Stationery Office, London, UK, 1972. ISBN 0-11-771422-4. viii + 78 pp. LCCN QA76.73.A24.

Woodward:1972:ARU

[660] Philip M. Woodward and Susan G. Bond. Algol 68-R; users guide. Her Majesty’s Stationery Office, London, UK, 1972. ISBN 0-11-771422-4. viii + 78 pp. LCCN QA76.73.A24 W66.

Woodward:1972:PEA

[661] P. M. Woodward. Practical experience with ALGOL 68. Software—Practice and Experience, 2(1):7–19, January/March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ahmad:1973:CSN

[662] M. Ahmad. The computing speed of a new machine. The Computer Journal, 16(2):161–164, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/2/161.full.pdf+html;

http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/160161.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/tiff/162.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/
163.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/164.tif.

Anonymous:1973:DBA

[663] Anonymous. Differences between Algol 68-R and 68. RRE Internal Document ??, Royal Radar Establishment, Malvern, Worcestershire, UK, June 1973.

Atkins:1973:MRA

[664] M. Stella Atkins. Mutual recursion in Algol 60 using restricted compilers. Communications of the ACM, 16(1):47–48, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bekic:1973:IA

[665] H. Bekić and Olivier Lecarme. An Introduction to Algol 68, volume 7 (Part 3) of Annual review in automatic programming. Pergamon, New York, NY, USA and Oxford, UK, 1973. ISBN 0-08-017128-1. ISSN 0066-4138. v + 143–197 pp. LCCN ????

Benson:1973:AAS

[666] A. Benson and D. J. Evans. Algorithm 80: An algorithm for the solution of periodic quindiagonal systems of linear equations. The Computer Journal, 16(3):278–279, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/278.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/279.tif.

Bron:1973:AAF

[667] Coen Bron and Joep Kerbosch. ACM Algorithm 457: Finding all cliques of an undirected graph. Communications of the ACM, 16(9):575–577, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:1973:AAB

[668] Shi-Kuo Chang. ACM Algorithm 445: Binary pattern reconstruction from projections [Z]. Communications of the ACM, 16(3):185–186, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [686].

Coulson:1973:ASP

[669] V. G. Coulson, I. D. Hill, and V. F. Hillier. Algol survey program. The Computer Journal, 16(1):13–15, February 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/16/1/13.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160013.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/14.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/
15.tif.

Domschke:1973:AAP

[670] W. Domschke. Algorithmen: ALGOL-Prozeduren für zwei Verfahren zur Suche negativer Zyklen in bewerteten Diagraphen. (German) [ALGOLprocedures for two methods for detecting negative cycles in valued digraphs]. Computing, 11(2):169–173, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Domschke:1973:TNA

[671] Wolfgang Domschke. Two new algorithms for minimal cost flow problems. Computing, 11(3):275–285, September 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Erkio:1973:EAV

[672] Hannu Erkio. An extension of ALGOL with variable precision floatingpoint arithmetic. Series A. Report 1973/2, University of Helsinki, Department of Computer Science, Helsinki, 1973. ii + 20 pp.

Gallaher:1973:AAM

[673] L. J. Gallaher. ACM Algorithm 440: A multidimensional Monte Carlo quadrature with adaptive stratified sampling [D1]. Communications of the ACM, 16(1):49–50, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1973:AAE

[674] Walter Gautschi. ACM Algorithm 471: Exponential integrals [S13]. Communications of the ACM, 16(12):761–763, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Griffiths:1973:DOS

[675] J. G. Griffiths. Drawing opaque solids using an incremental plotter. The Computer Journal, 16(1):35–38, February 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/16/1/35.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160035.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/36.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/
37.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/38.tif.

Hansen:1973:RAH

[676] Wilfred J. Hansen. A revised ALGOL 68 hardware representation for ISO-code and EBCDIC. Report ??, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1973. vi + 30 pp.

Haynes:1973:SPS

[677] Leonard S. Haynes. Structure of a Polish String language for an Algol 60 language processor. ACM SIGPLAN Notices, 8(11):131–140, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herriot:1973:APC

[678] J. G. Herriot and C. H. Reinsch. ALGOL 60 procedures for the calculation of interpolating natural quintic spline functions. Technical Report STAN-CS-73-402 (PB229616/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, November 1973. 40 pp.

Hill:1973:AAN

[679] G. W. Hill and A. W. Davis. ACM Algorithm 442: Normal deviate [S14]. Communications of the ACM, 16(1):51–52, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1973:AAS

[680] G. W. Hill. ACM Algorithm 465: Student’s t frequency [S14]. Communications of the ACM, 16(11):690, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hopcroft:1973:AAE

[681] John E. Hopcroft and Robert Endre Tarjan. ACM Algorithm 447: Efficient algorithms for graph manipulation [H]. Communications of the ACM, 16(6):372–378, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hunter:1973:AAA

[682] D. B. Hunter and Julia M. Williams. ACM Algorithm 455: Analysis of skew representations of the symmetric group. Communications of the ACM, 16(9):571–572, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jabs:1973:AFA

[683] Arthur Jabs. Approximation formula and Algol program of the Lorentzinvariant momentum-space integral for particles of equal masses. Computer Physics Communications, 5(3):217–224, March 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/001046557390101X.

Krutar:1973:RGV

[684] R. A. Krutar. Restricted global variables in Algol 60. ACM SIGPLAN Notices, 8(12):15–17, December 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:1973:AAA

[685] C. Lam and J. McKay. ACM Algorithm 469: Arithmetic overa finite field [A1]. Communications of the ACM, 16(11):699, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lau:1973:RAB

[686] John Lau. Remark on “Algorithm 445: Binary Pattern Reconstruction from Projections [Z]”. Communications of the ACM, 16(3):186, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [668].

Liggins:1973:AAC

[687] D. Liggins. Algorithm 79: An algorithm for the capacitated transportation problem. The Computer Journal, 16(3):276–278, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/276.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/277.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/278.tif.

Malcolm:1973:MIA

[688] Michael A. Malcolm. A machine independent ALGOL procedure for accurate floating-point summation. Technical Report STAN-CS-73-374 (AD-764 275), Stanford University, Department of Computer Science, Stanford, CA, USA, June 1973. iii + 6 pp. URL https://apps.dtic.

mil/sti/tr/pdf/AD0764275.pdf.

Mickunas:1973:PGS

[689] M. Dennis Mickunas and Victor B. Schneider. A parser-generating system for constructing compressed compilers. Communications of the ACM, 16(11):669–676, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Padgett:1973:TDD

[690] M. I. Padgett. Tree driven data input and its validation. The Computer Journal, 16(4):315–321, November 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/4/315.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/160315.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/315.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/tiff/316.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/
317.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/319.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/320.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_16/
Issue_04/tiff/321.tif.

Peto:1973:SAA

[691] Richard Peto. Statistical algorithms: Algorithm AS 56: Permutational significance testing. Applied Statistics, 22(1):112–118, March 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/56.

Reinsch:1973:AAE

[692] Christian H. Reinsch. ACM Algorithm 464: Eigenvalues of a real, symmetric, tridiagonal matrix [F2]. Communications of the ACM, 16(11): 689, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rossler:1973:LCC

[693] Roland Rössler. Language-comparison: a comparison of the properties of the programming languages ALGOL 68, CAMAC-IML, CORAL 66, PAS 1, Pearl, PL/I, PROCOL, RTL/2 in relation to real-time programming. ????, Erlangen, West Germany, 1973. xix + 470 pp. LCCN ????

Sethi:1973:NIP

[694] Ravi Sethi. A note on implementing parallel assignment instructions. Information Processing Letters, 2(4):91–95, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spath:1973:ACO

[695] H. Späth. Algorithmen: Clustering of one-dimensional ordered data. Computing, 11(2):175–177, June 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Syslo:1973:AAE

[696] Maciej M. Syslo. ACM Algorithm 459: The elementary circuits of a graph [H]. Communications of the ACM, 16(10):632–633, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [776].

Tinhofer:1973:KAA

[697] G. Tinhofer. Der Out-Of-Kilter-Algorithmus für antisymmetrische Netzwerke. (German) [The out-of-kilter-algorithm for antisymmetric networks]. Computing, 11(1):45–50, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

vandeRiet:1973:ASP

[698] R. P. van de Riet. The automatic solution of partial differential equations by means of Taylor series using formula-manipulation methods. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(28):33–36, December 1973. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

VanLehn:1973:SUM

[699] Kurt A. VanLehn. SAIL user manual. Technical Report STAN-CS-73-373 (AIM-204, AD765353/BWC), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1973. 124 pp. URL http:

//i.stanford.edu/TR/CS-TR-73-373.html.

Vos:1973:RAC

[700] H. Vos. Remark on “Algorithm 300: Coulomb wave functions”. Communications of the ACM, 16(5):308–309, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [299].

Watkins:1973:RAG

[701] Richard P. Watkins. Remark on “Algorithm 412: Graph plotter”. Communications of the ACM, 16(8):489–490, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [538].

Wegner:1973:BPG

[702] E. Wegner. Beamstrukturierte Programme. (German) [Beam structured programs]. Report PC2, D83, Informatik Technische Universtät Berlin, Berlin, West Germany, 1973. ?? pp.

Wirth:1973:PLP

[703] Niklaus Wirth. The programming language PASCAL, (revised report). Report 5, ETH Zürich, Zürich, Switzerland, July 1973. iv

+ 49 pp. URL http://pascal.hansotten.com/uploads/wirth/The_
Programming_Language_Pascal_1973.pdf;   https://www.research-
collection.ethz.ch/handle/20.500.11850/68910.

Balfour:1974:AAM

[704] A. Balfour. An alternative approach to mutual recursion in Algol 60 using restricted compilers. Communications of the ACM, 17(4):210, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barron:1974:BRBa

[705] D. W. Barron. Book review: Informal introduction to ALGOL 68, C. H. Lindsey and S. G. van der Meulen. North-Holland, Amsterdam, 1971. No. of pages: 382. Price: £7.00. Also available in paperback. Software—Practice and Experience, 4(3):304, July/September 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Batson:1974:MVM

[706] A. P. Batson and R. E. Brundage. Measurements of the virtual memory demands of Algol-60 programs (extended abstract). ACM SIGMETRICS Performance Evaluation Review, 3(4):121–126, December 1974. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Broderick:1974:BRB

[707] W. R. Broderick. Book reviews: Spline Analysis, by Martin H. Schultz, 1972; Introduction to Programming, by A. Sharp; Algol Programming, by R. W. Sharp. The Computer Journal, 17(1):92, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/92.tif.

Broderick:1974:BRS

[708] W. R. Broderick. Book reviews: Spline Analysis, by Martin H. Schultz, 1972; Introduction to Programming, by A. Sharp; Algol Programming, by R. W. Sharp. The Computer Journal, 17(1):92, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/92.tif.

Bron:1974:RAM

[709] C. Bron. Remark on “Algorithm 426: Merge sort algorithm”. Communications of the ACM, 17(12):706–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [602].

Brown:1974:WSA

[710] P. J. Brown. Writing software in ALGOL. Software—Practice and Experience, 4(2):139–144, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Buchanan:1974:ACP

[711] Jack R. Buchanan and David C. Luckham. On automating the construction of programs. Technical Report STAN-CS-74-433 (AIM-236, AD784513), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. 65 pp. URL http://i.stanford.edu/

TR/CS-TR-74-433.html.

Buchanan:1974:SAP

[712] Jack R. Buchanan. A study in automatic programming. Technical Report STAN-CS-74-458 (AIM-245, AD784816), Stanford University, Department of Computer Science, Stanford, CA, USA, May 1974. vii + 148 pp.

Cohen:1974:NTD

[713] J. Cohen, L. Trilling, and P. Wegner. A nucleus of a theorem-prover described in ALGOL-68. International Journal of Computer and Information Sciences, 3(1):1–31, March 1974. CODEN IJCIAH. ISSN 0091-7036.

Cohen:1974:TLE

[714] Jacques Cohen and Carl Zuckerman. Two languages for estimating program efficiency. Communications of the ACM, 17(6):301–308, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Finkel:1974:APS

[715] Raphael A. Finkel, Russell H. Taylor, Robert C. Bolles, Richard P. Paul, and Jerome A. Feldman. AL, a programming system for automation. Technical Report STAN-CS-74-456 (AIM-243, AD-A003815), Stanford University, Department of Computer Science, Stanford, CA, USA,

November 1974. ?? pp. URL http://i.stanford.edu/TR/CS-TR-74-
456.html.

Freeman:1974:BFA

[716] P. Freeman. B 74-34 Fortran and Algol: a programmed course for students of science and technology. IEEE Transactions on Computers, C-23(5):559–560, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=1672583;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&arnumber=1672583.

Germano:1974:LAC

[717] G. Germano and A. Maggiolo-Schettini. Loops in Algol 60 and in category theory. ACM SIGACT News, 6(4):21–23, October 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gottler:1974:AEE

[718] Herbert Göttler. Algol 68: eine exemplarische Einführung. (German) [Algol 68: a walkthrough introduction]. Arbeitsberichte des Instituts für Mathematische Maschinen und Datenverarbeitung; Bd. 7, Nr. 5. Institut für Mathematische Maschinen und Datenverarbeitung, Erlangen, West Germany, 1974. 111 pp. LCCN QA76.73.A24 G63.

Greibach:1974:SRW

[719] S. A. Greibach. Some restrictions on W-grammars. International Journal of Computer and Information Sciences, 3(4):289–327, December 1974. CODEN IJCIAH. ISSN 0091-7036.

Hansen:1974:THL

[720] Wilfred J. Hansen. Transportation of higher-level language programs: exemplified by an ALGOL 68 transportation representation. Report ??, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1974. iv + 39 pp.

Herriot:1974:APC

[721] John G. Herriot and Christian H. Reinsch. Algol 60 procedures for the calculation of interpolating natural quintic spline functions. Technical Report STAN-CS-74-402 (PB229616/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, January 1974. iv + 40 pp.

Hopcroft:1974:EPT

[722] John Hopcroft and Robert Tarjan. Efficient planarity testing. Journal of the ACM, 21(4):549–568, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See comments and errata in [795].

Hopgood:1974:BRB

[723] F. R. A. Hopgood. Book review: ALGOL 60 compilation and assessment, B. A. Wichmann, Academic Press, London, 1973. No. of pages:

320. Price: £6.50. Software—Practice and Experience, 4(4):405–406, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Koppelaar:1974:CRA

[724] Henk Koppelaar. Certification and remark on “Algorithm 191: Hypergeometric”. Communications of the ACM, 17(10):589–590, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [140].

Learner:1974:IAT

[725] A. Learner and Anthony J. Powell. An Introduction to Algol 68 through Problems. Macmillan Publishing Company, New York, NY, USA, 1974. ISBN 0-333-16620-5 (paperback), 1-349-02226-8, 1-349-02227-6, 1-349-02228-4. iv + 124 pp. LCCN QA76.73.A24.

Lyche:1974:AAP

[726] Tom Lyche and Larry L. Schumaker. ACM Algorithm 480: Procedures for computing smoothing and interpolating natural splines [E1]. Communications of the ACM, 17(8):463–467, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKay:1974:AAT

[727] John McKay and E. Regener. ACM Algorithm 482: Transitivity sets [G7]. Communications of the ACM, 17(8):470, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meulenvander:1974:PA

[728] S. G. Meulen, van der and Peter Kühling. Programmieren in ALGOL 68. (German) [Programming in ALGOL 68]. de Gruyter, New York, NY, USA and Berlin, West Germany, 1974. ???? pp. LCCN QA76.73.A24 M48.

Palme:1974:STE

[729] Jacob Palme. SIMULA as a tool for extensible program products. ACM SIGPLAN Notices, 9(2):24–40, February 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pasedach:1974:VSA

[730] K. Pasedach. Ein Verfahren zur schnellen Approximation von Summen mit vielen Summanden. (German) [Method for the fast approximation of sums with many terms]. Computing, 12(4):323–331, December 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Peck:1974:RAS

[731] J. E. L. Peck, J. M. Watt, and Michel Sintzoff. Revised ALGOL 68 syntax chart. ACM SIGPLAN Notices, 9(7):39, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rothmaier:1974:TCU

[732] B. Rothmaier. Der Triplex-Algol Compiler Der Univac 1108. Interner Bericht Des Inst. F. Prakt. Math. 74/1, Universität Karlsruhe, Karlsruhe, West Germany, 1974. ?? pp.

Stiller:1974:ABU

[733] Gerd Stiller. ALGOL 68, Begriffe und Ausdrucksmittel. (German) [AL-GOL 68, terms and means of expression]. Schriftenreihe Informationsverarbeitung. BSB Teubner, Leipzig, East Germany, 1974. 164 pp. LCCN QA76.73.A24 S74.

Swinehart:1974:CMP

[734] Daniel Carl Swinehart. COPILOT: a multiple process approach to interactive programming systems. Technical Report STAN-CS-74-412 (AIM-230, AD786721), Stanford University, Department of Computer Science, Stanford, CA, USA, July 1974. ?? pp. URL https://apps.dtic.mil/

sti/tr/pdf/AD0786721.pdf.

Ulery:1974:SLO

[735] Dana L. Ulery and H. M. Khalil. A survey of language-oriented systems for numerical linear algebra. The Computer Journal, 17(1):82–88, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/1/
82.full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_17/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/83.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/
84.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/87.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/
88.tif.

Ulery:1974:SLS

[736] Dana L. Ulery and H. M. Khalil. A survey of language-oriented systems for numerical linear algebra. The Computer Journal, 17(1):82–88, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/83.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/84.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/85.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/88.tif.

Valentine:1974:CNA

[737] S. H. Valentine. Comparative notes on Algol 68 and PL/I. The Computer Journal, 17(4):325–331, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/4/325.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/325.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/326.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_04/tiff/327.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/tiff/328.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_17/
Issue_04/tiff/329.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/330.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/331.tif.

Vowels:1974:AFI

[738] R. A. Vowels. ALGOL 60 and FORTRAN IV. J. Wiley and Sons Australasia, Sydney, Australia; New York, NY, USA, 1974. ISBN 0-471-91192-5. 173 pp. LCCN QA76.73.A24 V68.

Woodward:1974:ARU

[739] Philip M. Woodward and Susan G. Bond. Algol 68-R Users Guide. Her Majesty’s Stationery Office, London, UK, second edition, 1974. ISBN 0-11-771600-6 (paperback). viii + 99 pp. LCCN QA76.73.A24 W66 1974.

Woodward:1974:BRB

[740] P. M. Woodward. Book review: ALGOL 68 implementation, Ed. J. E. L. Peck, proceedings of the IPIP working conference on ALGOL 68 implementation, Munich, 20–24 July 1970, North-Holland, Amsterdam, 1971. No. of pages: 375. Price: £7.35. Software—Practice and Experience, 4 (2):192, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woolley:1974:LSL

[741] John D. Woolley and Leland R. Miller. LINUS: a structured language for instructional use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1):125–128, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Wright:1974:F

[742] F. H. G. Wright, II and Ralph E. Gorin. FAIL. Technical Report STAN-CS-74-407 (AIM-226, AD778310), Stanford University, Department of Computer Science, Stanford, CA, USA, April 1974. ?? pp. URL http:

//i.stanford.edu/TR/CS-TR-74-407.html.

Abdali:1975:LMPa

[743] S. K. Abdali. A lambda-calculus model of programming languages. I. simple constructs. Computer Languages, 1(4):287–301, ???? 1975. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

AFCETGA:1975:MDL

[744] Association française de cybernétique économique et technique.Groupe Algol. Manuel du langage algorithmique, Algol 68, volume 1373 of Actualités scientifiques et industrielles. Hermann, Paris, France, 1975. ISBN 2-7056-1373-0. 492 pp. LCCN Q111 .A3 no. 1373 QA76.73.A24.

Ammeraal:1975:DPLd

[745] Leendert Ammeraal. On the design of programming languages including mini ALGOL 68. Report IW 35/75, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Nether-

lands, 1975. 6 pp. URL https://ir.cwi.nl/pub/9464.

Anonymous:1975:FCS

[746] Anonymous. Foundations of computer science: [an advanced course on the foundations of computer science organized by the Mathematical Centre, held at the University of Amsterdam, May 20-31, 1974]. C, January 1975. ?? pp.

Anonymous:1975:ICA

[747] Anonymous. International Conference on Algol 68 Oklahoma State University. American Mathematical Monthly, 82(5):553, May 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baecker:1975:AR

[748] H. D. Baecker. Areas and record-classes. The Computer Journal, 18(3): 223–226, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_03/tiff/223.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_18/
Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/226.tif.

Baecker:1975:ARC

[749] H. D. Baecker. Areas and record-classes. The Computer Journal, 18(3): 223–226, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/223.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_03/tiff/223.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_18/
Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/226.tif.

Barron:1975:BRBa

[750] D. W. Barron. Book review: Algol 68-R Users Guide (2nd edn), P. M. Woodward and S. G. Bond, H.M.S.O., 1974. No. of pages: 99. Price: £0.90. Software—Practice and Experience, 5(2):217, April 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bierbaum:1975:EIN

[751] F. Bierbaum. Einsatz Der Intervallarithmetik Bei Der Numerischen Konvergenz Von Algol-60 Programmen. In K. Nickel, editor, Interval Mathematics, volume 29 of Lecture Notes In Computer Science, pages 160– 168. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975.

Brundage:1975:CPD

[752] Robert E. Brundage and Alan P. Batson. Computational processor demands of Algol-60 programs. Operating Systems Review, 9(5):161–168, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bukac:1975:SAA

[753] J. Bukač. Statistical algorithms: Algorithm AS 85: Critical values of the sign test. Applied Statistics, 24(2):265–267, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/85.

Burkard:1975:AVG

[754] Rainer E. Burkard and Herbert Enge. Algorithmus 25. Verfahren zur Gemischt-Ganzzahligen, konvexen Optimierung. (German) [Method for mixed integer convex programming]. Computing, 14(4):389–396, December 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Busch:1975:EAM

[755] W. Busch, R. Esser, W. Hackbusch, and U. Herrmann. Extrapolation applied to the method of characteristics for a first order system of two partial differential equations. I. the initial value problem. Numerische Mathematik, 24(4):331–353, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

CDC:1975:AVR

[756] CDC. ALGOL 68 Version 1 Reference Manual. Control Data Services B.V., Rijswijk, The Netherlands, b (1.1) edition, 1975. 115 pp. URL

https://archive.org/details/bitsavers_cdcTomHunton1ReferenceManualRevB_
8756301.

Denert:1975:GGL

[757] E. Denert, G. Ernst, and H. Wetzel. GRAPHEX68: Graphical language features in ALGOL 68. Computers and Graphics, 1(2–3):195–202, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Flon:1975:RSP

[758] Lawrence Flon. On research in structured programming. ACM SIG-PLAN Notices, 10(10):16–17, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ford:1975:APC

[759] W. S. Ford. Algorithm 90: Periodic cubic spline interpolation with equidistant nodes. The Computer Journal, 18(2):183–184, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/183.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/184.tif.

Griss:1975:RSC

[760] Martin L. Griss. The REDUCE system for computer algebra. In Proc. ACM 75, pages 261–262. ACM Press, New York, NY 10036, USA, 1975.

Hext:1975:ARO

[761] J. B. Hext. Array reference operations. ACM SIGPLAN Notices, 10 (3):113–118, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1975:AEM

[762] I. D. Hill. Algorithm 88: An electoral method. The Computer Journal, 18(1):89–90, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/90.tif.

Hill:1975:WAAa

[763] I. D. Hill, R. S. Scowen, and B. A. Wichmann. Writing algorithms in ALGOL 60. ACM SIGPLAN Notices, 10(2):14–31, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1975:WAAb

[764] I. D. Hill, R. S. Scowen, and B. A. Wichmann. Writing algorithms in ALGOL 60. Software—Practice and Experience, 5(3):229–244, July/ September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Housden:1975:SCT

[765] R. J. W. Housden. On string concepts and their implementation. The Computer Journal, 18(2):150–156, May 1975. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/2/150.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/150.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/151.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_02/tiff/152.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/tiff/153.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_18/
Issue_02/tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/155.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/156.tif.

HughesJones:1975:AGC

[766] R. Hughes Jones, P. R. Graves-Morris, and G. J. Makinson. Algorithm 85: Generation of Chisholm approximants. The Computer Journal, 18 (1):81–83, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/82.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/83.tif.

Keir:1975:CNR

[767] Roy A. Keir. Compatible number representations. In IEEE SCA ’75 [1279], pages 82–87. LCCN QA76.6.S919 1975. URL http://www.acsel-

lab.com/arithmetic/arith3/papers/ARITH3_Keir_2.pdf. IEEE or-

der number CH1017-3C.

Koch:1975:AAM

[768] Wilfried Koch and Christoph Oeters. An abstract ALGOL 68 machine and its application in a machine independent compiler. In GI — 5. Jahrestagung, volume 34 of Lecture Notes in Computer Science, pages 642–653. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 3-540-37929-0. ISSN 1611-3349.

Lawrie:1975:GPL

[769] Duncan H. Lawrie, T. Layman, D. Baer, and J. M. Randal. GLYPNIR — A programming language for Illiac IV. Communications of the ACM, 18(3):157–164, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leibowitz:1975:MOU

[770] Stephen Leibowitz. Memory optimization using block structures. ACM SIGPLAN Notices, 10(11):22–25, November 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1975:CPF

[771] John Gregg Lewis. Certification of “Algorithm 349: Polygamma functions with arbitrary precision”. ACM Transactions on Mathematical Software, 1(4):380–382, December 1975. CODEN ACMSCU. ISSN 0098-3500. See [440].

Mardia:1975:SAAa

[772] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 80: Spherical statistics. Applied Statistics, 24(1):144–146, March 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/80.

Mardia:1975:SAAd

[773] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 86: The von Mises distribution function. Applied Statistics, 24(2):268– 272, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876

(electronic). URL http://lib.stat.cmu.edu/apstat/86.

Reiser:1975:BDS

[774] John F. Reiser. BAIL: a debugger for SAIL. Technical Report STAN-CS-75-523 (AIM-270, AD-A019467), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1975. ?? pp. URL

http://i.stanford.edu/TR/CS-TR-75-523.html.

Shearn:1975:DES

[775] D. C. S. Shearn. Discrete event simulation in ALGOL 68. Software— Practice and Experience, 5(3):279–293, July/September 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Syslo:1975:RAE

[776] Maciej M. Syslo. Remark on “Algorithm 459: The elementary circuits of a graph”. Communications of the ACM, 18(2):119–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [696].

VanEmdeBoas:1975:NRE

[777] Peter Van Emde Boas and Paul M. B. Vitányi. A note on the recursive enumerability of some classes of recursively enumerable languages.

Report IW 35/75, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1975. 8 pp. URL

https://ir.cwi.nl/pub/9462/9462D.pdf.

vanWijngaarden:1975:RRA

[778] Adriaan van Wijngaarden, B. J. Mailloux, J. E. L. Peck, C. H. A. Koster, M. Sintzoff, C. H. Lindsey, L. G. L. T. Meertens, and R. G. Fisker. Revised report on the algorithmic language ALGOL 68. Acta Informatica, 5(1–3):1–236, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ziegler:1975:IIA

[779] B. Ziegler. Inversion einer im allgemeinen nicht positiv definiten symmetrischen Matrix. (German) [Inversion of a matrix which is symmetric but in general not positive definite]. Computing, 14(1–2):131–139, March 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zosel:1975:MPV

[780] Mary E. Zosel. A modest proposal for vector extensions to ALGOL. ACM SIGPLAN Notices, 10(3):62–71, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zvegintzov:1975:PAD

[781] N. Zvegintzov. Problems with analytic description of text formats. The Computer Journal, 18(2):148–149, May 1975. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/2/148.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/148.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/149.tif.

Ammann:1976:CGP

[782] Urs Ammann. On code generation in a PASCAL compiler. Report 13, ETH Zürich, Zürich, Switzerland, 1976. i + 41 pp. URL http://pascal.

hansotten.com/uploads/wirth/Pascal-On-Code-Generation-in-a-
Pascal-Compiler-eth-3056-01.pdf.

Anonymous:1976:EAJ

[783] Anonymous. Evaluation of ALGOL 68, Jovial J3B, PASCAL, SIMULA 67, and TACPOL versus TINMAN requirements for a common high order programming language. Report ??, Softech, Inc. and Defense Technical Information Center, Waltham, MA, USA and Fort Belvoir, VA, USA, October 1976. ?? pp.

Anonymous:1976:LCC

[784] Anonymous. A language comparison. a comparison of the properties of the programming languages ALGOL 68, CAMAC-IML, Coral 66, PAS 1, Pearl, PL/1, PROCOL, RTL?? Report ??, Purdue Laboratory for Applied Industrial Control, Purdue University, Lafayette, IN, USA, October 1976. ?? pp.

Bachmann:1976:PPA

[785] Karl Heinz Bachmann. Die Programmierungssprachen Pascal und AL-GOL 68. (German) [The programming languages Pascal and ALGOL 68], volume 10 of Elektronisches Rechnen und Regeln. Akademie-Verlag, Berlin, Germany, 1976. 220 pp. LCCN QA76.73.P2 B3.

Backus:1976:MRA

[786] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-

60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/4/364.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/364.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/365.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/366.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/367.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/369.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/370.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/
371.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/372.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_04/tiff/373.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/374.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/376.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/377.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/tiff/

378.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_04/tiff/379.tif.

Baron:1976:TI

[787] W. Baron. Zur trigonometrischen Interpolation. (German) [On trigonometric interpolation]. Computing, 16(4):319–328, December 1976. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Batson:1976:DDA

[788] A. P. Batson, R. E. Brundage, and J. P. Kearns. Design data for Algol-60 machines. ACM SIGARCH Computer Architecture News, 4(4):151–154, January 1976. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Boas:1976:LFP

[789] Peter van Emde Boas. Least fixed points and the recursion theorem. Dummy note., January 1976.

Boas:1976:SAM

[790] Peter van Emde Boas. Some applications of the Meyer–McCreight algorithm in abstract complexity theory: (prepublication). Dummy note., January 1976.

Brailsford:1976:ADC

[791] D. F. Brailsford and R. D. Knott. ALGOL 68 dialect conversion — a syntactic approach. In ????, editor, Proceedings of a conference on “Applications of Algol 68”, University of East Anglia, 22– 25 March 1976, pages 192–203. ????, ????, 1976. URL https://

www.researchgate.net/publication/28692800_Algol_68_Dialect_
Conversion_--_a_Syntactic_Approach.

Branquart:1976:OTP

[792] Paul Branquart. An Optimized translation process and its application to ALGOL 68, volume 38 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNCSD9. ISBN 0-387-07545-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 334 pp. LCCN QA76.73.A24

O67. URL http://link.springer-ny.com/link/service/series/
0558/tocs/t0038.htm;   http://www.springerlink.com/content/
978-0-387-07545-7;   http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=38.

Brewer:1976:DSB

[793] Richard K. Brewer. Documentation standards for beginning students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):69–73, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

DeMorgan:1976:SAR

[794] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A supplement to the Algol 60 revised report. The Computer Journal, 19(3):276–288, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/276.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_03/tiff/277.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/tiff/278.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_03/tiff/279.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/280.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/281.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/tiff/
282.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_03/tiff/284.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/tiff/285.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_19/
Issue_03/tiff/286.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/287.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/288.tif.

Deo:1976:NHT

[795] Narsingh Deo. Note on Hopcroft and Tarjan’s planarity algorithm. Journal of the ACM, 23(1):74–75, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [722].

Earnshaw:1976:GPA

[796] R. A. Earnshaw. Graph plotting in ALGOL 68-r. Software—Practice and Experience, 6(1):51–60, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Eijgenraam:1976:TIC

[797] P. Eijgenraam, J. A. Van De Griend, and L. S. C. Statema. Triplex-Algol 60 on the IBM 370/158 of the Centraal Rekeninstituut. Report 76-1, Institute of Applied Math. and Computer Science, University of Leiden, Leiden, The Netherlands, 1976. ?? pp.

elLozy:1976:RAC

[798] Mohamed el Lozy. Remark on “Algorithm 299: Chi-Squared Integral [S15]”. ACM Transactions on Mathematical Software, 2(4):393–395, December 1976. CODEN ACMSCU. ISSN 0098-3500. See [301, 1081].

Ford:1976:DSN

[799] B. Ford and D. K. Sayers. Developing a single numerical algorithm library for different machine ranges. ACM Transactions on Mathematical Software, 2:115–131, 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gegusch:1976:DTF

[800] T. Gegusch. Eine Dezimale Triplex-Langzahl-Arithmetik Für Den Univac-Algol-60-Compiler. Diplomarbeit, Inst. F. Prakt. Math., University Karlsruhe, 1976. ?? pp.

Goodenough:1976:EAJ

[801] John B. Goodenough, Clement L. McGowan, and John R. Kelly. Evaluation of ALGOL 68, JOVIAL J3B, PASCAL, SIMULA 67, and TACPOL vs Tinman requirements for a common high order programming language — final report for period: July 1976–October 1976. Report ??, US Army Electronics Command and Defense Technical Information Center, Fort Monmouth, NJ, USA and Fort Belvoir, VA, USA, 1976. 165 pp.

Haskell:1976:USF

[802] R. E. Haskell, D. E. Boddy, and G. A. Jackson. Use of structured flowcharts in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):67–74, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Herriot:1976:CIN

[803] John G. Herriot. Calculation of interpolating natural spline functions using de Boor’s package for calculating with B-splines. Technical Report STAN-CS-76-569 (P261 814/AS), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. i + 44 pp. URL

http://i.stanford.edu/TR/CS-TR-76-569.html.

Jones:1976:EASa

[804] B. Jones. An extended Algol-60 for shaded computer graphics. Computer Graphics, 10(1):18–23, Spring 1976. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Jones:1976:EASb

[805] Ben Jones. An extended Algol-60 for Shaded Computer Graphics. ACM SIGPLAN Notices, 11(6):18–23, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knott:1976:NSP

[806] Gary D. Knott. A numbering system for permutations of combinations. Communications of the ACM, 19(6):355–356, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuo-Petravic:1976:TPT

[807] L. G. Kuo-Petravic, M. Petravic, and K. V. Roberts. Tranal — a program for the translation of Symbolic Algol into Symbolic Algol II. Computer Physics Communications, 11(1):5–25, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0010465576900369.

Madison:1976:CPL

[808] A. Wayne Madison and Alan P. Batson. Characteristics of program localities. Communications of the ACM, 19(5):285–294, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Matuszek:1976:CAS

[809] David Matuszek. The case for the assert statement. ACM SIGPLAN Notices, 11(8):36–37, August 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meertens:1976:SST

[810] L. Meertens. A space-saving technique for assigning ALGOL 68 multiple values. Information Processing Letters, 5(4):97–99, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pagan:1976:PGA

[811] Frank G. Pagan. A Practical Guide to Algol 68. Wiley series in computing. John Wiley, London, UK, 1976. ISBN 0-471-65746-8 (hardcover), 0-471-65747-6 (paperback). ISSN 0277-2647. ix + 213 pp. LCCN QA76.73.A24 P34.

Reiser:1976:S

[812] John F. Reiser. SAIL. Technical Report STAN-CS-76-574 (AIM-289), Stanford University, Department of Computer Science, Stanford, CA,

USA, August 1976. 178 pp. URL http://i.stanford.edu/pub/cstr/
reports/cs/tr/76/574/CS-TR-76-574.pdf;   http://i.stanford.
edu/TR/CS-TR-76-574.html.

Riha:1976:AEA

[813] W. Riha and K. R. James. Algorithm 29. efficient algorithms for doubly and multiply restricted partitions. Computing, 16(1–2):163–168, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Smith:1976:SLP

[814] Carol Smith and Jon Rickman. Selecting languages for pedagogical tools in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):39–47, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Smith:1976:ST

[815] Nancy W. Smith. SAIL tutorial. Technical Report STAN-CS-76-575 (AIM-290, AD-A042494), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1976. 54 pp. URL http://i.

stanford.edu/TR/CS-TR-76-575.html.

Tanenbaum:1976:TA

[816] Andrew S. Tanenbaum. A tutorial on Algol 68. ACM Computing Surveys, 8(2):155–190, June 1976. CODEN CMSVAN. ISSN 0010-4892. See also [881].

vanWijngaarden:1976:RRAa

[817] Adriaan van Wijngaarden, B. J. (Barry James) Mailloux, J. E. L. (John E. L. (Edward Lancelot)) Peck, C. H. A. (Cornelis H. A.) Koster, M. Sintzoff, C. H. Lindsey, Lambert G. L. Meertens, and R. G. Fisker, editors. Revised Report on the Algorithmic Language ALGOL 68. Number 50 in Mathematical Centre tracts; 50. Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1976. ISBN 90-6196-089-4. 236 pp. LCCN QA76.73.A24 R45 1976b. URL

https://ir.cwi.nl/pub/18575.

vanWijngaarden:1976:RRAb

[818] Adriaan van Wijngaarden, B. J. (Barry James) Mailloux, J. E. L. (John E. L. (Edward Lancelot)) Peck, C. H. A. (Cornelis H. A.) Koster, M. Sintzoff, C. H. Lindsey, Lambert G. L. Meertens, and R. G. Fisker, editors. Revised Report on the Algorithmic Language ALGOL 68. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07592-5, 3-540-07592-5. 236 pp. LCCN QA76.73.A24 R45 1976.

URL https://algol68genie.nl/en/blog/revised-report-algol-
68-html;   https://web.archive.org/web/20070927191700/http:
//burks.brighton.ac.uk/burks/language/other/a68rr/rrtoc.htm.

vonBochmann:1976:SEL

[819] Gregor von Bochmann. Semantic evaluation from left to right. Communications of the ACM, 19(2):55–62, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wichmann:1976:TAC

[820] B. A. Wichmann and B. Jones. Testing ALGOL 60 compilers. Software—Practice and Experience, 6(2):261–270, April/June 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Winiger:1976:NOP

[821] G. Winiger. A note on one-pass CASE statement compilation. ACM SIG-PLAN Notices, 11(1):32–36, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bajpai:1977:FA

[822] Avi C. Bajpai. Fortran y Algol. Editorial Limusa, Mexico, DF, Mexico, 1977. 265 pp.

Banatre:1977:POC

[823] J. P. Banatre. Producing optimised code for coercions. Information Processing Letters, 6(2):56–59, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Batson:1977:SSL

[824] Alan P. Batson and Robert E. Brundage. Segment sizes and lifetimes in Algol 60 programs. Communications of the ACM, 20(1):36–44, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bennett:1977:IPI

[825] M. W. Bennett. Implementation of a PDP11/ICL1900 cross assembler in Algol 68R. ACM SIGPLAN Notices, 12(6):153–156, June 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birrell:1977:SMA

[826] Andrew D. Birrell. Storage management for ALGOL 68. ACM SIGPLAN Notices, 12(6):82–94, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Braid:1977:GMA

[827] I. C. Braid and R. C. Hillyard. Geometric modelling in ALGOL 68. ACM SIGPLAN Notices, 12(6):168–174, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brailsford:1977:RTPa

[828] D. F. Brailsford, E. Foxley, K. C. Mander, and D. J. Morgan. Run-time profiling of Algol 68-R programs using DIDYMUS and SCAMP. ACM SIGPLAN Notices, 12(6):27–35, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brownlee:1977:ABI

[829] J. Nevil Brownlee. An Algol-based implementation of SNOBOL 4 patterns. Communications of the ACM, 20(7):527–529, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Colin:1977:PPS

[830] Andrew John Theodore Colin. Programming and problem-solving in Algol 68. Macmillan computer science series. Macmillan Publishing Company, New York, NY, USA, 1977. ISBN 0-333-21716-0 (paperback); 0-333-23115-5 (paperback). viii + 251 pp. LCCN QA76.73.A24 C64.

Cousot:1977:SDDa

[831] Patrick Cousot and Radhia Cousot. Static determination of dynamic properties of generalized type unions. ACM SIGPLAN Notices, 12(3): 77–94, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:1977:SDDb

[832] Patrick Cousot and Radhia Cousot. Static determination of dynamic properties of generalized type unions. ACM SIGSOFT Software Engineering Notes, 2(2):77–94, March 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Craven:1977:DTA

[833] P. G. Craven. Derivatives without tears in Algol 68. ACM SIGPLAN Notices, 12(6):19–26, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Day:1977:AGP

[834] A. C. Day. Algorithm 95: Generating a parity testing table. The Computer Journal, 20(3):280, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_20/Issue_03/tiff/280.tif.

Delves:1977:AFG

[835] L. M. Delves and L. F. Abd.El Al. Algorithm 97: The fast Galerkin algorithm for the solution of linear Fredholm equations. The Computer Journal, 20(4):374–376, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/374.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/375.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/376.tif.

Delves:1977:ALE

[836] L. M. Delves. Algorithm 96: a linear equation solver for Galerkin and least squares methods. The Computer Journal, 20(4):371– 374, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_04/tiff/371.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_20/
Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/374.tif.

Delves:1977:ALN

[837] L. M. Delves. Algol 68 as a language for numerical software. In Cowell [1289], pages 95–126. ISBN 0-387-08446-0, 3-540-08446-0, 3-540-37071-

4. ISSN 0302-9743. LCCN QA297 .W65 1976. URL http://link.

springer.com/10.1007/3-540-08446-0.

Delves:1977:LES

[838] L. M. Delves. Algorithm 96: A linear equation solver for Galerkin and least squares methods. The Computer Journal, 20(4):371– 374, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_04/tiff/371.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_20/
Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/374.tif.

DeMorgan:1977:SAR

[839] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A supplement to the ALGOL 60 revised report. ACM SIGPLAN Notices, 12(1):52–66, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dershowitz:1977:IRP

[840] Nachum Dershowitz and Zohar Manna. Inference rules for program annotation. Technical Report STAN-CS-77-631 (AIM-303, AD-A050806), Stanford University, Department of Computer Science, Stanford, CA, USA, October 1977. 46 pp. URL http://i.

stanford.edu/TR/CS-TR-77-631.html;   https://apps.dtic.mil/
sti/tr/pdf/ADA050806.pdf.

Donahue:1977:LCU

[841] James E. Donahue. Locations considered unnecessary. Acta Informatica, 8(3):221–242, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Edwards:1977:PLS

[842] Roy Edwards. Is Pascal a logical subset of Algol 68 or not? Part One. ACM SIGPLAN Notices, 12(6):184–191, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Evans:1977:ACB

[843] D. J. Evans and S. A. Smith. Algorithm 99: On the construction of balanced binary trees for parallel processing. The Computer Journal, 20(4): 378–379, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_04/tiff/378.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/379.
tif.

Foster:1977:ADF

[844] J. M. Foster and P. D. Foster. Abstract data and functors. ACM SIG-PLAN Notices, 12(6):161–167, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gardner:1977:TA

[845] P. J. Gardner. A transportation of ALGOL 68C. ACM SIGPLAN Notices, 12(6):95–101, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1977:TDS

[846] Dick Grune. Towards the design of a super-language of ALGOL 68 for the Standard Prelude (excerpt). ACM SIGPLAN Notices, 12(6):78–81, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamlet:1977:IAC

[847] Richard Hamlet. Ignorance of ALGOL 68 considered harmful. ACM SIGPLAN Notices, 12(4):51–56, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1977:RSH

[848] Wilfrid J. Hansen and Hendrik Boom. The report on the standard hardware representation for ALGOL 68. ACM SIGPLAN Notices, 12(5): 80–87, May 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haubitz:1977:PZA

[849] I. Haubitz. Programme zum Zeichnen von allgemeinen Flächenstücken. (German) [Programs drawing pieces of surfaces]. Computing, 18(4):295– 315, December 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haynes:1977:AAC

[850] Leonard S. Haynes. The architecture of an ALGOL 60 computer implemented with distributed processors. ACM SIGARCH Computer Architecture News, 5(7):95–104, March 1977. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hedrick:1977:AIO

[851] G. E. Hedrick. ALGOL68 instruction at Oklahoma State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):16–20, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Hemker:1977:CTA

[852] Pieter W. Hemker. Criteria for transportable Algol libraries. In Cowell [1289], pages 145–157. ISBN 0-387-08446-0, 3-540-08446-0, 3-540-37071-4. ISSN 0302-9743. LCCN QA297 .W65 1976. URL http://link.

springer.com/10.1007/3-540-08446-0.

Hennell:1977:QTE

[853] M. A. Hennell, D. Hedley, and M. R. Woodward. Quantifying the test effectiveness of Algol 68 programs. ACM SIGPLAN Notices, 12(6):36– 41, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hibbard:1977:SA

[854] P. G. Hibbard. A sublanguage of ALGOL 68. ACM SIGPLAN Notices, 12(5):71–79, May 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holdsworth:1977:SIA

[855] D. Holdsworth. System implementation in Algol 68-r. Software—Practice and Experience, 7(3):331–339, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoskins:1977:ANS

[856] W. D. Hoskins and G. E. McMaster. Algorithm 98: A note on the solution of certain tridiagonal systems of linear equations. The Computer Journal, 20(4):376–378, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_20/Issue_04/tiff/376.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/378.tif.

Housden:1977:CSP

[857] R. J. W. Housden and N. Kotarski. Character string pattern matching in Algol 68. ACM SIGPLAN Notices, 12(6):144–152, June 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Houssais:1977:VAI

[858] Bernard Houssais. Verification of an Algol 68 implementation. ACM SIGPLAN Notices, 12(6):117–128, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunter:1977:LVL

[859] R. B. Hunter, A. D. McGettrick, and R. Patel. LL versus LR parsing with illustrations from ALGOL 68. ACM SIGPLAN Notices, 12(6):49– 53, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jansen:1977:RLF

[860] J. K. M. Jansen. Remark on “Algorithm 259: Legendre Functions for Arguments Larger than One”. ACM Transactions on Mathematical Software, 3(2):204–250, June 1977. CODEN ACMSCU. ISSN 0098-3500. See [210].

Janssen:1977:EPI

[861] Theo Janssen and Peter van Emde Boas. The expressive power of intensional logic in the semantics of programming languages: (preprint). Dummy note., January 1977.

Janssen:1977:PTR

[862] Theo Janssen and Peter van Emde Boas. On the proper treatment of referencing, dereferencing and assignment: (preprint). Dummy note., January 1977.

Janssen:1977:SMG

[863] Theo Janssen. Simulation of a Montague grammar: (preprint). Dummy note., January 1977.

Kaletzky:1977:POL

[864] A. Kaletzky and D. W. Lewin. Problem orientated language for logic design. The Computer Journal, 20(2):163–169, May 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/2/163.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/163.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_02/tiff/165.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/tiff/166.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_20/

Issue_02/tiff/167.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/169.tif.

Kemp:1977:WEF

[865] P. Kemp. Writing the elementary function procedures for the AL-GOL68C compiler. In Cowell [1289], pages 127–144. ISBN 0-387-08446-0, 3-540-08446-0, 3-540-37071-4. ISSN 0302-9743. LCCN QA297 .W65

1976. URL http://link.springer.com/10.1007/3-540-08446-0.

Kerner:1977:SAI

[866] I. O. Kerner. A sublanguage of ALGOL 68 and its implementation. In Ershov and Koster [1290], pages 246–250. CODEN LNCSD9. ISBN 0-387-08065-1 (New York), 3-540-08065-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S96 1975; QA76.7.V7813

1975. URL http://link.springer-ny.com/link/service/series/
0558/tocs/t0047.htm;   http://www.springerlink.com/content/
978-0-387-08065-9;   http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=47. English translation from

the Russian of most of the proceedings (Trudy) of the symposium.

Koch:1977:BAI

[867] Wilfried Koch and Christoph Oeters. The Berlin ALGOL 68 implementation. ACM SIGPLAN Notices, 12(6):102–108, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Korhecz:1977:KSP

[868] Imre Korhecz. Ketnyelvu szamitogepes programok: ALGOL 60 es FOR-TRAN programozasi nyelvu programok tervezese, irasa es elemzese. (Hungarian) [Bilingual computer programs: design, writing and analysis of programs in the ALGOL 60 and FORTRAN programming languages]. Egyetemi Szamitokozpont, Budapest, Hungary, 1977. ISBN 963-451-020-

5. 368 pp. Two volumes.

Lindsey:1977:IIA

[869] C. H. Lindsey and S. G. van der Meulen. Informal introduction to ALGOL 68. North-Holland, Amsterdam, The Netherlands, second edition, 1977. ISBN 0-7204-0504-1, 0-7204-0726-5 (paperback). viii + 361 pp. LCCN QA76.73.A24 M47 1977. URL http://books.google.com/

books?id=8moZAQAAIAAJ.

Loveman:1977:PIS

[870] David B. Loveman. Program improvement by source-to-source transformation. Journal of the ACM, 24(1):121–145, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pagan:1977:AIL

[871] Frank G. Pagan. ALGOL 68 as an implementation language for portable interpreters. ACM SIGPLAN Notices, 12(6):54–62, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Poigne:1977:SPR

[872] A. Poigné and H. Huwig. Sprachkonzepte in der Programmierung: eine rekursionsorientierte Einführung in eine ALGOL-68-ähnlich Sprache. (German) [Language concepts in programming: a recursion-oriented introduction to an ALGOL-68-like language]. Bericht - Abteilung Informatik, Universität Dortmund; Nr. 46. Abt. Informatik, Univ. Dortmund, Dortmund, West Germany, 1977. 130 columns pp. LCCN QA76.73.A24 P64.

Prudom:1977:SPC

[873] A. Prudom and M. A. Hennell. Some problems concerning the automatic translation of Fortran to Algol 68. ACM SIGPLAN Notices, 12(6):138– 143, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ratcliff:1977:ASP

[874] Bryan Ratcliff. ALGOL 68 and structured programming for learnerprogrammers. ACM SIGPLAN Notices, 12(6):157–160, June 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rayward-Smith:1977:UPL

[875] V. J. Rayward-Smith. Using procedures in list processing. ACM SIG-PLAN Notices, 12(6):179–183, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scowen:1977:DFA

[876] R. S. Scowen. The diagnostic facilities in Algol and Fortran compilers. Report ??, Division of Numerical Analysis and Computing, National Physical Laboratory, Teddington, Middlesex, 1977. 43 pp.

Shave:1977:SUT

[877] M. J. R. Shave. The simulation and use of a TABLE data type in Algol W. Software—Practice and Experience, 7(6):733–745, November/ December 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Simonet:1977:ADS

[878] M. Simonet. An attribute description of a subset of Algol 68. ACM SIGPLAN Notices, 12(6):129–137, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stiller:1977:AD

[879] Gerd Stiller. ALGOL 68, Datenorganisation. (German) [ALGOL 68, Data Organization]. Reihe Datenverarbeitung. R. Oldenbourg, München, Germany, 1977. ISBN 3-486-39931-4. 160 pp. LCCN QA76.73.A24 S74 1977.

Szoke:1977:SRN

[880] Péter Szöke. Some remarks on new instances and garbage collection. ACM SIGPLAN Notices, 12(6):42–48, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanenbaum:1977:CTA

[881] Andrew S. Tanenbaum. Corrigenda: “A Tutorial on Algol 68”. ACM Computing Surveys, 9(3):255–256, September 1977. CODEN CMSVAN. ISSN 0010-4892. See [816].

Taupin:1977:ACP

[882] Daniel Taupin. The ALGOL 68 compiler of Paris-XI University (Orsay). ACM SIGPLAN Notices, 12(6):109–116, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomson:1977:ECT

[883] C. M. Thomson. Error checking, tracing, and dumping in an ALGOL 68 checkout compiler. ACM SIGPLAN Notices, 12(7):106–111, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderMeulen:1977:AMB

[884] S. G. van der Meulen. ALGOL 68 might-have-beens. ACM SIGPLAN Notices, 12(6):1–18, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderMeulen:1977:DVL

[885] S. G. van der Meulen and M. Veldhorst. Datastructuren voor lineaire ruimten “Torrix”. (Dutch) [Data structures for linear spaces “Torrix”]. Technical report RUU-CS-77-01, Universiteit Utrecht, Utrecht, The Netherlands, 1977. ?? pp.

VanLeeuwen:1977:REL

[886] Jan Van Leeuwen. Recursively enumerable languages and van Wijngaarden grammars. Indagationes Mathematicae (Proceedings), 80(1):29–39, 1977. ISSN 1385-7258 (print), 1878-5972 (electronic).

Vansina:1977:DDM

[887] C. F. Vansina. Description of dynamic manipulation of ALGOL 68 values using a generative grammar. ACM SIGPLAN Notices, 12(6):63–70, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanVliet:1977:TMI

[888] J. C. van Vliet. Towards a machine-independent transput section. ACM SIGPLAN Notices, 12(6):71–77, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanWijngaarden:1977:RRA

[889] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, C. H. A. Koster, Michel Sintzoff, C. H. Lindsey, L. G. L. T. Meertens, and R. G. Fisker. Revised report on the algorithmic language ALGOL 68. ACM SIGPLAN Notices, 12(5):1–70, May 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL https://dl.acm.org/doi/10.1145/954652.1781176;   https:
//five.sh/files/pdf/Technology/RR-ALGOL-68.pdf.

Vowels:1977:AFI

[890] R. A. Vowels. ALGOL 60 and FORTRAN IV. John Wiley, New York, NY, USA, 1977. ISBN 0-471-91192-5 (paperback). x + 173 pp. LCCN QA76.73.A24 V68.

West:1977:MRT

[891] P. R. West. A Multi Radar Tracking simulation using ALGOL 68’R. ACM SIGPLAN Notices, 12(6):175–178, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1977:WCW

[892] Niklaus Wirth. What can we do about the unnecessary diversity of notation for syntactic definitions? Communications of the ACM, 20(11): 822–823, November 1977. CODEN CACMA2. ISSN 0001-0782 (print),

1557-7317 (electronic). URL http://pascal.hansotten.com/uploads/
wirth/What%20can%20we%20do%20Notation%20EBNF.pdf.

Zimmermann:1977:MSH

[893] Gerhard Zimmermann. Microprogram structures for high level language elements. ACM SIGMICRO Newsletter, 8(3):47–54, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.acm.

org/doi/10.1145/1014196.803315.

Atkinson:1978:CCA

[894] L. V. Atkinson and J. J. McGregor. CONA: a conversational Algol system. Software—Practice and Experience, 8(6):699–708, November/ December 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Banerjee:1978:CCD

[895] Jayanta Banerjee, Richard I. Baum, and David K. Hsiao. Concepts and capabilities of a database computer. ACM Transactions on Database Systems, 3(4):347–384, December 1978. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://www.

acm.org/pubs/articles/journals/tods/1978-3-4/p347-banerjee/
p347-banerjee.pdf;   http://www.acm.org/pubs/citations/journals/
tods/1978-3-4/p347-banerjee/.

Blaser:1978:SAR

[896] J. Blaser. Ein System Von Algol-60-Unterprogrammen Zur Realisierung Einer Gleitpunktrechnung Mit Frei Wählbarer Mantissenlänge. Diplomarbeit, Inst. F. Angew. Math., University Freiburg I. Br., 1978. ?? pp.

Bos:1978:ANA

[897] H. J. Bos and D. T. Winter. Aflink: a new Algol 68–Fortran interface: (preprint). Report NN 17/78, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, De-

cember 1978. 24 pp. URL https://ir.cwi.nl/pub/9115.

Braid:1978:SCS

[898] I. C. Braid. On storing and changing shape information. Computer Graphics, 12(3):252–256, August 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Brailsford:1978:IAP

[899] D. F. Brailsford and A. N. Walker. Introductory Algol 68 programming. Ellis Horwood, Chichester, UK, 1978. ISBN 0-470-26746-1, 0-85312-127-3. 281 pp. LCCN ????

Carroll:1978:CFT

[900] Johnny Glen Carroll. A comparison of the formatted transput of ALGOL 68, PL/I, and FORTRAN. Thesis (m.s.), Oklahoma State University, Stillwater, OK, USA, 1978. vii + 197 pp.

Cheatham:1978:BDJ

[901] T. E. Cheatham, Jr. A brief description of JOVIAL. ACM SIGPLAN Notices, 13(8):201–202, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Colin:1978:PPS

[902] Andrew J. T. Colin. Programming and Problem-Solving in Algol 68. Computer Science Series. Macmillan Education UK, London, UK, 1978. ISBN 0-333-23115-5 (paperback), 1-349-03561-0, 1-349-03562-9. viii + 251 pp. LCCN QA76.73.A24 C64.

DeMorgan:1978:MAS

[903] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified Algol 60 and the step-until element. The Computer Journal, 21(3):282–285, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/282.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/282.tif.

Foxley:1978:MRT

[904] Eric Foxley and D. J. Morgan. Monitoring the run-time activity of Algol 68-R programs. Software—Practice and Experience, 8(1):29–34, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gries:1978:ASH

[905] David Gries. ACM SIGPLAN history of programming languages conference ALGOL 60 language summary. ACM SIGPLAN Notices, 13(8):

1, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hazlewood:1978:AAC

[906] L. Hazlewood. Algorithm 101: An algorithm for cautious adaptive quadrature. The Computer Journal, 21(2):180–183, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/180.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_02/tiff/182.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/tiff/183.
tif.

Janssen:1978:EPI

[907] Theo Janssen and Peter van Emde Boas. The expressive power of intensional logic in the semantics of programming languages. Dummy note., January 1978.

Janssen:1978:PTR

[908] Theo Janssen and Peter van Emde Boas. On the proper treatment of referencing, dereferencing and assignment. Dummy note., January 1978.

Kaucher:1978:HPA

[909] Edgar W. Kaucher, Rudi Klatte, and Christian Ullrich. Höhere Programmiersprachen ALGOL, FORTRAN, PASCAL in einheitlicher und ubersichtlicher Darstellung. Reihe Informatik; Bd. 24. Bibliographisches Institut, Mannheim, Germany, 1978. ISBN 3-411-01544-6. 258 pp.

Kline:1978:VDP

[910] Russell B. Kline, Gary D. Hamor, Kenneth L. Krause, and Larry E. Druffel. Visual demonstration of program execution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1): 16–18, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/CSA Technical Symposium on Computer Science Education.

Lengauer:1978:FAF

[911] Thomas Lengauer and Robert Endre Tarjan. A fast algorithm for finding dominators in a flow graph. Technical Report STAN-CS-78-650 (AD-A054144), Stanford University, Department of Computer Science, Stanford, CA, USA, March 1978. ii + 40 pp.

Maslov:1978:ASP

[912] A. N. Maslov. Algol 68. Structure of Programs. Moscow State University Publishers, Moscow, USSR, 1978. ???? pp.

McGettrick:1978:AFS

[913] Andrew D. McGettrick. ALGOL 68: a first and second course, volume 8 of Cambridge computer science texts. Cambridge University Press, Cambridge, UK, 1978. ISBN 0-521-21412-2. a0−521−29143−7b paperback. xii + 348 pp. LCCN QA76.73.A24 M3.

McGettrick:1978:ARA

[914] Andrew D. McGettrick. An Introduction to the Formal Definition of ALGOL 68, volume 9 (Part 1) of Annual Review in Automatic Programming. Pergamon, New York, NY, USA and Oxford, UK, 1978. ISBN 0-08-023056-3. ISSN 0066-4138. 1–84 pp. LCCN QA1 A62 v. 9, pt. 1.

Meek:1978:FPAa

[915] Brian L. (Brian Lawrence) Meek. Fortran, PL/I and the Algols. Macmillan computer science series. Macmillan Publishing Company, New York, NY, USA, 1978. ISBN 0-333-24385-4. viii + 291 pp. LCCN QA76.73.F25 M43 1978b.

Meek:1978:FPAb

[916] B. L. (Brian Lawrence) Meek. Fortran, PL/I, and the Algols. Elsevier North-Holland Publishing Co., Inc., New York, NY, USA, 1978. ISBN 0-444-19464-9. viii + 291 pp. LCCN QA76.73.F25 .M43.

Meek:1978:FPAc

[917] B. L. (Brian Lawrence) Meek. Fortran, PL/I, and the Algols. Macmillan computer science series. Macmillan Publishing Company, New York, NY, USA, 1978. ISBN 0-333-24385-4. viii + 291 pp.

Naur:1978:ESL

[918] Peter Naur. The European side of the last phase of the development of ALGOL 60. ACM SIGPLAN Notices, 13(8):15–44, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peck:1978:ASP

[919] John E. Peck. The ALGOL 68 story: a personal account by a member of the design team. ACS Bulletin, ??(??):4–6, November 1978. URL

https://bernd-oppolzer.de/ALGOL68-The-Algol-68-Story.pdf;

https://softwarepreservation.computerhistory.org/ALGOL/paper/
The%20Algol%2068%20Story.pdf.

Perlis:1978:ASD

[920] Alan J. Perlis. The American side of the development of Algol. ACM SIGPLAN Notices, 13(8):3–14, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosin:1978:ASH

[921] Robert F. Rosin. ACM SIGPLAN history of programming languages conference PL/I language summary. ACM SIGPLAN Notices, 13(8):225– 226, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smillie:1978:IAA

[922] K. W. Smillie. Introducing APL, ALGOL and FORTRAN. Report ??, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1978. i + 61 pp.

Tanenbaum:1978:CPA

[923] A. S. Tanenbaum. A comparison of Pascal and Algol 68. The Computer Journal, 21(4):316–323, November 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/4/316.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/316.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_04/tiff/318.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/tiff/319.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_21/
Issue_04/tiff/320.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/321.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/322.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/tiff/
323.tif.

vanderMeulen:1978:TPS

[924] S. G. van der Meulen and M. Veldhorst. TORRIX: A programming system for operations on vectors and matrices over arbitrary fields and of variable size. Volume I. Tract 86, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1978.

ISBN 90-6196-152-1. 246 pp. URL https://jemarch.net/torrix-1.
pdf.

VanEmdeBoas:1978:NRE

[925] Peter Van Emde Boas and Paul M. B. Vitányi. A note on the recursive enumerability of some classes of recursively enumerable languages. Information Sciences, 14(1):89–91, January 1978. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

vanWijngaarden:1978:RBA

[926] A. van Wijngaarden, Immo Ottomar Kerner, et al. Revidierter Bericht über die algorithmische Sprache ALGOL 68. (German) [Revised report on the algorithmic language ALGOL 68], volume 27 of Elektronisches Rechnen und Regeln: Sonderband. Akademie-Verlag, Berlin, Germany, 1978. x + 249 pp. LCCN QA76.73.A24 R4515.

Vowels:1978:FIA

[927] R. A. Vowels. FORTRAN IV y ALGOL 60. Editorial Limusa, Mexico, DF, Mexico, 1978. 202 pp.

Williams:1978:SSF

[928] M. H. Williams. Static semantic features of Algol60 and Basic. The Computer Journal, 21(3):234–242, August 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/3/234.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/234.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_03/tiff/236.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_21/
Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/239.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/240.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/tiff/
241.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/242.tif.

Yuldashev:1978:ARC

[929] Z. Kh. Yuldashev. Algorithms for the realization of computer interval arithmetic for computers with a nonsymmetric set of machine numbers in Algol 60 (in Russian). Algorithms and Programs — Inform. Bull., 1 (21):21–22, 1978.

Axford:1979:ASL

[930] T. H. Axford, Diana Burkhardt, W. P. Dodd, Susan Laflin, D. G. Parkyn, and P. Ramsay. ATOL: a simple language with powerful data structuring facilities. ACM SIGPLAN Notices, 14(3):5–15, March 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banatre:1979:EDC

[931] Jean-Pierre P. Banâtre, Jean-Paul P. Routeau, and L. Trilling. An eventdriven compiling technique. Communications of the ACM, 22(1):34–42, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berry:1979:TES

[932] Daniel M. Berry and Richard L. Schwartz. Type equivalence in strongly typed languages: one more look. ACM SIGPLAN Notices, 14(9):35–41, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bird:1979:REV

[933] R. S. Bird. Recursion elimination with variable parameters. The Computer Journal, 22(2):151–154, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/22/2/151.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/151.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_02/tiff/153.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/tiff/154.
tif.

Brailsford:1979:IAP

[934] D. F. Brailsford and A. N. Walker. Introductory Algol 68 programming, volume 2 of Computers and their applications. Ellis Horwood, Chichester, UK, 1979. ISBN 0-470-26746-1 (Halsted), 0-85312-127-3. 281 pp. LCCN QA76.73.A24 B72.

Clarke:1979:PLC

[935] Edmund Melson Clarke, Jr. Programming language constructs for which it is impossible to obtain good Hoare axiom systems. Journal of the ACM, 26(1):129–147, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

elLozy:1979:RAS

[936] Mohamed el Lozy. Remark on “Algorithm 395: Student’s t-distribution” and remark on “Algorithm 396: Student’s quantiles [S14]”. ACM Transactions on Mathematical Software, 5(2):238–239, June 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [492, 493, 1003, 1081].

Foxley:1979:BRB

[937] Eric Foxley. Book review: Algol 68: a first and second course. Andrew D. McGettrick, C.U.P., Cambridge, 1978. No. of pages: 360. Price: £12.50 (hard cover), £5.50 (paperback). Software—Practice and Experience, 9 (5):425, May 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Glanc:1979:ECD

[938] Alois Glanc. On expressing control and data structures in predicate logic language ALGOL M. ACM SIGPLAN Notices, 14(4):32–34, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1979:HSTa

[939] Susan L. Graham, William N. Joy, and O. Roubine. Hashed symbol tables for languages with explicit scope control. ACM SIGPLAN Notices, 14(8):50–57, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1979:HSTb

[940] Susan L. Graham, William N. Joy, and O. Roubine. Hashed symbol tables for languages with explicit scope control. In ACM SIGPLAN SCC ’79 [1292], pages 50–57. LCCN QA 76.7 A1095 v.14 pt.2.

Grune:1979:SSA

[941] Dick Grune. Some statistics on ALGOL 68 programs. ACM SIGPLAN Notices, 14(7):38–46, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gutknecht:1979:FAC

[942] M. H. Gutknecht. Fast algorithms for the conjugate periodic function. Computing, 22(1):79–91, March 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hennell:1979:ETN

[943] M. A. Hennell and D. Hedley. An experimental testbed for numerical software. II. ALGOL 68. The Computer Journal, 22(1):53–56, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/1/
53.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_01/tiff/54.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/
55.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/56.tif.

Hennell:1979:RPP

[944] M. A. Hennell. The role of procedural parameters in high order languages. ACM SIGPLAN Notices, 14(7):47–50, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Janssen:1979:PCQ

[945] Theo Janssen. On problems concerning the quantification rules in Montague grammar: (preprint). Dummy note., January 1979.

Kawai:1979:LSS

[946] Satoru Kawai. Lattice structure segmentation of ALGOL-like programs. Software—Practice and Experience, 9(6):485–498, June 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Langmaack:1979:FTP

[947] Hans Langmaack, Wolfram Lippe, and Franz Wagner. The formal termination problem for programs with finite ALGOL 68-modes. Information Processing Letters, 9(3):155–159, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

LeBlanc:1979:ISC

[948] Richard J. LeBlanc and Charles N. Fischer. On implementing separate compilation in block-structured languages. ACM SIGPLAN Notices, 14 (8):139–143, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lippe:1979:EPG

[949] W. M. Lippe and F. Simon. Zur Elimination von Prozedurschachtelungen. (German) [On the elimination of procedure nestings]. Computing, 23(2):119–138, June 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

McGettrick:1979:AAM

[950] Andrew D. McGettrick. Aspects of the ALGOL 68 mode structure. ACM SIGPLAN Notices, 14(7):62–77, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1979:FPA

[951] B. Meek. Fortran, PL-I and the Algols. American Elsevier, New York, NY, USA, September 1979. ISBN 0-444-19464-9. ???? pp. LCCN

????  URL http://www.cbooks.com/sqlnut/SP/search/gtsumt?
source=&isbn=0444194649.

Middleton:1979:AGA

[952] Tony Middleton. Automatic generation of aggregate operators in AL-GOL 68. Software—Practice and Experience, 9(8):685–690, August 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oliver:1979:PEA

[953] J. R. Oliver and R. S. Newton. Practical experience with Algol 68-Rt. The Computer Journal, 22(2):114–118, May 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/22/2/114.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/114.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_02/tiff/116.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/tiff/117.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_22/
Issue_02/tiff/118.tif.

Pagan:1979:AMD

[954] F. G. Pagan. Algol 68 as a metalanguage for denotational semantics. The Computer Journal, 22(1):63–66, February 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/22/1/63.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/63.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/65.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/66.tif.

Pagan:1979:SSU

[955] Frank G. Pagan. Semantic specification using two-level grammars: Blocks, procedures and parameters. Computer Languages, 4(3-4):171– 185, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Schneider:1979:DSI

[956] M. J. Schneider. Die Definition Der Sprache ‘Langzahl-Triplex-Algol 60’ Und Ihre Implementierung Durch Einen Präcompiler (Diplomarbeit). Freiburger Intervall-Ber. 79/3, Universität Freiburg, Freiburg, Germany, 1979. ?? pp.

Schonfelder:1979:APA

[957] J. L. Schonfelder and J. T. Thomason. Arbitrary precision arithmetic in Algol 68. Software—Practice and Experience, 9(3):173–182, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schwartz:1979:SVA

[958] Richard L. Schwartz and Daniel M. Berry. A semantic view of ALGOL 68. Computer Languages, 4(1):1–15, ???? 1979. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Smart:1979:MEP

[959] Robert K. Smart. Micro Extensible Programming Language. ACM SIG-PLAN Notices, 14(5):54–63, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Temme:1979:AAP

[960] N. M. Temme. An algorithm with ALGOL 60 program for the computation of the zeros of ordinary Bessel functions and those of their derivatives. Journal of Computational Physics, 32(2):270–279, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901347.

VanRemoortere:1979:BRR

[961] P. Van Remoortere. Book review: Realzeitrechner-Modelle: B. Walke: Published by R. Oldenburg Verlag, München, 1978, 240 seiten, 81 Abbildungen, 8 Tabellen, 5 ALGOL-Programme, DM.44,00. ISBN 3-486-21551-5. Mathematics and Computers in Simulation, 21(4):390, December 1979. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-

7166 (electronic). URL https://www.sciencedirect.com/science/
article/pii/0378475479900417.

vanVliet:1979:AT

[962] J. C. van Vliet. ALGOL 68 transput, volume 110–111 of Mathematical Centre tracts. Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1979. ISBN 90-6196-178-5 (paperback): v. 1, 90-6196-179-3 (paperback): v. 2. ???? pp. LCCN QA76.73.A24 V55.

Zuse:1979:REC

[963] Konrad Zuse. Die Rolle der ETH bei der Computerentwicklung. Dem Andenken an Eduard Stiefel. (German) [The role of ETH in computer development: In memory of Eduard Stiefel]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30(2):399–403, March 1979. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Aretz:1980:SSA

[964] F. E. J. Kruseman Aretz, Jan L. A. van de Snepscheut, H. Grasdijk, and J. M. H. Smeets. SATHE: Some aspects of an ALGOL implementation. Software—Practice and Experience, 10(7):563–573, July 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Boom:1980:CCS

[965] H. J. Boom and E. de Jong. A critical comparison of several programming language implementations. Software—Practice and Experience, 10(6): 435–473, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brailsford:1980:BRB

[966] David F. Brailsford. Book review: Informal introduction to Algol 68 (revised edn), C. H. Lindsey and S. G. van der Meulen, North Holland, 1977. No. of pages: 361. Price: £14.00. Software—Practice and Experience, 10(2):161–162, February 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Conway:1980:BRI

[967] D. E. Conway. Book review: Introductory Algol 68 Programming by D. F. Brailsford, A. N. Walker. Mathematical Gazette, 64(429):224, 1980. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

http://www.jstor.org/stable/3615158.

Culik:1980:WFL

[968] Karel Čulik. What is a flowchart loop and about structured programming. ACM SIGPLAN Notices, 15(1):45–57, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Evans:1980:ADP

[969] D. J. Evans and Shirley A. Williams. Analysis and detection of parallel processable code. The Computer Journal, 23(1):66–72, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/1/
66.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/67.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
68.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_01/tiff/71.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/
72.tif.

Ford:1980:BRB

[970] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_23/Issue_02/tiff/146.tif.

Ford:1980:BRW

[971] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_23/Issue_02/tiff/146.tif.

Garside:1980:API

[972] G. R. Garside and P. E. Pintelas. An ALGOL 68 package for implementing graph algorithms. The Computer Journal, 23(3):237–242, August 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/3/237.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
23/Issue_03/tiff/237.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_03/tiff/238.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/tiff/239.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_23/
Issue_03/tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_23/Issue_03/tiff/241.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_23/Issue_03/tiff/242.tif.

Gobin:1980:FHP

[973] Marc A. Gobin. File handling in programming languages. ACM SIG-PLAN Notices, 15(10):42–47, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guenther:1980:PSI

[974] G. Guenther and G. Marquardt. A programming system for interval arithmetic in Algol 68. In K. Nickel, editor, Interval Mathematics 1980, pages 355–366. Academic Press, New York, USA, 1980.

Hammerlin:1980:PKA

[975] Günther Hämmerlin and Larry L. Schumaker. Procedures for kernel approximation and solution of Fredholm integral equations of the second kind. Numerische Mathematik, 34(2):125–141, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Igarashi:1980:SAP

[976] Y. Igarashi. The size of arrays for a prime implicant generating algorithm. The Computer Journal, 23(1):73–77, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/23/1/73.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/73.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_01/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_23/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_23/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_
01/tiff/77.tif.

Ince:1980:ABA

[977] Darrel C. Ince and Keith Robson. An Algol 68 based algebraic manipulation system. Software—Practice and Experience, 10(6):427–430, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1980:IIL

[978] D. C. Ince. An interpretative implementation of limited entry decision tables in ALGOL 68. ACM SIGPLAN Notices, 15(2):43–45, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joshi:1980:LCP

[979] A. K. Joshi, L. S. Levy, and K. Yueh. Local constraints in programming languages. I. syntax. Theoretical Computer Science, 12(3):265–290, November 1980. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kumar:1980:PRB

[980] Gopa Kumar and C. Thomas Nute. Program restructuring for block structured languages. ACM SIGMETRICS Performance Evaluation Review, 9(2):67–79, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McClintock:1980:TCA

[981] John McClintock and Paul Klint. Technical correspondence: Algol-W approach to line number maintenance. Communications of the ACM, 23 (12):729, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGettrick:1980:AFS

[982] Andrew D. McGettrick. ALGOL 68: a First and Second Course, volume 8 of Cambridge computer science texts. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-21412-2 (hardcover), 0-521-29143-7 (paperback). xii + 348 pp. LCCN QA76.73.A24 M3.

Pagan:1980:NSA

[983] Frank G. Pagan. Nested sublanguages of ALGOL 68 for teaching purposes. ACM SIGPLAN Notices, 15(7–8):72–81, July/August 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siero:1980:AAC

[984] P. L. J. Siero. APL and Algol68, the correspondence and the differences, especially in applications of graph-analysis. ACM SIGPLAN Notices, 15 (12):56–61, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Starostenko:1980:MAF

[985] Vitalii Ivanovich Starostenko and Antonina Nikolaevna Zavorotko. Metodika i Algol (Fortran)–programmy ustoichivogo resheniia obratnykh lineinykh i nelineinykh zadach gravimetrii. “Nauk. dumka”, Kiev, USSR, 1980. 102 pp.

vanderMeulen:1980:ADT

[986] S. G. van der Meulen. Abstract data types for numerical applications and some thoughts on precision. Presentation at IFIP WG 2.5 (Numerical Software) meeting, Harwell, 1980, January 14, 1980.

URL https://web.archive.org/web/20061009123241/https://www.
nsc.liu.se/~boein/ifip/projects/p14.txt.

xxx:1980:IAS

[987] ???? Intval — Ein Algol 68 System Zur Intervallarithmetik. Rrzn-Publ. No. Spr. A68 8, Regionales Rechenzentrum Für Niedersachsen, Universität Hannover, Hannover, West Germany, 1980. ?? pp.

Anonymous:1981:RBA

[988] Anonymous. Relationships between and among models. ACM SIGPLAN Notices, 16(1):77–82, January 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backus:1981:CSB

[989] John Backus. Is computer science based on the wrong fundamental concept of ‘program’? An extended concept (invited address). In de Bakker and van Vliet [1293], pages 133–165. ISBN 0-444-86285-4. LCCN QA76.7

.I575 1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Bajpai:1981:FAP

[990] Avinash Chandra Bajpai. Fortran and Algol: a programmed course for students of science and technology. A series of programmes on mathematics for scientists and technologists. John Wiley, New York, NY, USA, 1981. ISBN ???? 273 pp. LCCN ????

Bauer:1981:ASP

[991] F. L. Bauer and H. Wössner. Algorithmische Sprache und Programmentwicklung. Unter Mitarbeit von H. Partsch und P. Pepper. (German) [Algorithmic language and program development. With the collaboration of H. Partsch and P. Pepper]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-09853-4, 3-540-09853-4. xv + 513 pp. LCCN QA76.6 .B395 1981.

Bossavit:1981:DVP

[992] Alain Bossavit and Bertrand Meyer. The design of vector programs. In de Bakker and van Vliet [1293], pages 99–114. ISBN 0-444-86285-

4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/34328/
34328D.pdf.

Broy:1981:DPP

[993] M. Broy, P. Pepper, and M. Wirsing. On design principles for programming languages: an algebraic approach. In de Bakker and van Vliet [1293], pages 203–219. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981.

URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Darlington:1981:SDA

[994] John Darlington. The structured description of algorithm derivations (invited address). In de Bakker and van Vliet [1293], pages 221–250. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.

nl/pub/34328/34328D.pdf.

deChampeaux:1981:SEL

[995] Dennis de Champeaux and Jos de Bruin. Symbolic evaluation of LISP functions with side effects for verification. In de Bakker and van Vliet [1293], pages 271–291. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981.

URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Dekker:1981:FMA

[996] K. Dekker. Formula manipulation in ALGOL 68 and application to Routh’s algorithm. Computing, 26(2):167–187, June 1981. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

eertens:1981:MAC

[997] L. G. L. T. eertens and J. C. van Vliet. On the MC Algol 68 compiler. Report IW 90/91, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, December 1981.

10 pp. URL https://ir.cwi.nl/pub/9310.

Fokkinga:1981:NST

[998] Maarten M. Fokkinga. On the notion of strong typing. In de Bakker and van Vliet [1293], pages 305–320. ISBN 0-444-86285-4. LCCN QA76.7

.I575 1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Freak:1981:FPT

[999] R. A. Freak. A Fortran to Pascal translator. Software—Practice and Experience, 11(7):717–732, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Grumberg:1981:PRF

[1000] 0rna Gr”umberg, Nissim Francez, Johann A. Makowsky, and Willem P. de Roever. A proof rule for fair termination of guarded commands. In de Bakker and van Vliet [1293], pages 399–416. ISBN 0-444-86285-

4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/34328/
34328D.pdf.

Grune:1981:VGA

[1001] D. Grune. From VW-grammar to ALEPH. In de Bakker and van Vliet [1293], pages 185–202. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981.

URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Guenther:1981:PSI

[1002] G. Guenther and G. Marquardt. A programming system for interval arithmetic in Algol 68. In J. C. Van Vliet and H. Wupper, editors, Proceedings of the International Conference on Algol 68, Rechenzentrum Der Ruhr-Universität Bochum, March 30–31, 1981, volume 134(2) of Math. Centre Tracts, pages 201–215. ??, ??, 1981.

Hill:1981:RSD

[1003] G. W. Hill. Remark on “Algorithm 395: Student’s t-distribution”. ACM Transactions on Mathematical Software, 7(2):247–249, June 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [495, 496, 936].

Hill:1981:RSQ

[1004] G. W. Hill. Remark on “Algorithm 396: Student’s t-quantiles”. ACM Transactions on Mathematical Software, 7(2):250–251, June 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [496].

Ince:1981:DTA

[1005] Darrel C. Ince. Decision tables in Algol 68. Software—Practice and Experience, 11(7):753–756, July 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ince:1981:IFS

[1006] Darrel C. Ince. The implementation of file structures in some high level languages. ACM SIGPLAN Notices, 16(11):49–55, November 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonkers:1981:ASS

[1007] H. B. M. Jonkers. Abstract storage structures. In de Bakker and van Vliet [1293], pages 321–343. ISBN 0-444-86285-4. LCCN QA76.7 .I575

1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Klint:1981:FLD

[1008] Paul Klint. Formal language definitions can be made practical. In de Bakker and van Vliet [1293], pages 115–132. ISBN 0-444-86285-

4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/34328/
34328D.pdf.

Kuiper:1981:OSB

[1009] R. Kuiper. An operational semantics for bounded nondeterminism equivalent to a denotational one. In de Bakker and van Vliet [1293], pages 373–398. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https:

//ir.cwi.nl/pub/34328/34328D.pdf.

Lenstra:1981:LFP

[1010] A. K. Lenstra. Lattices and factorization of polynomials. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 15(3):15–16, August 1981. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Meertens:1981:IDB

[1011] Lambert G. L. T. Meertens. Issues in the design of a beginners’ programming language. In de Bakker and van Vliet [1293], pages 167–184. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.

nl/pub/34328/34328D.pdf.

Naur:1981:AVW

[1012] Peter Naur. Aad van Wijngaarden’s contributions to ALGOL 60 (invited address). In de Bakker and van Vliet [1293], pages 293–304. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/

34328/34328D.pdf.

Paul:1981:F

[1013] M. Paul. Foreword. In de Bakker and van Vliet [1293], pages xxi–xxiv. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.

nl/pub/34328/34328D.pdf.

Reynolds:1981:EAI

[1014] John C. Reynolds. The essence of ALGOL (invited address). In de Bakker and van Vliet [1293], pages 345–372. ISBN 0-444-86285-4. LCCN QA76.7

.I575 1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Roberts:1981:NCG

[1015] J. D. Roberts. Naming by colours: A graph-theoretic approach to distributed structure. In de Bakker and van Vliet [1293], pages 59–76. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.

nl/pub/34328/34328D.pdf.

Sato:1981:H

[1016] Masahiko Sato and Masami Hagiya. HYPERLISP. In de Bakker and van Vliet [1293], pages 251–269. ISBN 0-444-86285-4. LCCN QA76.7 .I575

1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Schild:1981:PPB

[1017] Rudolf Schild. PORTAL — a PASCAL-based real-time programming language. In de Bakker and van Vliet [1293], pages 49–58. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/

34328/34328D.pdf.

Smillie:1981:PNA

[1018] K. W. Smillie. Programming notes: APL, FORTRAN, ALGOL, PAS-CAL. Report ??, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1981. ii + 65 pp.

Turski:1981:ART

[1019] Władysław M. Turski. ALGOL 68 revisited twelve years later or from AAD to ADA (invited address). In de Bakker and van Vliet [1293], page 417. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.

cwi.nl/pub/34328/34328D.pdf.

Warren:1981:OIA

[1020] Henry S. Warren, Jr. Optimization of inductive assertions. In de Bakker and van Vliet [1293], pages 77–98. ISBN 0-444-86285-4. LCCN QA76.7

.I575 1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Wasserman:1981:PAL

[1021] Anthony I. Wasserman, Reind P. van de Riet, and Martin L. Kersten. PLAIN: An algorithmic language for interactive information systems. In de Bakker and van Vliet [1293], pages 29–48. ISBN 0-444-86285-4. LCCN

QA76.7 .I575 1981. URL https://ir.cwi.nl/pub/34328/34328D.pdf.

Zemanek:1981:RPV

[1022] Heinz Zemanek. The role of Professor A. van Wijngaarden in the history of IFIP (invited opening address). In de Bakker and van Vliet [1293], pages 1–47. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https:

//ir.cwi.nl/pub/34328/34328D.pdf.

AboElNaga:1982:HAM

[1023] N. M. Abo El Naga and J. A. Field. A hardware Algol machine. IEEE Micro, 2(4):37–47, November 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Atkinson:1982:PAA

[1024] Malcolm Atkinson, Ken Chisholm, and Paul Cockshott. PS-algol: an Algol with a persistent heap. ACM SIGPLAN Notices, 17(7):24–31, July 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bauer:1982:ALP

[1025] F. L. Bauer and H. Wössner. Algorithmic Language and Program Development. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11148-4, 3-540-11148-4. xvi + 497 pp. LCCN QA76.6 .B39513 1982. In memorian Klaus Samelson (1918–1980). In collaboration with H. Partsch and P. Pepper. Translated from the German original [991].

Cornelius:1982:BRB

[1026] B. J. Cornelius. Book review: Introductory Algol 68 programming, D. F. Brailsford and A. N. Walker, Ellis Horwood, Chichester. No. of pages:

281. Price: £14.00 (hardback), £5.95 (paperback). Software—Practice and Experience, 12(12):1166–1167, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Damm:1982:IOH

[1027] W. Damm. The IO- and OI-hierarchies (tree language theory). Theoretical Computer Science, 20(2):95–207, May 1982. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ford:1982:ORV

[1028] Gary Ford and Brian Hansche. Optional, repeatable, and varying type parameters. ACM SIGPLAN Notices, 17(2):41–48, February 1982. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gunn:1982:CTT

[1029] Hamish I. E. Gunn. Compile time type checking of structure field accessing. Information Processing Letters, 14(1):22–25, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hemker:1982:LMR

[1030] Piet Hemker and Paul de Zeeuw. A library of multigrid routines: ALGOL 68. Dummy note., January 1982.

Langmaack:1982:TPF

[1031] Hans Langmaack. On termination problems for finitely interpreted ALGOL-like programs. Acta Informatica, 18(1):79–108, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Leverett:1982:ASD

[1032] Bruce W. Leverett and Peter G. Hibbard. An adaptive system for dynamic storage allocation. Software—Practice and Experience, 12(6):543– 555, June 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Messerschmidt:1982:CCO

[1033] J. Messerschmidt and R. Wilhelm. Constructors from composed objects. Computer Languages, 7(2):53–59, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Osipov:1982:IAE

[1034] L. A. (Lev Aleksandrovich) Osipov. IAzyk analitik i ego sravnenie s iazykami ALGOL i FORTRAN. Bibliotechka programmista. “Nauka,” Glav. red. fiziko-matematicheskoi lit-ry, Moskva, USSR, 1982. ISBN ???? 160 pp. LCCN ????

Osipov:1982:JAE

[1035] L. A. (Lev Aleksandrovich) Osipov. Jazyk analitik i ego sravnenie s jazykami Algol i Fortran. Biblioteka programmista. Nauka, Moscow, USSR, 1982. ISBN ???? 160 pp. LCCN ????

Reed:1982:GPM

[1036] Karl Reed. On a general property of memory mapping tables. ACM SIGPLAN Notices, 17(4):81–86, April 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scheer:1982:AFA

[1037] Linda Sue Scheer. Ada, FORTRAN, ALGOL, JOVIAL, Pascal, PL/I, and LISP compared to Ada design requirements. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1982. x + 121 pp.

Scowen:1982:SSP

[1038] R. S. Scowen and Z. J. Ciechanowicz. Seven sorts of programs. ACM SIGPLAN Notices, 17(3):74–79, March 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Veldhorst:1982:ASM

[1039] M. Veldhorst. An analysis of sparse matrix storage schemes. Tract 150, Mathematisch Centrum (2 Boerhaavestraat 49), University of Amsterdam, Amsterdam, The Netherlands, 1982. ISBN 90-6196-242-0. 254 pp.

URL https://jemarch.net/torrix-2.pdf.

Anonymous:1983:CMC

[1040] Anonymous. Centre for mathematics and computer science: facts & figures. Dummy note., September 1983.

Anonymous:1983:CVW

[1041] Anonymous. Centrum voor wiskunde en informatica: enkele statistische en organisatorische gegevens. (Dutch) [Center for Mathematics and Computer Science: some statistical and organizational data]. Dummy note., September 1983.

Atkinson:1983:APP

[1042] M. P. Atkinson, P. J. Bailey, K. J. Chisholm, P. W. Cockshott, and R. Morrison. An approach to persistent programming. The Computer Journal, 26(4):360–365, November 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/26/4/360.full.pdf+html;

http://www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
04/tiff/360.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_26/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_04/tiff/362.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_26/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_26/
Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_04/tiff/365.tif. Also published in/as: In

“Readings in Object-Oriented Database Systems” edited by S. Zdonik and D. Maier, Morgan Kaufman, 1990.

Barnard:1983:RDP

[1043] D. T. Barnard. Recursive descent parsing using implementation languages requiring definition before use. Information Processing Letters, 17(5):255–258, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cormack:1983:ESS

[1044] G. V. Cormack. Extensions to static scoping. ACM SIGPLAN Notices, 18(6S):187–191, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldmann:1983:AG

[1045] Harry Feldmann. Acceptive grammars. ACM SIGPLAN Notices, 18(7): 25–44, July 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardy:1983:NSP

[1046] Norman Hardy. Nested scopes in Pascal and Algol 68. ACM SIGPLAN Notices, 18(4):65, April 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Irons:1983:SDC

[1047] Edgar T. Irons. A syntax directed compiler for ALGOL 60 (reprint). Communications of the ACM, 26(1):14–16, 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [57].

Maher:1983:API

[1048] B. Maher and D. H. Sleeman. Automatic program improvement: Variable usage transformations. ACM Transactions on Programming Languages and Systems, 5(2):236–264, April 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Metcalf:1983:WF

[1049] Michael Metcalf. Why Fortran? ACM ForTec Forum, 2(1):13–14, March 1983. CODEN ???? ISSN 0735-3731.

Morrison:1983:EHL

[1050] R. Morrison, A. J. Cole, P. J. Bailey, M. A. Wolfe, and M. Shearer. Experience with a high level language that supports interval arithmetic. In IEEE SCA6 ’83 [1295], pages 74–78. ISBN 0-8186-4476-1 (paperback), 0-8186-8476-3 (hardcover), 0-8186-6476-2 (microfiche), 0-8186-0034-9 (hardcover). LCCN QA 76.9 C62 S95 1983. URL http://

www.acsel-lab.com/arithmetic/arith6/papers/ARITH6_Morrison_
Cole_Bailey_Wolfe_Shearer.pdf. IEEE catalog number 83CH1892-9.

IEEE Computer Society order number 476.

Mulders:1983:SOO

[1051] Huub Mulders. Some observations on the in- and output in high-level languages. ACM SIGPLAN Notices, 18(9):55–58, September 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nadrchal:1983:IAS

[1052] Jaroslav Nadrchal, Ivan Sklenář, Václav Kříž, and Jaroslav Král. An implementation of Algol 68 for a small computer. Software—Practice and Experience, 13(7):597–660, July 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smillie:1983:PNA

[1053] K. W. Smillie. Programming notes: APL, FORTRAN, ALGOL, PAS-CAL. Report ??, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1983. ii + 105 pp.

Temme:1983:NCC

[1054] N. M. Temme. The numerical computation of the confluent hypergeometric function U(a, b, z). Numerische Mathematik, 41(1):63–82, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Woodward:1983:GAU

[1055] Philip M. Woodward and Susan G. Bond. Guide to ALGOL 68 for users of RS systems. E. Arnold, London, UK, 1983. ISBN 0-7131-3490-9. x + 148 pp. LCCN QA76.73.A24 W67 1983.

Zellweger:1983:IHLa

[1056] Polle T. Zellweger. An interactive high-level debugger for control-flow optimized programs. ACM SIGPLAN Notices, 18(8):159–172, August 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zellweger:1983:IHLb

[1057] Polle T. Zellweger. An interactive high-level debugger for control-flow optimized programs. ACM SIGSOFT Software Engineering Notes, 8 (4):159–172, August 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Anonymous:1984:PSM

[1058] Anonymous. Proceedings seminar 1981-1982 mathematical structures in field theories: [lectures given during the seminar at the academic year 1981-1982 at Amsterdam]. C, January 1984. ?? pp.

Anonymous:1984:VCV

[1059] Anonymous. Vakantiecursus 1984 Centum voor Wiskunde en Informatica: Hewet — plus wiskunde. (Dutch) [Holiday course 1984 Centre for Mathematics and Computer Science: Hewet — plus mathematics]. C, January 1984. ?? pp.

Cardelli:1984:BPT

[1060] Luca Cardelli. Basic polymorphic typechecking. Computing Science Technical Report 112, AT&T Bell Laboratories, Murray Hill, NJ,

USA, ???? 1984. 17 pp. URL http://lucacardelli.name/Papers/
BasicTypechecking%20(TR%201984).pdf.

Deikalo:1984:NSP

[1061] G. F. (Gennadiĭ Fedorovich) Deĭkalo et al. Novye sredstva programmirovaniëiìa dlëiìa ES çEVM: translëiìator s ëiìazyka Algol 68 i dialogovaëiìa sistema JEC. “Finansy i statistika”, Moskva, USSR, 1984. 206 pp. LCCN QA76.73.A24 N68 1984.

Deykalo:1984:NPI

[1062] G. Deykalo, A. N. Terekhov, et al. New Programming Instruments for ES EVM. Finansy i Statistika Publishers, Moscow, USSR, 1984. ISBN ???? ???? pp.

Dongarra:1984:EPS

[1063] J. J. Dongarra and C. B. Moler. EISPACK — A package for solving matrix eigenvalue problems. In Cowell [1296], pages 68–87. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

ISO:1984:PLA

[1064] ISO. Programming language: ALGOL 60: ISO 1538:1984 (en), volume 1538 of ISO recommendation. International Organization for Standardization, Geneva, Switzerland, 1984. ISBN ???? ?? pp. LCCN ???? URL

https://www.iso.org/obp/ui/en/#iso:std:iso:1538:ed-1:v1:en.

Jabs:1984:AFA

[1065] Arthur Jabs. Approximation formula and Algol program of the Lorentz-invariant momentum-space integral for particles of equal masses. Computer Physics Communications, 35(1–3):C–175–C–176, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0010465584824236.

Kallenberg:1984:TSH

[1066] W. C. M. Kallenberg. Testing statistical hypotheses: worked solutions. C, January 1984. ?? pp.

Kuo-Petravic:1984:TPT

[1067] L. G. Kuo-Petravic, M. Petravic, and K. V. Roberts. Tranal — a program for the translation of symbolic Algol I into symbolic Algol II. Computer Physics Communications, 35(1–3):C–356, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S001046558482576X.

Ollongren:1984:CEI

[1068] Alexander Ollongren. Classroom experience with interactive formula manipulation. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 18–19(4–1):31–37, November/February 1984. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Sheldon:1984:CTDb

[1069] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bγ) and (a, bγ − γ). (III) Fortran translations of the Algol programs ‘mandy’ and ‘barbara’. Computer Physics Communications, 35(1–3):C–91–C–93, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0010465584823528.

Verhelst:1984:PIP

[1070] P. W. E. Verhelst and N. F. Verster. PEP: an interactive programming system with an ALGOL-like programming language. Software—Practice and Experience, 14(2):119–133, February 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Woodward:1984:AHS

[1071] M. R. Woodward. The application of Halstead’s software science theory to Algol 68 programs. Software—Practice and Experience, 14(3):263– 276, March 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Allison:1985:PDS

[1072] Lloyd Allison. Programming denotational semantics. II. The Computer Journal, 28(5):480–486, November 1985. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/28/5/480.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_
05/tiff/480.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_28/Issue_05/tiff/481.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_05/tiff/482.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_05/tiff/483.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_28/
Issue_05/tiff/484.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_05/tiff/485.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_05/tiff/486.tif.

Anonymous:1985:ATA

[1073] Anonymous. AI tools automate software translation. Electronics, 58(38): 59–61, September 23, 1985. ISSN 0883-4989.

Anonymous:1985:PCC

[1074] Anonymous. Parallel computers and computations: [a series of eight lectures organized at the University of Utrecht, fall 1983]. C, January 1985. ?? pp.

Anonymous:1985:VCV

[1075] Anonymous. Vakantiecursus 1985 Centrum voor Wiskunde en Informatica: variatierekening. (Dutch) [Holiday course 1985 Center for Mathematics and Computer Science: calculus of variations]. C, January 1985. ?? pp.

DeBruin:1985:CVN

[1076] R. De Bruin and C. G. Van Der Laan. The creation of a virtual NAG– Algol 68 program library. Software—Practice and Experience, 15(10): 963–972, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Delves:1985:DAD

[1077] L. M. Delves and S. C. Mawdsley. DAP-Algol: a development system for parallel algorithms. The Computer Journal, 28(2):148–153, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/28/2/148.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_28/
Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_02/tiff/150.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/
151.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
28/Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/153.tif.

Delves:1985:DDS

[1078] L. M. Delves and S. C. Mawdsley. DAP-Algol: a development system for parallel algorithms. The Computer Journal, 28(2):148– 153, May 1985. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_28/Issue_02/tiff/148.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/149.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_28/
Issue_02/tiff/150.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_28/Issue_02/tiff/151.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_28/Issue_02/tiff/152.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_28/Issue_02/tiff/
153.tif.

Fogiel:1985:HGC

[1079] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FORTRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, 1985. ISBN 0-87891-561-3 (paperback). v + 122 pp.

Harris:1985:RD

[1080] Lawrence A. Harris and Yannich Wets. Resetting displays. ACM SIG-PLAN Notices, 20(8):73–77, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1985:RCS

[1081] I. D. Hill and M. C. Pike. Remark on “Algorithm 299: Chi-Squared Integral”. ACM Transactions on Mathematical Software, 11(2):185, June 1985. CODEN ACMSCU. ISSN 0098-3500. See [301, 798].

Kristensen:1985:MSE

[1082] Bent Bruun Kristensen, Ole Lehrmann Madsen, Birger Møller-Pedersen, and Kristen Nygaard. Multi-sequential execution in the BETA programming language. ACM SIGPLAN Notices, 20(4):57–70, April 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:1985:PAB

[1083] R. R. Martin and C. Anderson. A proposal for an ALGOL 68 binding of GKS. Computer Graphics Forum, 4(1):43–57, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shreeve:1985:BRG

[1084] Richard I. Shreeve. Book review: Guide to ALGOL 68 for Users of RS Systems by Philip M. Woodward, Susan G. Bond. Mathematical Gazette, 69(447):66–67, 1985. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL http://www.jstor.org/stable/3616477.

Sklenar:1985:PLA

[1085] Ivan Sklenar. Programming language assessments using the program profiles. ACM SIGPLAN Notices, 20(4):80–87, April 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1985:PLD

[1086] Niklaus Wirth. From programming language design to computer construction. Communications of the ACM, 28(2):160–164, February 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/toc/Abstracts/0001-0782/2789.html.

Wupper:1985:SSG

[1087] Hanno Wupper. Zur Spezifikation und Struktur grösserer Systeme: Untersuchungen am Beispiel von Algol 68 und GKS. (German) [On the specification and structure of larger systems: investigations using the example of Algol 68 and GKS], volume 4 of Rechenzentrums-Schriften. Rechenzentrum der Ruhr-Universität Bochum, Bochum, West Germany, 1985. ISSN 0721-2186. 234 pp. LCCN QA76.9.S88 W87 1985.

Yemini:1985:MVE

[1088] Shaula Yemini and Daniel M. Berry. A modular verifiable exceptionhandling mechanism. ACM Transactions on Programming Languages and Systems, 7(2):214–243, April 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0164-0925/3320.html.

Anonymous:1986:EUC

[1089] Anonymous. EUROGRAPHICS (UK) Conference. Computer Graphics Forum, 5(2):151–157 (or 131–157??), June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bahlke:1986:PSF

[1090] Rolf Bahlke and Gregor Snelting. The PSG system: From formal language definitions to interactive programming environments. ACM Transactions on Programming Languages and Systems, 8(4):547–576, October 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/0164-0925/
20890.html.

Clay:1986:NSQ

[1091] C. Clay. A new solution to the N ≤ 8 queens problem. ACM SIGPLAN Notices, 21(8):28–30, August 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

House:1986:ASR

[1092] R. T. House. Alternative scope rules for block-structured languages. The Computer Journal, 29(3):253–260, June 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/29/3/253.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
03/tiff/253.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_03/tiff/255.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/256.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_29/
Issue_03/tiff/257.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_03/tiff/258.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_29/Issue_03/tiff/259.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_03/tiff/
260.tif.

Kulkarni:1986:EEF

[1093] K. G. Kulkarni and M. P. Atkinson. EFDM: Extended functional data model. The Computer Journal, 29(1):38–46, February 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/29/1/38.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/38.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_29/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_29/Issue_01/tiff/44.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_29/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_29/Issue_
01/tiff/46.tif.

Lewi:1986:DSP

[1094] Johan Lewi and Jan Paredaens. Data structures of Pascal, Algol 68, PL/1, and Ada. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-15121-4, 0-387-15121-4. xii + 395 pp. LCCN ????

Morrison:1986:IGP

[1095] R. Morrison, A. L. Florianis, A. Dearle, and M. P. Atkinson. An integrated graphics programming environment. Computer Graphics Forum, 5(2):147–157, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patel:1986:IAP

[1096] R. R. Patel and R. B. Hunter. Implementing ALGOL 68 predicates. Software—Practice and Experience, 16(1):87–97, January 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

REA:1986:HGC

[1097] Research and Education Association. Handbook and guide for comparing and selecting the computer languages: BASIC, FORTRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, revised print. edition, 1986. ISBN ???? v + 122 pp. LCCN ????

Sathi:1986:PBC

[1098] Harbans L. Sathi. A project-based course in compiler construction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1):114–119, February 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 17th SIGCSE Symposium on Computer Science Education.

Stroustrup:1986:OC

[1099] Bjarne Stroustrup. An overview of C++. ACM SIGPLAN Notices, 21 (10):7–18, October 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1986:VIP

[1100] D. Taylor. Visionaries: An interview with Professor Michael O’Leary. ACM SIGPLAN Notices, 21(6):41–54, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandeLune:1986:ITT

[1101] Jan van de Lune. An introduction to Tauberian theory: from Tauber to Wiener. C, January 1986. ?? pp.

Walker:1986:IPP

[1102] A. N. Walker. The Ilem Project: Pascal to Algol in easy stages I: Outward bound. Software—Practice and Experience, 16(12):1097–1107, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baayen:1987:MVW

[1103] P. C. Baayen and J. Nuis. In memoriam: A. van Wijngaarden 1916—

1987. Web document, 1987. URL https://ir.cwi.nl/pub/1706/
1706D.pdf.

deGoede:1987:VOE

[1104] Erik de Goede and J. H. M. ten Thije Boonkkamp. Vectorization of the odd-even Hopscotch scheme and the alternating direction implicit scheme for the two-dimensional Burger’s equations. Dummy note., January 1987.

Fogiel:1987:HGC

[1105] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FORTRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, revised print. edition, 1987. ISBN ???? v + 122 pp. LCCN ????

Gannon:1987:TIM

[1106] John D. Gannon and Marvin V. Zelkowitz. Two implementation models of abstract data types. Computer Languages, 12(1):21–25, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Huijsman:1987:TAP

[1107] R. D. Huijsman, J. van Katwijk, C. Pronk, and W. J. Toetenel. Translating Algol 60 programs into Ada: Report on a feasability study. ACM SIGADA Ada Letters, 7(5):42–50, September/October 1987. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kulkarni:1987:IEF

[1108] K. G. Kulkarni and Malcolm P. Atkinson. Implementing an extended functional data model using PS-Algol. Software—Practice and Experience, 17(3):171–185, March 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meertens:1987:ARP

[1109] Lambert Meertens. An abstract reader prepared for IFIP WG 2.1. Dummy note., January 1987.

Robison:1987:IFP

[1110] A. D. Robison. The Illinois functional programming interpreter. ACM SIGPLAN Notices, 22(7):64–73, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/proceedings/plan/29650/p64-
robison/.

Rusch:1987:ERM

[1111] J. J. Rusch. Enkele resultaten met de schema’s van Roe en Steger & Warming bij transsone stromingen in een MG/eindige-volume methode. (Dutch) [Some results of the Roe and Steger & Warming schemes for transonic flows in a MG finite volume method]. Dummy note., January

Tennent:1987:QAL

[1112] R. D. Tennent. Quantification in Algol-like languages. Information Processing Letters, 25(2):133–137, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yemini:1987:ATE

[1113] Shaula Yemini and Daniel M. Berry. An axiomatic treatment of exception handling in an expression-oriented language. ACM Transactions on Programming Languages and Systems, 9(3):390–407, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http:

//www.acm.org/pubs/toc/Abstracts/0164-0925/24052.html. See re-

marks [1128].

Zakharov:1987:MCS

[1114] L. A. Zakharov, S. B. Pokrovsky, G. G. Stepanov, and S. V. Ten. A Multilanguage Compiling System. Computing Center of the Siberian Division of the USSR Academy of Sciences, Novosibirsk, USSR, 1987. ???? pp.

Anonymous:1988:MKS

[1115] Anonymous. Mark Kac seminar on probability and physics: syllabus: [seminar held during the academic year 1985–1987 at Amsterdam]. C, January 1988. ?? pp.

Anonymous:1988:SSP

[1116] Anonymous. STATAL: statistical procedures in Algol 60, Part 1. C, January 1988. ?? pp.

Anonymous:1988:VCV

[1117] Anonymous. Vakantiecursus 1988 Centrum voor Wiskunde en Informatica: differentierekening. (Dutch) [Holiday course 1988 Center for Mathematics and Computer Science: difference calculus]. C, January 1988. ?? pp.

Dearle:1988:SBS

[1118] A. Dearle and A. L. Brown. Safe browsing in a strongly typed persistent environment. The Computer Journal, 31(6):540–544, December 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/31/6/540.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_31/
Issue_06/tiff/540.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_06/tiff/541.tif;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_31/Issue_06/tiff/542.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_06/tiff/
543.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_06/tiff/544.tif.

deRooy:1988:CES

[1119] M. de Rooy. Concurrent evaluation of side-effects. Dummy note., January 1988.

Fogiel:1988:HGC

[1120] Max Fogiel. Handbook and guide for comparing and selecting computer languages BASIC, FORTRAN, PASCAL, COBOL, PL/1, APL, ALGOL-60, C. Research and Education Association, New York, NY, USA, 1988. ISBN ???? v + 122 pp. LCCN ????

Hemker:1988:DCN

[1121] Piet Hemker and Barry Koren. Defect correction and nonlinear multigrid for the steady Euler equations. Dummy note., January 1988.

Meyer:1988:TFA

[1122] A. R. Meyer and K. Sieber. Towards fully abstract semantics for local variables. In ACM [1297], pages 191–203. ISBN ???? LCCN

???? URL http://www.acm.org:80/pubs/citations/proceedings/
plan/73560/p191-meyer/.

Morgan:1988:PPA

[1123] Carroll Morgan. Procedures, parameters, and abstraction: separate concerns. Science of Computer Programming, 11(1):17–28 (or 17–27??), October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nijholt:1988:BPA

[1124] Anton (Antinus) Nijholt. BNF Programming, Analyzing, and Compiling, chapter 8, pages 201–237. North-Holland, Amsterdam, The Netherlands, 1988. ISBN 0-444-70463-9. LCCN QA76.6 .N55

1988.  URL https://www.sciencedirect.com/science/chapter/
bookseries/pii/B9780444704634500153.

Philbrow:1988:PAD

[1125] P. C. Philbrow, I. Armour, M. P. Atkinson, and J. Livingstone. PS-Algol’s device-independent output statement. ACM SIGPLAN Notices, 23(6):107–114, June 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderHorst:1988:SSP

[1126] Rob van der Horst and Richard Gill. STATAL: statistical procedures in Algol 60, Part 3. C, January 1988. ?? pp.

vanWijngaarden:1988:DRA

[1127] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Draft report on the algorithmic language ALGOL 68. Report MR 93, Mathematisch Centrum, University of Amsterdam, Amsterdam, The Netherlands, March 1988. ?? pp. URL https://archive.

computerhistory.org/resources/text/algol/algol_bulletin/AS26/

. Second printing.

Yemini:1988:TCA

[1128] Shaula Yemini and Daniel M. Berry. Technical correspondence: “An Axiomatic Treatment of Exception Handling in an Expression-Oriented Language”. ACM Transactions on Programming Languages and Systems, 10(3):503–504, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [1113].

Zaki:1988:ARM

[1129] M. Zaki and M. M. Elboraey. Analysis of reliability models for interconnecting MIMD systems. The Computer Journal, 31(4):304–312, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/31/4/304.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/304.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/305.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/306.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_31/
Issue_04/tiff/307.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_31/Issue_04/tiff/309.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/
310.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
31/Issue_04/tiff/311.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_31/Issue_04/tiff/312.tif.

Anonymous:1989:PVS

[1130] Anonymous. Proceedings van het symposium wiskunde en de computer: [Amsterdam, 1989]. (Dutch) [Proceedings of the symposium mathematics and the computer: [Amsterdam, 1989]]. C, January 1989. ?? pp.

Anonymous:1989:VCV

[1131] Anonymous. Vakantiecursus 1989 Centrum voor Wiskunde en Informatica: wiskunde in de Gouden Eeuw. (Dutch) [Holiday course 1989 Center for Mathematics and Computer Science: mathematics in the Golden Age]. C, January 1989. ?? pp.

Rapin:1989:PON

[1132] Ch. Rapin. Procedural objects in Newton. ACM SIGPLAN Notices, 24(9):133–141, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tennent:1989:EDS

[1133] R. D. Tennent. Elementary data structures in ALGOL-like languages. Science of Computer Programming, 13(1):73–110, December 1989. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Merrett:1990:SOC

[1134] T. H. Merrett. Scoping is not orthogonal to completeness. The Computer Journal, 33(4):381, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_33/Issue_04/tiff/381.tif.

Nawrocki:1990:DOA

[1135] J. R. Nawrocki and C. H. A. Koster. On display optimization for Algollike languages. Computer Languages, 15(1):27–39, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Philbrow:1990:EEH

[1136] P. C. Philbrow and M. P. Atkinson. Events and exception handling in PS-algol. The Computer Journal, 33(2):108–125, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/33/2/108.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_
02/tiff/108.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_33/Issue_02/tiff/109.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/110.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/111.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_33/
Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/113.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_33/Issue_02/tiff/114.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/

115.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/117.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/118.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_33/
Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_02/tiff/120.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_33/Issue_02/tiff/121.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/
122.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
33/Issue_02/tiff/123.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_33/Issue_02/tiff/124.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_02/tiff/125.
tif.

Robson:1990:RAM

[1137] J. M. Robson. Random access machines with multi-dimensional memories. Information Processing Letters, 34(5):265–266, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stancin:1990:BRS

[1138] Marina Stancin. Book review: STATAL: statistical procedures in Algol 60 by R. Van der Hor. Bulletin mathématique de la Société des Sciences Mathématiques de Roumanie, 34 (82)(3):288, 1990. ISSN 1220-3874,

2065-0264. URL http://www.jstor.org/stable/43678351.

Anonymous:1991:SCV

[1139] Anonymous. Studiedag cultuurgeschiedenis van de automatisering [GMFW-CWI colloquium history of computing]: [26.04.1991]. (Dutch) [Study day on the cultural history of automation [GMFW-CWI colloquium history of computing]: [26.04.1991]]. Dummy note., January 1991.

Consel:1991:SDS

[1140] Charles Consel and Olivier Danvy. Static and dynamic semantics processing. In ACM [1298], pages 14–24. ISBN ???? LCCN

???? URL http://www.acm.org:80/pubs/citations/proceedings/
plan/99583/p14-consel/.

Khanna:1991:LPS

[1141] S. Khanna. Logic programming for software verification and testing. The Computer Journal, 34(4):350–357, August 1991. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/34/4/350.full.pdf+html;

http://www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_
04/tiff/350.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/351.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_34/Issue_04/tiff/352.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/353.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_34/
Issue_04/tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_34/Issue_04/tiff/355.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_34/Issue_04/tiff/356.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_34/Issue_04/tiff/
357.tif.

Tennent:1991:CPW

[1142] R. D. Tennent and J. K. Tobin. Continuations in possible-world semantics. Theoretical Computer Science, 85(2):283–303, August 12, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tennent:1991:LVN

[1143] R. D. Tennent. Local variables and non-interference in Algol-like languages. Lecture Notes in Computer Science, 530:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1992:MKS

[1144] Anonymous. Mark Kac seminar on probability and physics: syllabus [a selection of reports of lectures delivered during the academic years] 1987– 1992. C, January 1992. ?? pp.

Anonymous:1992:PSL

[1145] Anonymous. Proceedings seminar 1986–1987: lectures on Kac–Moody algebras [presented in the seminar Mathematical structures in field theories’]. C, January 1992. ?? pp.

Anonymous:1992:VCV

[1146] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 1992: systeemtheorie. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 1992: systems theory]. C, January 1992. ?? pp.

Boute:1992:EDF

[1147] Raymond T. Boute. The Euclidean definition of the functions div and mod. ACM Transactions on Programming Languages and Systems, 14(2): 127–144, April 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-

4593 (electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/128862.html.

Koster:1992:SHA

[1148] C. H. A. Koster. A shorter history of Algol68. Web archive document,

1992 (??). URL https://web.archive.org/web/20060810103448/
http://npt.cc.rsu.ru/user/wanderer/ODP/ALGOL68.txt.

Lambert:1992:AOC

[1149] Kenneth A. Lambert. An ALGOL object code interpreter in scheme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24(3):39–44, September 1992. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meertens:1992:ASE

[1150] Lambert Meertens, Steven Pemberton, and Guido van Rossum. The ABC structure editor: structure-based editing for the ABC programming environment. Dummy note., January 1992.

Needham:1992:LDC

[1151] Roger M. Needham. Later developments at Cambridge: Titan, CAP, and the Cambridge Ring. IEEE Annals of the History of Computing, 14(4):57–58, October/December 1992. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/dlib.computer.org/an/books/an1992/pdf/a4057.pdf;   http://
www.computer.org/annals/an1992/a4057abs.htm.

Stansifer:1992:CE

[1152] Ryan Stansifer. The calculation of Easter. ACM SIGPLAN Notices, 27 (12):61–65, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also [1176].

vanAsch:1992:HHM

[1153] A. G. van Asch and F. van der Blij. Hoeken en hun maat. (Dutch) [Angles and their measure]. C, January 1992. ?? pp.

Alberts:1993:ICH

[1154] Gerard Alberts. Introduction to the conference [on the history of ALGOL 68]. In Alberts [1301], pages 5–8. URL https://ir.cwi.nl/pub/10854/ 10854D.pdf. Historical Note AM-NH9301.

Alberts:1993:PC

[1155] Gerard Alberts. Program of the conference. In Alberts [1301], page 73.

URL https://ir.cwi.nl/pub/10854/10854D.pdf. Historical Note

AM-NH9301.

Anonymous:1993:MVK

[1156] Anonymous. Meetkunde, van kunst tot kunde. (Dutch) [Geometry: from art to science]. C, January 1993. ?? pp.

Anonymous:1993:VCV

[1157] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 1993: het reele getal. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 1993: the real number]. C, January 1993. ?? pp.

Bauer:1993:HPL

[1158] Friedrich L. Bauer. History of programming languages, a survey. In Al-

berts [1301], pages 9–22. URL https://ir.cwi.nl/pub/10854/10854D.

pdf. Historical Note AM-NH9301.

Bowlden:1993:IA

[1159] Henry J. Bowlden. An introduction to ALGOL 68. ACM SIGPLAN Notices, 28(3):345–346, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.acm.org:80/pubs/citations/proceedings/plan/154766/p345-
bowlden/.

Bulyonkov:1993:AYU

[1160] Mikhail A. Bulyonkov, Alexandr F. Rar, and Andrei N. Terekhov. Algol 68 — 25 years in the USSR. In Alberts [1301], pages 45–51. URL https:

//ir.cwi.nl/pub/10854/10854D.pdf. Historical Note AM-NH9301.

Guttag:1993:BIC

[1161] John Guttag. A brief introduction to CLU. ACM SIGPLAN Notices, 28 (3):351–352, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.org:

80/pubs/citations/proceedings/plan/154766/p351-guttag/.

Jervis:1993:PPL

[1162] Robert Jervis. The Parasol programming language. Dr. Dobb’s Journal of Software Tools, 18(11):34–41, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Koster:1993:MA

[1163] Cees Koster. The making of Algol 68. In Alberts [1301], pages 27–

39. URL https://ir.cwi.nl/pub/10854/10854D.pdf. Historical Note

AM-NH9301.

Lee:1993:BIP

[1164] Peter Lee. A brief introduction to Pascal. ACM SIGPLAN Notices, 28 (3):363–364, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.org:

80/pubs/citations/proceedings/plan/154766/p363-lee/.

Lindsey:1993:HA

[1165] C. H. Lindsey. A history of ALGOL 68. ACM SIGPLAN Notices, 28(3): 97–132, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://algol68-lang.org/

docs/lindsey-history-of-algol68.pdf   http://www.acm.org:80/
pubs/citations/proceedings/plan/154766/p97-lindsey/.

Lindsey:1993:HAE

[1166] Charles Lindsey. The history of ALGOL 68 (extended abstract). In

Alberts [1301], pages 23–26. URL https://ir.cwi.nl/pub/10854/

10854D.pdf. Historical Note AM-NH9301.

Meertens:1993:DEL

[1167] Lambert Meertens. The design of elegant languages. In Alberts [1301],

pages 53–64. URL https://ir.cwi.nl/pub/10854/10854D.pdf. His-

torical Note AM-NH9301.

OHearn:1993:RPL

[1168] P. W. O’Hearn and R. D. Tennent. Relational parametricity and local variables. In ACM [1300], pages 171–184. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/pubs/citations/proceedings/
plan/158511/p171-o_hearn/. ACM order number 549930.

Peck:1993:AVW

[1169] John Peck. Aad van Wijngaarden and the Mathematisch Centrum, a personal recollection. In Alberts [1301], pages 65–71. URL https://ir.

cwi.nl/pub/10854/10854D.pdf. Historical Note AM-NH9301.

Plauger:1993:PLG

[1170] P. J. Plauger. Programming language guessing games. Dr. Dobb’s Journal of Software Tools, 18(11):16–22, October 1993. CODEN DDJOEB. ISSN 1044-789X.

Schwarz:1993:BIC

[1171] Jerry Schwarz. A brief introduction to C++. ACM SIGPLAN Notices, 28(3):349–350, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.org:

80/pubs/citations/proceedings/plan/154766/p349-schwarz/.

vanderMeulen:1993:OFP

[1172] Sietse van der Meulen. An orthogonal first programming language. In Alberts [1301], pages 41–44. URL https://ir.cwi.nl/pub/10854/ 10854D.pdf. Historical Note AM-NH9301.

Weeks:1993:OAP

[1173] Stephen Weeks and Matthias Felleisen. On the orthogonality of assignments and procedures in Algol. In ACM [1300], pages 57–70. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN

QA76.7 .A15 1993. URL http://www.acm.org:80/pubs/citations/
proceedings/plan/158511/p57-weeks/. ACM order number 549930.

Wilkes:1993:CPF

[1174] Maurice V. Wilkes. Computing perspectives: From Fortran and Algol to object-oriented languages. Communications of the ACM, 36(7):21–23, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/0001-0782/
159553.html.

Alberts:1994:WPH

[1175] Gerard Alberts. Wiskunde en praktijk in historisch perspectief: syllabus. (Dutch) [Mathematics and practice in historical perspective: syllabus]. C, January 1994. ?? pp.

Bevan:1994:LEC

[1176] Stephen J. Bevan. Letter to the editor: SIGPLAN Notices 27(12) 1992: “The Calculation of Easter”. ACM SIGPLAN Notices, 29(9):8, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1152].

Boxma:1994:PEP

[1177] Onno Boxma and Ger Koole. Performance evaluation of parallel and distributed systems: solution methods: proceedings of the third QMIPS [Quantitative Modeling in Parallel Systems] workshop, [Torino, Italy, 25-26.09.1993], part 2. C, January 1994. ?? pp.

Sieber:1994:FAS

[1178] K. Sieber. Full abstraction for the second order subset of an ALGOLlike language. Lecture Notes in Computer Science, 841:608–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:VKK

[1179] Anonymous. Vakantiecursus 1995: kegelsneden en kwadratische vormen. (Dutch) [Holiday course 1995: conic sections and quadratic shapes]. C, January 1995. ?? pp.

Brookes:1995:CTT

[1180] S. Brookes. A category-theoretic treatment of a parallel Algol-like language. Lecture Notes in Computer Science, 953:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHearn:1995:PLV

[1181] P. W. O’Hearn and R. D. Tennent. Parametricity and local variables. Journal of the ACM, 42(3):658–709, May 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/210425.html.

Reynolds:1995:UFC

[1182] John C. Reynolds. Using functor categories to generate intermediate code. In ACM [1302], pages 25–36. ISBN 0-89791-692-1.

LCCN QA 76.7 A11 1995. URL http://www.acm.org:80/pubs/
citations/proceedings/plan/199448/p25-reynolds/. ACM order

number: 549950.

Walters:1995:MEL

[1183] Pum Walters and Jasper Kamperman. A model for I/O in equational languages with don’t care non-determinism. Dummy note., January 1995.

Anonymous:1996:PSM

[1184] Anonymous. Proceedings seminar 1989–1990 mathematical structures in field theories. C, January 1996. ?? pp.

Anonymous:1996:VC

[1185] Anonymous. Vakantiecursus 1996: chaos. (Dutch) [Holiday course 1996: chaos]. C, January 1996. ?? pp.

Barnes:1996:WDK

[1186] J. Barnes. We don’t know nothing [programming languages]. Ada User Journal, 17(4):168–171, December 1996. CODEN AUJOET. ISSN 1381-6551.

Bulyonkov:1996:PAS

[1187] Mikhail A. Bulyonkov and Dmitry V. Kochetov. Practical aspects of specialization of Algol-like programs. Lecture Notes in Computer Science, 1110:17–32, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL https://link.springer.com/chapter/10.1007/3-
540-61580-6_2.

Dybvig:1996:SPL

[1188] R. Kent Dybvig. The SCHEME Programming Language: ANSI Scheme. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1996. ISBN 0-13-454646-6. xii + 248 pp. LCCN QA76.73.S34 D93 1996. URL

http://www.scheme.com/tspl2d/.

Koster:1996:MA

[1189] C. H. A. Koster. The making of Algol 68. Lecture Notes in Computer Science, 1181:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHearn:1996:NAC

[1190] Peter W. O’Hearn. Note on Algol and conservatively extending functional programming. Journal of Functional Programming, 6(1):171– 180, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-

7653 (electronic). URL https://www.cambridge.org/core/product/
050F6652EB31D730ECC20A16E5B0E8EE.

Reynolds:1996:DPL

[1191] John C. Reynolds. Design of the programming language Forsythe. Report CMU-CS-96-146 (AD-A311094), School of Computer Science, Carnegie Mellon University, Pittsburgh, PA 15213, USA, June 28, 1996.

ii + 68 pp. URL https://apps.dtic.mil/sti/pdfs/ADA311094.pdf.

Reprinted in Algol-like Languages, edited by P. O’Hearn and R. D. Tennent, Birkhäuser, 1996.

Sieber:1996:FAS

[1192] Kurt Sieber. Full abstraction for the second order subset of an ALGOLlike language. Theoretical Computer Science, 168(1):155–212, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/store/
tcs/cas_sub/browse/browse.cgi?year=1996&volume=168&issue=1&
aid=2268.

Anonymous:1997:VCV

[1193] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 1997: rekenen op het toeval. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 1997: counting on chance]. C, January 1997. ?? pp.

Naur:1997:RRA

[1194] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Kafz, J. McCarthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language Algol 60. In Algol-like Languages, pages 19–49. Birkhäuser Boston Inc., Cambridge, MA, USA, 1997.

Pescio:1997:FWN

[1195] Carlo Pescio. A few words with Niklaus Wirth. Software Development,

5(6):??, June 1997. URL http://pascal.hansotten.com/uploads/
wirth/interview%20wirth%20June%201997%20carlio%20pescio.txt.

Abramsky:1998:LSS

[1196] S. Abramsky and G. McCusker. Linearity, sharing and state: a fully abstract game semantics for idealized Algol with active expressions. Theoretical Computer Science, 197(1–2):241–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1998:VCV

[1197] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 1998: Meetkunde, oud en nieuw. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 1998: Geometry, old and new]. C, January 1998. ?? pp.

Hamlet:1998:WCW

[1198] Dick Hamlet. What can we learn by testing a program? ACM SIG-SOFT Software Engineering Notes, 23(2):50–52, March 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Abramsky:1999:FAI

[1199] Samson Abramsky and Guy McCusker. Full abstraction for Idealized Algol with passive expressions. Theoretical Computer Science, 227 (1–2):3–42, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.nl/gej-

ng/10/41/16/151/17/17/abstract.html;   http://www.elsevier.
nl/gej-ng/10/41/16/151/17/17/article.pdf.

Anonymous:1999:VCV

[1200] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 1999: onbewezen vermoedens. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 1999: unproven suspicions]. C, January 1999. ?? pp.

Roeckl:1999:CPS

[1201] C. Roeckl and D. Sangiorgi. A λ-calculus process semantics of concurrent idealised ALGOL. Lecture Notes in Computer Science, 1578:306–321, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thielecke:1999:CFJ

[1202] Hayo Thielecke. Continuations, functions and jumps. ACM SIGACT News, 30(2):33–42, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL http://www.cs.bham.ac.uk/~

hxt/research/Logiccolumn8.ps.

Dzhaparidze:2000:IOP

[1203] Kacha Dzhaparidze. Introduction to option pricing in a securities market. C, January 2000. ?? pp.

OHearn:2000:APL

[1204] Peter W. O’Hearn and John C. Reynolds. From Algol to polymorphic linear lambda-calculus. Journal of the ACM, 47(1):167–223, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-1/p167-o_hearn/.

Plauger:2000:JAV

[1205] P. J. Plauger. Just another version of Algol. C/C++ Users Journal, 18 (1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:VCV

[1206] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 2001: Experimentele wiskunde. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 2001: Experimental mathematics]. C, January 2001. ?? pp.

Clark:2002:IFA

[1207] David Clark, Chris Hankin, and Sebastian Hunt. Information flow for Algol-like languages. Computer Languages, Systems and Structures, 28 (1):3–28, April 2002. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Markoff:2002:EDP

[1208] John Markoff. Edsger Dijkstra, 72, physicist who shaped computer era, dies. New York Times, ??(??):??, August 10, 2002. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

www.nytimes.com/2002/08/10/obituaries/10DIJK.html.

Randell:2002:ASE

[1209] Brian Randell. From Algol 60 to software engineering. Talk at Dijkstra Tribute, Sun Microsystems, London, UK., November 14, 2002.

Ritchie:2002:FLL

[1210] Dennis M. Ritchie. Five little languages and how they grew: Talk at HOPL. Report ??, Bell Laboratories, Murray Hill, NJ 07974, USA,

March 19, 2002. ?? pp. URL https://www.bell-labs.com/usr/dmr/

www/hopl.html. This talk was presented at the Second History of Programming Languages conference, Cambridge, Mass., April, 1993.

Sebesta:2002:CPL

[1211] Robert W. Sebesta. Concepts of Programming Languages. Addison-Wesley, Reading, MA, USA, fifth edition, 2002. ISBN 0-201-75295-6, 0-321-17645-6. xiii + 698 pp. LCCN QA76.7 .S43 2002.

Anonymous:2003:VCV

[1212] Anonymous. Vakantiecursus [Centrum voor Wiskunde en Informatica] 2003: wiskunde in het dagelijks leven. (Dutch) [Holiday course [Center for Mathematics and Computer Science] 2003: mathematics in everyday life]. C, January 2003. ?? pp.

Aretz:2003:DZAa

[1213] F. E. J. Kruseman Aretz. The Dijkstra–Zonneveld ALGOL 60 compiler for the Electrologica X1. Dummy note., January 2003.

Aretz:2003:DZAb

[1214] F. E. J. Kruseman Aretz. The Dijkstra–Zonneveld ALGOL 60 compiler for the Electrologica X1 historical note SEN, 2. ????, ????, June 2003. ?? pp.

Knuth:2003:SPC

[1215] Donald E. Knuth. Selected Papers on Computer Languages, volume 139 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-381-2 (hardcover), 1-57586-382-0 (paperback). xvi + 594 pp. LCCN QA76.7 .K63 2002.

KrusemanAretz:2003:DZA

[1216] F. E. J. Kruseman Aretz. The Dijkstra–Zonneveld ALGOL 60 compiler for the Electrologica X1: historical note SEN, 2. Note SEN-N0301, Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, June

30, 2003. vii + 328 pp. URL http://archive.computerhistory.org/
resources/text/algol/algol_bulletin/EX/SN-N0301.PDF.

Windhouwer:2003:FGS

[1217] Menzo A. indhouwer. Feature Grammar Systems — Incremental Maintenance of Indexes to Digital Media Warehouses. Ph.D. thesis, Faculty of Science (FNWI) — Informatics Institute (IVI), University of Amsterdam, Amsterdam, The Netherlands, November 2003. viii + 191 pp.

URL http://dare.uva.nl/document/86747;   http://hdl.handle.
net/11245/1.214540.

Haigh:2004:IBF

[1218] Thomas Haigh. An interview with Brian Ford, Steve Hague and Sven Hammarling. Web site, June 29–30, 2004. URL https://history. siam.org/pdfs2/ford_final.pdf. Interview conducted for the Society for Industrial and Applied Mathematics, and the transcript and original tapes were later donated to the Computer History Museum in Mountain View, CA.

Abraham:2005:APS

[1219] Erika Abraham. An Assertional Proof System for Multithreaded Java — Theory and Tool Support. Ph.D. thesis, ????, ????, January 2005. ?? pp.

URL http://hdl.handle.net/1887/584.

Vinju:2005:ATS

[1220] Jurgen Vinju. Analysis and transformation of source code by parsing and rewriting. Ph.D. thesis, Faculty of Science (FNWI) — Informatics Institute (IVI), University of Amsterdam, Amsterdam, The Netherlands,

November 15, 2005. ?? pp. URL http://hdl.handle.net/11245/1.
252129.

Anonymous:2006:PFS

[1221] Anonymous. Proceedings of the fifty-second European study group with industry, Amsterdam, The Netherlands, 31.01-04.02.2005. C, January 2006. ?? pp.

Anonymous:2006:VCV

[1222] Anonymous. Vakantiecursus 2006 Centrum voor Wiskunde en Informatica: Actuele wiskunde, Amsterdam en Eindhoven, 2006. (Dutch) [Holiday course 2006 Center for Mathematics and Computer Science: Current mathematics, Amsterdam and Eindhoven, 2006]. C, January 2006. ?? pp.

Economopoulos:2006:GLP

[1223] Giorgos Robert Economopoulos. Generalised LR Parsing Algorithms. Ph.D. thesis, Royal Holloway, London, UK, January 2006. ?? pp. URL

https://www.proquest.com/pqdtglobal/docview/301691838/.

Herder:2006:MSP

[1224] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Modular system programming in MINIX 3. ;login: the USENIX Association newsletter, 31(2):19–28, April 2006. CODEN LOGNEM. ISSN 1044-6397. URL http://www.minix3.

org/;   http://www.minix3.org/docs/login-2006.pdf;   http:/
/www.usenix.org/publications/login/2006-04/openpdfs/herder.
pdf.

Barnes:2007:SIBb

[1225] John G. P. Barnes. SP1: introducing the best of Ada 2005. ACM SIGADA Ada Letters, 27(3):3, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Hermida:2007:FFP

[1226] C. Hermida and R. D. Tennent. A fibrational framework for possibleworld semantics of Algol. Theoretical Computer Science, 375(1–3):3–19, May 1, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wirth:2007:MO

[1227] Niklaus Wirth. Modula-2 and Oberon. In Proceedings of the Third ACM SIGPLAN Conference on History of Programming Languages, HOPL III, pages 3.1–3.10. ACM Press, New York, NY 10036, USA, 2007. ISBN

1-59593-766-8. URL https://people.inf.ethz.ch/wirth/Articles/
Modula-Oberon-June.doc.

Acs:2008:CAB

[1228] Éva Ács, A. K. Borsodi, and Gy. Záray. Comparative algological and bacteriological examinations on biofilms developed on different substrata in a shallow soda lake. Aquatic Ecology, 42(4):521–531, December 2008. CODEN AQECF9. ISSN 1386-2588 (print), 1573-5125 (electronic). URL

https://link.springer.com/article/10.1007/s10452-007-9132-0.

Anonymous:2008:VWP

[1229] Anonymous. Vakantiecursus 2008 — Wiskunde en profil: het gezicht van de wiskunde. (Dutch) [Holiday course 2008 — Mathematics and profile: the face of mathematics]. C, August 2008. ?? pp.

Lampson:2008:LSE

[1230] Butler W. Lampson. Lazy and speculative execution in computer systems. ACM SIGPLAN Notices, 43(9):1–2, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Murawski:2008:TOI

[1231] Andrzej S. Murawski and Igor Walukiewicz. Third-order Idealized Algol with iteration is decidable. Theoretical Computer Science, 390(2–3):214– 229, January 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ahmed:2009:SDR

[1232] Amal Ahmed, Derek Dreyer, and Andreas Rossberg. State-dependent representation independence. ACM SIGPLAN Notices, 44(1):340–353, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2009:VTU

[1233] Anonymous. Vakantiecursus 2009 — Tel uit je winst — wiskunde in geld en spelen. (Dutch) [Holiday course 2009 — Count your winnings — mathematics in money and games]. C, January 2009. ?? pp.

Apt:2009:VSC

[1234] Krzysztof Apt, Ernst-Rüdiger Olderog, and Frank de Boer. Verification of Sequential and Concurrent Programs. S, October 2009. ?? pp.

KrusemanAretz:2010:DCP

[1235] F. E. J. Kruseman Aretz. Design and correctness proof of an emulation of the floating-point operations of the Electrologica X8: a case study. Computer Science Report 1002, Technische Universiteit Eindhoven, Eindhoven, The Netherlands, March 30, 2010. 58 pp. URL http:

//repository.tue.nl/674735.

Nofre:2010:UAU

[1236] David Nofre. Unraveling Algol: US, Europe, and the creation of a programming language. IEEE Annals of the History of Computing, 32(2):58–68, April/June 2010. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https://www.researchgate.net/

profile/David_Nofre/publication/224107164_Unraveling_Algol_
US_Europe_and_the_Creation_of_a_Programming_Language.

Anonymous:2011:VS

[1237] Anonymous. Vakantiecursus 2011 — Symmetrie. (Dutch) [Holiday course 2011 — Symmetry]. C, January 2011. ?? pp.

Bauer:2011:AVG

[1238] Friedrich L. Bauer. Die ALGOL-Verschwörung. (German) [The ALGOL conspiracy]. Informatik Spektrum, 35(2):141–149, December 2011. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Gil:2011:BMC

[1239] Amparo Gil, Javier Segura, and Nico Temme. Basic Methods for Computing Special Functions, chapter ??, page ?? S, January 2011.

Stoer:2011:GST

[1240] Josef Stoer. The Gatlinburg Symposia and their influence on the development of numerical linear algebra. IMAGE, ??(??):13–25,

Spring 2011. URL http://www.maths.manchester.ac.uk/~higham/
conferences/householder/image46_stoer.pdf. From the start of §2:

“The Gatlinburg Symposia, officially started by Alston S. Householder in 1961, were influenced in part by the existence of two earlier and similar symposia, the first organized by Olga Taussky in 1951 in Los Angeles and the second by Wallace Givens in 1957 in Detroit. ”.

Zaytsev:2011:UFL

[1241] Vadim Zaytsev and Ralf Lämmel. A unified format for language documents. In Brian Malloy, Steffen Staab, and Mark van den Brand, editors, Software Language Engineering Third International Conference, SLE 2010, Eindhoven, The Netherlands, October 12–13, 2010, Revised Selected Papers, pages 206–225. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., December 2011. ISBN 3-642-19439-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .S54 2011.

Bauer:2012:AVG

[1242] Friedrich L. Bauer. Die ALGOL-Verschwörung. (German) [The ALGOL conspiracy]. Informatik Spektrum, 35(2):141–149, December 2012. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic). URL

http://link.springer.com/article/10.1007/s00287-011-0585-0.

Grassmann:2012:CBR

[1243] Winfried K. Grassmann. A computer built with relays and a mechanical memory, and Algol. The Computer Journal, 55(11):1309–1316, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/55/11/
1309.full.pdf+html.

Braden:2013:BAS

[1244] Robert Braden. Burroughs Algol at Stanford University, 1960–1963. IEEE Annals of the History of Computing, 35(4):69–73, October/ December 2013. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). From page 72: “This Anecdote is dedicated to the memory of George Forsythe, whose vision and gentle wisdom initiated and steered the revolution I have described. Between 1960 and 1963, he created at Stanford an environment in which a Burroughs B220 computer system, the Burroughs Algebraic Compiler, and the BALGOL language combined to excite and train a group of Stanford people diving into the emerging computer age.”.

Rutishauser:2013:NAS

[1245] Hanna Rutishauser. Numerik, ALGOL und die Schweizer Hochalpen. Zur Arbeit an der Biografie von Heinz Rutishauser (1918–1970). (German) [Numerics, ALGOL and the Swiss High Alps. On the work on the biography of Heinz Rutishauser (1918–1970)]. Informatik Spektrum, 36(5):??, ???? 2013. CODEN INSKDW. ISSN 0170-6012 (print), 1432-122X (elec-

tronic). URL http://link.springer.com/article/10.1007/s00287-
013-0730-z.

Alberts:2014:ACP

[1246] Gerard Alberts. Algol culture and programming styles [guest editor’s introduction]. IEEE Annals of the History of Computing, 36(4):2–5, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Alberts:2014:UVL

[1247] G. Alberts and E. G. Daylight. Universality versus locality: The Amsterdam style of Algol implementation. IEEE Annals of the History of Computing, 36(4):52–63, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durnova:2014:EAE

[1248] H. Durnova. Embracing the Algol effort in Czechoslovakia. IEEE Annals of the History of Computing, 36(4):26–37, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Durnova:2014:TPW

[1249] H. Durnova and G. Alberts. Think piece: Was Algol 60 the first algorithmic language? IEEE Annals of the History of Computing, 36(4):104, 102–103, October/December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mounier-Kuhn:2014:AFU

[1250] P. Mounier-Kuhn. Algol in France: From universal project to embedded culture. IEEE Annals of the History of Computing, 36(4):6–25, October/ December 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Terekhov:2014:AIU

[1251] Andrey Terekhov. ALGOL 68 and its impact on the USSR and Russian programming. In Tomilin [1304], pages 97–106. ISBN 1-4799-1799-0, 1-4799-1800-8. LCCN QA75.5 .I578 2014. URL https://sorucom.iis.

nsk.su/files/news/sorucom2014.pdf.

Gruner:2015:TOI

[1252] Andreas Grüner. Testing Object Interaction. Ph.D. thesis, ????, ????,

December 2015. ?? pp. URL http://hdl.handle.net/1887/16243.

Rot:2015:EC

[1253] Jurriaan Rot. Enhanced coinduction. Ph.D. thesis, ????, ????, October

2015. ?? pp. URL http://hdl.handle.net/1887/35814.

Haigh:2016:HRW

[1254] Thomas Haigh and Mark Priestley. Historical reflections: Where code comes from: architectures of automatic control from Babbage to Algol. Communications of the ACM, 59(1):39–44, January 2016. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//cacm.acm.org/magazines/2015/1/195721/fulltext.

Osvald:2016:GGT

[1255] Leo Osvald, Grégory Essertel, Xilun Wu, Lilliam I. González Alayón, and Tiark Rompf. Gentrification gone too far? Affordable 2nd-class values for fun and (co-)effect. ACM SIGPLAN Notices, 51(10):234–251, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higham:2018:UN

[1256] Nicholas J. Higham. The unwinding number. SIAM News, 51(8):2, Octo-

ber 2018. ISSN 0036-1437. URL https://sinews.siam.org/Details-
Page/the-unwinding-number.

Bright:2020:ODP

[1257] Walter Bright, Andrei Alexandrescu, and Michael Parker. Origins of the D programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):73:1–73:38, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/3386323.

Brock:2020:DDR

[1258] David C. Brock. Discovering Dennis Ritchie’s lost dissertation. Web

site, June 19, 2020. URL https://computerhistory.org/blog/
discovering-dennis-ritchies-lost-dissertation/.

Mody:2020:BRS

[1259] Cyrus C. M. Mody. Book review: Subrata Dasgupta. The Second Age of Computer Science: From Algol Genes to Neural Nets. Isis, 111(2):439– 440, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nofre:2021:PEP

[1260] David Nofre. The politics of early programming languages: IBM and the Algol Project. Historical Studies in the Natural Sciences, 51(3):379–??, June 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.ucpress.edu/hsns/article/51/3/
379/117357/The-Politics-of-Early-Programming-LanguagesIBM-and.

Korner:2022:FYP

[1261] Philipp Körner, Michael Leuschel, João Barbosa, Vítor Santos Costa, Verónica Dahl, Manuel Hermenegildo, Jose Morales, Jan Wielemaker, Daniel Diaz, Salvador Abreu, and Giovanni Ciatto. Fifty years of prolog and beyond. Theory and Practice of Logic Programming, 22(6):776–858, March 2022.

Castellan:2023:GCM

[1262] Simon Castellan and Pierre Clairambault. The geometry of causality: Multi-token geometry of interaction and its causal unfolding. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 24:1–24:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic).

URL https://dl.acm.org/doi/10.1145/3571217.

Reinsch:2023:SVD

[1263] Christian Reinsch and Mathias Richter. Singular value decomposition in extended double precision arithmetic. Numerical Algorithms, 93(3): 1137–1155, July 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01459-9.

Anonymous:2025:AFE

[1264] Anonymous. An Algol 68 front end for GCC. Web site, January 1, 2025.

URL https://lwn.net/Articles/1003916/.

Marchesi:2025:AFE

[1265] Jose E. Marchesi. Algol 68 front end. Web site, 2025. URL https://

gcc.gnu.org/wiki/Algol68FrontEnd.

Marchesi:2025:GAC

[1266] Jose E. Marchesi. The GNU Algol 68 Compiler. Free Software Foundation, Inc., Cambridge, MA, USA, 2025. iv + 74 pp.

McJones:2026:HA

[1267] Paul McJones. History of ALGOL. Web site, May 2, 2026. URL https:

//softwarepreservation.computerhistory.org/ALGOL/. This useful

site has clearly been in progress for some time, and contains many links to publications about the development of the Algol language family.

SPG:2026:AID

[1268] Software Preservation Group. ALGOL 68 implementations and dialects. Computer History Museum Web site, May 2, 2026. URL https://

softwarepreservation.computerhistory.org/ALGOL/algol68impl.
html.

SPG:2026:ALT

[1269] Software Preservation Group. Applications, libraries, and test suites. Computer History Museum Web site, February 27, 2026.

URL https://softwarepreservation.computerhistory.org/ALGOL/
applications.html.

Bulyonkov:20xx:AYU

[1270] Mikhail A. Bulyonkov, Alexandre F. Rar, and Andrey N. Terekhov. Algol 68 — 25 years in the USSR. Russian Virtual Computer Museum

Web site, 20xx. URL https://www.computer-museum.ru/english/
algol68.htm.

Dijkstra:20xx:EA

[1271] Edsger W. Dijkstra. To the EDITOR ALGOL 68. Circulated privately.,

20xx. URL http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD230.
PDF.

Yoneda:1969:NAL

[1272] Nobuo Yoneda. New algorithmic language ALGOL 68, volume 5–?? ????, ????, Japan, 1969. ?? pp.

Jakobson:1964:PSA

[1273] Roman Jakobson, editor. Proceedings of Symposia in Applied Mathematics. Volume XII: Structure of Language and Its Mathematical Aspects. American Mathematical Society, Providence, RI, USA, 1964. Roman Jakobson, editor.

Randell:1964:AIT

[1274] Brian Randell and L. J. (Lawford John) Russell. ALGOL 60 implementation; the translation and use of ALGOL 60 programs on a computer, volume 5 of A.P.I.C. studies in data processing. Academic Press, New York, USA, 1964. xiv + 418 pp. LCCN QA76.5 .R35 1964. With a foreword by E. W. Dijkstra. Published for the Automatic Programming Information Centre, Brighton College of Technology, England. Russian translation in 1967.

Bauer:1967:DA

[1275] Friedrich L. Bauer, Alston S. Householder, Frank W. J. Olver, Heinz Rutishauser, Klaus Samelson, and Eduard Stiefel. Description of AL-GOL 60, volume 1a of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-540-03826-4, 3-642-86934-3 (e-book), 3-642-86936-X (print), 3-662-38103-6. xii + 326 pp. LCCN QA76.5 .R87. URL

http://www.springerlink.com/content/978-3-642-86934-1.

Klerer:1967:DCU

[1276] Melvin Klerer and Granino A. Korn, editors. Digital Computer User’s Handbook. McGraw-Hill, New York, NY, USA, 1967. LCCN QA76.5 .K524.

Anonymous:1969:PIC

[1277] Anonymous, editor. Proceedings of an informal conference on ALGOL 68 implementation, 29-30 August 1969. Department of Computer Science, University of British Columbia, Vancouver, BC, Canada, 1969. LCCN QA76 .P68 1969.

Peck:1971:AIP

[1278] J. E. L. (John Edward Lancelot) Peck, editor. ALGOL 68 implementation; proceedings of the IFIP Working Conference on ALGOL 68 Implementation, Munich, July 20–24, 1970. North-Holland, Amsterdam, The Netherlands, 1971. ISBN 0-7204-2045-8. LCCN QA76.5 .I28 1970.

URL https://softwarepreservation.computerhistory.org/ALGOL/
paper/ALGOL_68-Implementation.pdf.

IEEE:1975:SCA

[1279] 3rd Symposium on Computer Arithmetic, November 19–20, 1975, Southern Methodist University, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1975. LCCN QA76.6.S919 1975. IEEE order number CH1017-3C.

King:1975:PIC

[1280] Peter R. King, editor. Proceedings of an International Conference on ALGOL 68 Implementation: Department of Computer Science, University of Manitoba, Winnipeg, June 18–20, 1974, volume 11 of Congressus numerantium. Utilitas Mathematica Publishers, Winnipeg, Manitoba, Canada, 1975. ISBN 0-919628-11-7. LCCN QA76.73.A24 I57 1974.

Rossler:1975:LEG

[1281] R. Rössler and K. Schenk, editors. LTPL-European Group Language-Comparison: a comparison of the properties of the programming languages ALGOL 68, Pearl … in relation to real-time programming. ????, Brussels, Belgium, 1975. ISBN 0-08-918884-5. LCCN ????

Schuman:1975:NDA

[1282] Stephen A. Schuman, editor. New directions in algorithmic languages 197: prepared for IFIP Working Group 2.1 on ALGOL [Proceedings of the second conference, held in Munich, West Germany on 25–29 August 1975]. Institut de Recherche d’Informatique et d’Automatique, Le Chesnay, France, 1975. ISBN 2-7261-0091-0 (paperback). LCCN QA76.7 .N40 1975; QA76.73 .A24 N49.

Bauer:1976:CCA

[1283] Friedrich Ludwig Bauer and Jürgen Eickel. Compiler Construction: an Advanced Course, volume 21 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1976. CODEN LNCSD9. ISBN 0-387-08046-5 (USA), 3-540-08046-5 (Berlin), 0-387-07542-9. ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 638 pp. LCCN QA76.6 .C628

1976. URL http://link.springer-ny.com/link/service/series/
0558/tocs/t0021.htm;   http://www.springerlink.com/content/
978-0-387-08046-8;   http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=21.

Jones:1976:CLL

[1284] Alan Jones and R. F. Churchhouse, editors. The computer in literary and linguistic studies: (proceedings of the Third International Symposium). University of Wales Press, Cardiff, Wales, 1976. ISBN 0-7083-0590-3 (hardcover). LCCN P98 .I58 1974; PN98.D.

Rayward-Smith:1976:PCA

[1285] V. J. Rayward-Smith, editor. Proceedings of the conference on Applications of ALGOL 68 held at the University of East Anglia, Norwich, England from Monday 22nd March, 1976 until Thursday 25th March 1976. [University of East Anglia, School of Computing Studies], Norwich, UK, 1976. ISBN 0-905706-00-5. LCCN QA76.73.A24 P76.

Rossler:1976:LCI

[1286] Roland Rössler, editor. A language comparison: International Purdue Workshop on Industrial Computer Systems; a comparison of the properties of the programming languages ALGOL 68, CAMAC-IML, CORAL 66, PAS 1, …. Purdue University, West Lafayette, IN, USA, 1976. LCCN ????

Schuman:1976:NDAb

[1287] Stephen A. Schuman, editor. New directions in algorithmic languages 1976: prepared for IFIP Working Group 2.1 on ALGOL [Proceedings of the second conference, held in St. Pierre-de-Chartreuse on 30 August–2 September 1976]. Institut de Recherche d’Informatique et d’Automatique, Rocquencourt, France, 1976. ISBN 2-7261-0124-0 (paperback). LCCN QA76.73.A24N49.

ACM:1977:PSA

[1288] ACM, editor. Proceedings of the Strathclyde ALGOL 68 Conference. Glasgow, UK; 29–31 March, 1977, volume 12(6) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cowell:1977:PNS

[1289] Wayne Cowell, editor. Portability of Numerical Software, Workshop, Oak Brook, Illinois, June 21–23, 1976, volume 57 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08446-0, 3-540-08446-0, 3-540-37071-4. ISSN 0302-9743. LCCN QA297 .W65 1976. URL http:

//link.springer.com/10.1007/3-540-08446-0.

Ershov:1977:MAL

[1290] A. P. (Andreýi Petrovich) Ershov and C. H. A. (Cornelis H. A.) Koster, editors. Methods of Algorithmic Language Implementation, volume 47 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. CODEN LNCSD9. ISBN 0-387-08065-1 (New York), 3-540-08065-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S96 1975; QA76.7.V7813 1975. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/t0047.
htm;   http://www.springerlink.com/content/978-0-387-08065-
9;   http://www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=47. English translation from the Russian of most of the proceedings (Trudy) of the symposium.

ACM:1978:ASH

[1291] ACM, editor. ACM SIGPLAN History of Programming Languages Conference. Los Angeles, CA, USA, 1–3 June, 1978, volume 13(8) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1979:PSS

[1292] Proceedings of the SIGPLAN Symposium on Compiler Construction: papers presented at the symposium, Denver, Colorado, August 6–10, 1979, SIGPLAN notices; v. 14, no. 8. ACM Press, New York, NY 10036, USA, 1979. LCCN QA 76.7 A1095 v.14 pt.2.

deBakker:1981:ALP

[1293] J. W. (Jacobus Willem) de Bakker and J. C. van Vliet, editors. Algorithmic languages: Proceedings of the International Symposium on …, Amsterdam, The Netherlands, 26–29 October 1981, a tribute to prof. A. van Wijngaarden on the occasion of his retirement from the Mathematical Centre. North-Holland, Amsterdam, The Netherlands, 1981. ISBN 0-444-86285-4. LCCN QA76.7 .I575 1981. URL https://ir.cwi.nl/

pub/34328/34328D.pdf.

Halpern:1981:ARAb

[1294] M. I. Halpern, Andrew D. McGettrick, and B. Cronhjort, editors. Introduction to the formal definition of ALGOL 68, volume 9 of Annual review in automatic programming. Pergamon, New York, NY, USA and Oxford, UK, 1981. ISBN 0-08-020242-X. LCCN QA76 .A63 vol. 9.

IEEE:1983:PSC

[1295] Proceedings: 6th Symposium on Computer Arithmetic, June 20–22, 1983, Aarhus University, Århus, Denmark. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. ISBN 0-8186-4476-1 (paperback), 0-8186-8476-3 (hardcover), 0-8186-6476-2 (microfiche), 0-8186-0034-9 (hardcover). LCCN QA 76.9 C62 S95

1983. URL http://www.acsel-lab.com/arithmetic/arith6/papers/.

IEEE catalog number 83CH1892-9. IEEE Computer Society order number 476.

Cowell:1984:SDM

[1296] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

ACM:1988:PPC

[1297] ACM, editor. POPL ’88. Proceedings of the conference on Principles of programming languages, January 13–15, 1988, San Diego, CA. ACM Press, New York, NY 10036, USA, 1988. ISBN ???? LCCN

???? URL http://www.acm.org/pubs/contents/proceedings/plan/
73560/index.html.

ACM:1991:PPE

[1298] ACM, editor. POPL ’91. Proceedings of the eighteenth annual ACM symposium on Principles of programming languages, January 21–23, 1991, Orlando, FL. ACM Press, New York, NY 10036, USA, 1991. ISBN ???? LCCN ????

ACM:1993:ASH

[1299] ACM, editor. ACM SIGPLAN HOPL-II. 2nd ACM SIGPLAN History of Programming Languages Conference (Preprints), volume 28(3) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, March 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

ACM:1993:CRT

[1300] ACM, editor. Conference record of the Twentieth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Charleston, South Carolina, January 10–13, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover).

LCCN QA76.7 .A15 1993. URL http://www.acm.org/pubs/contents/
proceedings/plan/158511/index.html. ACM order number 549930.

Alberts:1993:CHA

[1301] G. Alberts, editor. Conference on the history of ALGOL 68. Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, January

1993. URL https://ir.cwi.nl/pub/10854/10854D.pdf. Historical

Note AM-NH9301.

ACM:1995:CRP

[1302] ACM, editor. Conference record of POPL ’95, 22nd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: San Francisco, California, January 22–25, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-692-1.

LCCN QA 76.7 A11 1995. URL http://www.acm.org/pubs/contents/
proceedings/plan/199448/index.html. ACM order number: 549950.

Bergin:1996:HPL

[1303] Thomas J. Bergin, Jr. and Richard G. Gibson, Jr., editors. History of Programming Languages II. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1996. ISBN 0-201-89502-1. xvi + 864 pp. LCCN QA76.7 .H57 1996; QA76.7 .H558 1996. Drawn from the Second ACM SIGPLAN History of Programming Languages Conference.

Tomilin:2014:TIC

[1304] A. N. Tomilin, editor. 2014 Third International Conference on Computer Technology in Russia and in the Former Soviet Union (SoRu-Com). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-1799-0, 1-4799-1800-

8. LCCN QA75.5 .I578 2014. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=7031478.